### v1

# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

Abstract
In sequential event prediction, which finds applications in finance, retail, social networks, and healthcare, a crucial task is forecasting multiple future events within a specified time horizon. Traditionally, this has been addressed through autoregressive generation using next-event prediction models, such as Marked Temporal Point Processes. However, autoregressive methods use their own output for future predictions, potentially reducing quality as the prediction horizon extends. In this paper, we challenge traditional approaches by introducing a novel benchmark, HoTPP, specifically designed to evaluate a model’s ability to predict event sequences over a horizon. This benchmark features a new metric inspired by object detection in computer vision, addressing the limitations of existing metrics in assessing models with imprecise time-step predictions. Our evaluations on established datasets employing various models demonstrate that high accuracy in next-event prediction does not necessarily translate to superior horizon prediction, and vice versa. HoTPP aims to serve as a valuable tool for developing more robust event sequence prediction methods, ultimately paving the way for further advancements in the field. The code is available at https://github.com/ivan-chai/hotpp-benchmark.
1 Introduction
The world is full of events. Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information. In the era of AI, it is crucial to develop methods capable of handling these complex data streams. We refer to this type of data as Event Sequences (ESs). Event sequences differ fundamentally from other data types. Unlike tabular data [wang2022transtab], ESs include timestamps and possess an inherent order. In contrast to time series data [lim2021timeseriessurvey], ESs are characterized by irregular time intervals and structured descriptions of each event. These differences necessitate the development of specialized models and evaluation practices.
The primary task in the domain of ESs is sequence modeling, specifically predicting the next event type and its occurrence time. Indeed, the ability to accurately forecast sequential events is vital for applications such as stock price prediction, personalized recommendation systems, and early disease detection. The simplified domain, represented as pairs of event types and times, is typically called Marked Temporal Point Processes (MTPP) [rizoiu2017hawkes-tutorial]. Additionally, structured modeling of dependencies between different data fields [mcdermott2024eventstreamgpt] can be considered an extension of MTPP.
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events poses unique challenges distinct from traditional next-event prediction tasks. The conventional approach to this problem typically involves the autoregressive application of next-event prediction models. While these models have demonstrated effectiveness in predicting the immediate next event, their performance tends to decline as the prediction horizon extends. This results in increased computational costs and reduced predictive accuracy over longer sequences. Despite this, it is commonly assumed that the quality of next-event prediction directly influences the quality of the corresponding autoregressive model.
In this paper, we challenge this belief and demonstrate that autoregressive models can be suboptimal for long-horizon predictions in some cases. To address this, we introduce HoTPP (Horizon Temporal Point Process), a novel benchmark designed to evaluate the effectiveness of event sequence prediction models over horizons. Our benchmark leverages a new metric, T-mAP (Temporal mean Average Precision), inspired by object detection methodologies in computer vision. We demonstrate that T-mAP offers a more nuanced assessment of a model’s performance in scenarios where precise time step predictions are challenging. This innovation addresses the inherent limitations of traditional metrics, which often fail to capture the complexities involved in long-term event forecasting.
Thus, our main contributions can be summarized as follows:
  1. 1.
    We introduce HoTPP, a new open-source benchmark designed to facilitate research on long horizon Event Sequences (ESs) prediction. This benchmark integrates datasets and methods from various domains, including financial transactions, social networks, medicine, and recommender systems.
  2. 2.
    We demonstrate that widely used evaluation methods for ESs and MTPPs often overlook critical aspects of model performance. Notably, we show that simple statistical baselines can sometimes outperform popular methods in terms of Optimal Transport Distance (OTD). To address the shortcomings of OTD, we introduce a new evaluation metric for MTPP called Temporal mAP (T-mAP) (Fig. 1).
  3. 3.
    We show that the length of the predicted sequence significantly impacts long-horizon prediction performance and must be carefully tuned. We also provide a method for selecting this hyperparameter.
  4. 4.
    We compare popular approaches from different domains and show that intensity-based MTPP modeling methods can be suboptimal compared to simple approaches. We also show that autoregressive approaches can be outperformed by direct prediction of multiple future events in some cases.
Figure 1: Temporal mAP (T-mAP) evaluation pipeline. Unlike previous methods, T-mAP evaluates sequences of variable lengths over the prediction horizon. It also enhances performance assessment by analyzing label distributions rather than fixed predictions.
2 Background
2.1 Marked Temporal Point Processes
Marked Temporal Point Process (MTPP) is a stochastic process that consists of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ are event type labels. Common MTPP modeling approaches primarily focus on predicting the next event. A straightforward solution is to predict time and type independently. A more advanced approach involves splitting the original sequence into subsequences, one for each event type, and then independently modeling each subsequence’s time. In both approaches, the fundamental problem is modeling the arrival time of the next event. The stochastic process that generates these times is typically called a Temporal Point Process [rizoiu2017hawkes-tutorial]. For the details on MTPP modeling, refer to Appendix A.
2.2 Evaluation
MTPPs are typically evaluated based on the quality of next-event predictions. Time and type predictions are usually assessed independently. The quality of type prediction is characterized by the error rate, while time prediction error is measured using either the Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE). Recent studies have advanced the measurement of horizon prediction quality by employing the Optimal Transport Distance (OTD). The OTD metric is analogous to the edit distance between the predicted event sequence and the ground truth. Below, we provide the formal definition of OTD.
Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. If $M(S_{p},S_{gt})$ is a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$, OTD is computed as follows:
$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost. It is common to take $C_{ins}=C_{del}$ [mei2019imputing]. It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality. However, OTD has practical limitations, which are described in the following section.
3 Limitations of the Next-Event and OTD Metrics
Figure 2: Comparison of OTD and T-mAP metrics. In example 1.a, prefix evaluation of three events using OTD results in an incorrect false positive. T-mAP addresses this issue by comparing all events within the time horizon, as shown in example 1.b. Additionally, OTD evaluates only the label with the maximum probability. From the OTD perspective, cases 2.a and 2.b have the same quality. In contrast, T-mAP evaluates the entire distribution, yielding a low score of 0.33 in case 2.a and a high score of 1 in case 2.b.
A predominant evaluation metric in the MTPP domain is the quality of next-event prediction. This metric typically comprises two distinct components: the accuracy (or error rate) of the next label prediction and either Mean Absolute Error (MAE) or Root Mean Square Error (RMSE) for time step prediction. However, these metrics do not assess the model’s ability to predict multiple future events. For instance, autoregression uses its own outputs as inputs for the next step, which can potentially lead to cumulative prediction errors. Consequently, long-horizon evaluation metrics are necessary.
The OTD metric extends evaluation to sequences by comparing fixed-size prefixes, but this approach limits its flexibility in assessing models with imprecise time step predictions, as illustrated in Fig. 2.1. If the model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives. This outcome is inaccurate because allowing the model to predict more events would alter the evaluation. Conversely, if the time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of false positives. This is undesirable as it prevents the inclusion of additional ground truth events in the evaluation. Therefore, evaluating a dynamic number of events is essential to properly align with the time horizon in the ground truth.
Another limitation of the OTD metric is its inflexibility in evaluating predicted distributions of labels, as shown in Fig. 2.2. OTD evaluates only the labels with the highest probability, disregarding the full distribution. As a result, it relies on the calibration of model outputs and cannot assess the performance beyond a limited set of popular labels. However, models typically predict probabilities for all classes, allowing for an independent assessment of prediction quality for each label. Therefore, we aim to develop an evaluation metric that accurately reflects the performance across common and rare classes.
4 Temporal mAP: Prediction as Detection
In this section, we introduce a novel metric, Temporal mAP (T-mAP), inspired by object detection methods from computer vision [everingham2010pascalcvdetection], as illustrated in Fig. 1. The goal of object detection is to localize objects within an image and identify their types. In event sequences, we tackle a similar problem, but we consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
T-mAP incorporates concepts from OTD but addresses its limitations, as outlined in the previous section. Firstly, T-mAP evaluates label probabilities instead of relying on final predictions. Secondly, T-mAP restricts the prediction horizon rather than the number of events. These adjustments result in significant differences in formulation and computation, detailed below.
4.1 Definition
T-mAP is parameterized by the horizon length $T$ and the maximum allowed time error $\delta$. T-mAP compares predicted and ground truth sequences within the interval $T$ from the last observed event. Consider a simplified scenario with a single event type $l$. Assume the model predicts timestamps and presence scores (logits or probabilities) for several future events: $S^{l}_{p}=\{(t^{p}_{i},s^{p}_{i})\},1\leq i\leq n_{p}$. The corresponding ground truth sequence is $S^{l}_{gt}=\{t^{gt}_{j}\},1\leq j\leq n_{gt}$. For simplicity, assume the sequences $S^{l}_{p}$ and $S^{l}_{gt}$ are filtered to include only events within the horizon $T$.
For any threshold value $h$, we can select a subset of the predicted sequence $S^{l}_{p}$ with scores exceeding the threshold:
$S^{l}_{>}(h)=\{t^{p}_{i}:\exists(t^{p}_{i},s^{p}_{i})\in S^{l}_{p},s^{p}_{i}>h\}.$ (2)
A prediction $i$ can be matched with ground truth $j$ iff $|t^{p}_{i}-t^{gt}_{j}|<\delta$. T-mAP identifies the matching that maximizes precision and recall, matching with maximum cover $c$. The precision of this matching is calculated as $c/|S^{l}_{>}(h)|$ and the recall as $c/|S^{l}_{gt}|$. For any threshold $h$, we can count true positives (TP), false positives (FP), and false negatives (FN) across all predicted and ground truth sequences. Note that there are no true negatives, as the model cannot explicitly predict the absence of an event. Similarly to binary classification, we can define a precision-recall curve by varying the threshold $h$ and estimating Average Precision (AP), the area under the precision-recall curve. T-mAP is the mean AP across all possible event types $l$.
4.2 Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. In this section, we show how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm [jonker1988linsum]. The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
Figure 3: T-mAP computation pipeline.
For each pair of sequences $S^{l}_{p}$ and $S^{l}_{gt}$ we define a weighted bipartite graph $\mathcal{G}(S^{l}_{p},S^{l}_{gt})$ with $|S^{l}_{p}|$ vertices in the first and $|S^{l}_{gt}|$ vertices in the second part. For each pair of prediction $i$ and ground truth event $j$ with $|t^{p}_{i}-t^{gt}_{j}|<\delta$ we add an edge with weight $-s^{p}_{i}$, as shown in Fig. 3.a. Jonker-Volgenant algorithm finds the matching with the maximum number of edges in the graph, such that the resulting matching minimizes the total cost of selected edges, as shown in Fig. 3.b. We call this matching optimal matching and denote a set of all optimal matching possibilities as $M(\mathcal{G})$. For any threshold $h$, there is a subgraph $\mathcal{G}_{h}(S^{l}_{>}(h),S^{l}_{gt})$ with events whose scores are greater than $h$. The following theorem holds:
Theorem 4.1.
For any threshold $h$ there exists an optimal matching in the graph $\mathcal{G}_{h}$, that is a subset of an optimal matching in the full graph $\mathcal{G}$:
$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (3)
According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig. 3.c. The proof of the theorem and the complete algorithm for T-mAP evaluation are provided in Appendix B.
5 HoTPP Benchmark
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP differs from prior MTPP benchmarks by including simple statistical baselines and next-k models, which simultaneously predict multiple future events. The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, and reproducibility.
Simplicity and Extensibility. The benchmark code is clearly structured, separating the library, dataset-specific scripts, and configuration files. New methods can be implemented at multiple levels: configuration file, model architecture, loss function, metric, and training module. Core components can be reused by applying Hydra configuration library [Yadan2019Hydra].
Evaluation Stability. Unlike previous works [xue2023easytpp, xue2022hypro], we evaluate methods not only at the end of each sequence but also at intermediate points. This approach collects a larger set of predictions, reducing the final metric’s STD. This is particularly important for datasets with few sequences, such as StackOverflow and Amazon.
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library [Falcon_PyTorch_Lightning_2019] for training with a specified random seed, and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
A detailed benchmark description can be found in Appendix C.
Table 1: Evaluation results. The best result is shown in bold. Mean and STD values of 5 runs with different random seeds are reported.
Method Mean length Next-item Acc., % Next-item MAE OTD Val / Test T-mAP, % Val / Test Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.09$\pm$ 0.00 / 1.02$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.924$\pm$ 0.000 7.51$\pm$ 0.00 / 7.55$\pm$ 0.00 1.64$\pm$ 0.00 / 1.69$\pm$ 0.00 MAE-CE 14.0 38.18$\pm$ 0.03 0.570$\pm$ 0.001 7.18$\pm$ 0.01 / 7.20$\pm$ 0.01 4.63$\pm$ 0.33 / 4.50$\pm$ 0.27 MAE-CE-K 9.3 37.68$\pm$ 0.01 0.573$\pm$ 0.000 7.10$\pm$ 0.02 / 7.10$\pm$ 0.02 3.34$\pm$ 0.16 / 3.73$\pm$ 0.16 RMTPP 11.5 38.29$\pm$ 0.02 0.658$\pm$ 0.001 6.78$\pm$ 0.01 / 6.80$\pm$ 0.01 6.72$\pm$ 0.35 / 6.49$\pm$ 0.17 RMTPP-K 9.2 37.74$\pm$ 0.03 0.664$\pm$ 0.001 7.11$\pm$ 0.00 / 7.12$\pm$ 0.00 4.79$\pm$ 0.23 / 4.96$\pm$ 0.13 NHP 4.7 30.49$\pm$ x.xx 0.995$\pm$ x.xxx 7.68$\pm$ x.xx / 7.69$\pm$ x.xx 1.62$\pm$ x.xx / 1.51$\pm$ x.xx HYPRO 11.4 38.27$\pm$ x.xx 0.658$\pm$ x.xxx 6.81$\pm$ x.xx / 6.82$\pm$ x.xx 6.72$\pm$ x.xx / 6.51$\pm$ x.xx MIMIC-IV MostPopular 8.3 1.79$\pm$ 0.00 11.23$\pm$ 0.00 19.89$\pm$ 0.00 / 19.88$\pm$ 0.00 0.73$\pm$ 0.00 / 0.75$\pm$ 0.00 Last 5 4.2 0.88$\pm$ 0.00 4.77$\pm$ 0.00 19.77$\pm$ 0.00 / 19.74$\pm$ 0.00 2.31$\pm$ 0.00 / 2.38$\pm$ 0.00 MAE-CE 10.8 57.38$\pm$ 0.05 2.85$\pm$ 0.01 11.69$\pm$ 0.00 / 11.70$\pm$ 0.01 23.48$\pm$ 0.20 / 23.60$\pm$ 0.20 MAE-CE-K 15.3 56.71$\pm$ 0.22 2.89$\pm$ 0.02 14.04$\pm$ 0.24 / 14.01$\pm$ 0.25 21.26$\pm$ 0.16 / 21.03$\pm$ 0.11 RMTPP 15.9 57.35$\pm$ 0.05 2.85$\pm$ 0.01 12.93$\pm$ 0.03 / 12.93$\pm$ 0.03 20.61$\pm$ 0.15 / 20.85$\pm$ 0.23 RMTPP-K 16.0 56.69$\pm$ 0.03 2.79$\pm$ 0.00 13.91$\pm$ 0.04 / 13.88$\pm$ 0.03 23.01$\pm$ 0.04 / 22.85$\pm$ 0.06 NHP 8.5 24.71$\pm$ 0.95 6.72$\pm$ 0.35 18.73$\pm$ 0.19 / 18.69$\pm$ 0.20 8.55$\pm$ 1.76 / 8.57$\pm$ 1.75 HYPRO 16.0 57.18$\pm$ x.xx 2.83$\pm$ x.xx 12.93$\pm$ x.xx / 12.94$\pm$ x.xx 21.47$\pm$ x.xx / 21.50$\pm$ x.xx Retweet MostPopular 15.5 58.50$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 26.51$\pm$ 0.00 / 25.87$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 21.93$\pm$ 0.00 165.5$\pm$ 0.0 / 164.0$\pm$ 0.0 28.54$\pm$ 0.00 / 28.61$\pm$ 0.00 MAE-CE 15.0 59.92$\pm$ 0.04 18.28$\pm$ 0.03 173.6$\pm$ 4.3 / 173.2$\pm$ 4.6 33.10$\pm$ 4.84 / 31.35$\pm$ 4.71 MAE-CE-K 14.8 59.30$\pm$ 0.33 18.25$\pm$ 0.45 169.7$\pm$ 2.2 / 169.3$\pm$ 2.6 36.34$\pm$ 5.06 / 34.09$\pm$ 5.46 RMTPP 12.9 60.06$\pm$ 0.04 18.54$\pm$ 0.09 169.1$\pm$ 2.2 / 167.9$\pm$ 2.3 50.98$\pm$ 0.88 / 48.35$\pm$ 1.01 RMTPP-K 13.6 60.03$\pm$ 0.05 18.42$\pm$ 0.21 166.4$\pm$ 4.7 / 165.8$\pm$ 4.7 49.21$\pm$ 1.06 / 46.31$\pm$ 1.25 NHP 10.7 59.84$\pm$ 0.16 27.37$\pm$ 6.91 194.4$\pm$ 14.6 / 192.1$\pm$ 13.8 39.71$\pm$ 9.02 / 37.52$\pm$ 8.56 HYPRO 13.0 60.12$\pm$ x.xx 18.55$\pm$ x.xx 171.6$\pm$ x.x / 170.6$\pm$ x.x 51.79$\pm$ x.xx / 49.60$\pm$ x.xx Amazon MostPopular 15.7 33.46$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 9.34$\pm$ 0.00 / 9.02$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 0.323$\pm$ 0.000 6.70$\pm$ 0.00 / 6.67$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 MAE-CE 13.2 35.73$\pm$ 0.12 0.242$\pm$ 0.003 6.62$\pm$ 0.06 / 6.55$\pm$ 0.04 21.98$\pm$ 0.36 / 22.58$\pm$ 0.37 MAE-CE-K 13.3 35.10$\pm$ 0.08 0.246$\pm$ 0.000 6.74$\pm$ 0.01 / 6.70$\pm$ 0.01 21.84$\pm$ 0.05 / 22.35$\pm$ 0.07 RMTPP 15.3 35.74$\pm$ 0.04 0.294$\pm$ 0.001 6.68$\pm$ 0.02 / 6.62$\pm$ 0.03 19.92$\pm$ 0.26 / 20.33$\pm$ 0.33 RMTPP-K 15.8 35.10$\pm$ 0.13 0.300$\pm$ 0.001 6.95$\pm$ 0.01 / 6.89$\pm$ 0.01 17.58$\pm$ 0.12 / 17.83$\pm$ 0.13 NHP 14.4 22.38$\pm$ 6.15 0.378$\pm$ 0.073 7.68$\pm$ 0.12 / 7.67$\pm$ 0.14 21.85$\pm$ 0.86 / 21.34$\pm$ 0.86 HYPRO 14.9 35.87$\pm$ x.xx 0.293$\pm$ x.xxx 6.69$\pm$ x.xx / 6.64$\pm$ x.xx 20.40$\pm$ x.xx / 20.96$\pm$ x.xx StackOverflow MostPopular 11.8 42.90$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.33$\pm$ 0.00 / 5.86$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 0.950$\pm$ 0.000 14.87$\pm$ 0.00 / 14.92$\pm$ 0.00 8.61$\pm$ 0.00 / 6.64$\pm$ 0.00 MAE-CE 15.7 45.39$\pm$ 0.11 0.641$\pm$ 0.002 13.58$\pm$ 0.08 / 13.65$\pm$ 0.05 8.82$\pm$ 0.70 / 8.45$\pm$ 0.38 MAE-CE-K 15.3 44.84$\pm$ 0.24 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.52$\pm$ 0.06 12.15$\pm$ 0.99 / 11.18$\pm$ 0.79 RMTPP 12.5 45.44$\pm$ 0.13 0.700$\pm$ 0.007 12.97$\pm$ 0.03 / 13.18$\pm$ 0.05 13.44$\pm$ 0.23 / 12.95$\pm$ 0.11 RMTPP-K 12.0 44.84$\pm$ 0.08 0.687$\pm$ 0.001 12.93$\pm$ 0.01 / 13.13$\pm$ 0.02 14.79$\pm$ 0.45 / 14.19$\pm$ 0.26 NHP 10.4 44.44$\pm$ 0.06 0.758$\pm$ 0.021 13.32$\pm$ 0.16 / 13.59$\pm$ 0.16 12.05$\pm$ 0.64 / 11.35$\pm$ 0.42 HYPRO 11.8 45.42$\pm$ x.xx 0.716$\pm$ x.xxx 13.05$\pm$ x.xx / 13.25$\pm$ x.xx 13.51$\pm$ x.xx / 13.61$\pm$ x.xx
6 Experiments
We conduct experiments on five datasets of varying sizes and origins: Transaction financial dataset [agepred], MIMIC-IV medical dataset [johnson2020mimic], two social networks datasets, Retweet [zhao2015seismic] and StackOverflow [jure2014snap], and Amazon reviews dataset[jure2014amazon]. A detailed description of these datasets, along with the training and evaluation parameters, is provided in Appendix C.4.
We evaluate representative modeling methods from different groups:
  1. 1.
    Statistical baselines. MostPopular Generates a constant output with the most popular label from the prefix and the average time step. Last copies multiple previous events with adjusted timestamps.
  2. 2.
    Intensity-free approaches. We implement MAE-CE method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels [padhi2021tabgpt, mcdermott2024eventstreamgpt].
  3. 3.
    Intensity-based approaches. We implement RMTPP [du2016rmtpp] method as an example of the TPP approach with a traditional RNN. We also add NHP [mei2017nhp], based on a continuous time LSTM architecture.
  4. 4.
    Next-K approaches. We adapt MAE-CE and RMTPP for a direct prediction of multiple future events. This approach originates in time series analysis and has not been previously applied in the MTPP domain. The resulting methods are called MAE-CE-K and RMTPP-K, respectively.
  5. 5.
    Reranking. We also implement HYPRO [xue2022hypro], which generates multiple hypotheses and selects the best sequence using a contrastive approach.
Details on the methods can be found in Appendix A and Appendix C.1. Training specifics are provided in AppendixD. Hyperparameters for both methods and metrics are presented in Appendix E. The main evaluation results are presented in Table 1. Below, we will discuss these results from various perspectives and provide an additional analysis of the methods’ behavior.
6.1 OTD vs T-mAP
Evaluation results indicate that the OTD metric can yield low values even for simple statistical baselines. For instance, the Last baseline achieves the lowest OTD value on the Retweet dataset and outperforms Next-K methods on the Amazon dataset. In contrast, the T-mAP metric clearly differentiates between statistical baselines and deep models, as the former is incapable of predicting label probabilities or confidences.
6.2 Intensity-Based vs Intensity-Free
Previous works have compared either intensity-based neural methods [du2016rmtpp, xue2022hypro] or intensity-free approaches [padhi2021tabgpt, mcdermott2024eventstreamgpt]. This raises the question: which approach is superior? We compare the intensity-free MAE-CE method with intensity-based RMTPP and NHP approaches. Our results show that the intensity-free MAE-CE method excels in the next-event MAE prediction, as it optimizes this metric. In other scenarios, the results vary between datasets. Therefore, we conclude that both approaches warrant attention in future research.
6.3 Autoregression vs Direct Horizon Prediction
We compared MAE-CE and RMTPP methods with their Next-K counterparts. According to our results, autoregression generally performs better on Transactions, MIMIC-IV, and Amazon datasets. However, on social network datasets (Retweet and StackOverflow), Next-K models typically outperform autoregression in horizon prediction tasks. Despite this, little effort has been made to adapt Next-K models to MTPPs. Consequently, we recommend further research into this type of method.
We also observed that the Next-K approaches exhibit some interesting properties compared to autoregression. For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Fig. 4.
Figure 4: Entropy of label distributions as a function of the position in the generated sequence. Results for StackOverflow and Amazon datasets are provided in AppendixF.
The likely reason for this behavior is the dependency of the model’s future output on its own past errors. Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict. This suggests that Next-K models have the potential to better predict the distributions of future labels, a factor that should be considered in future research.
6.4 The Maximum Sequence Length
We observed that long-horizon prediction quality largely depends on the maximum number of allowed predictions in autoregression and Next-K models. As shown in Fig. 5, the optimal number of predicted events is usually less then the maximum number of events in the horizon. For example, in Transactions dataset, the optimal length ranges from 3 to 5, depending on the method. This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to manually limit the number of predictions in the target metric.
Figure 5: T-mAP dependency on the maximum length of the predicted sequence. Results for the StackOverflow and Amazon datasets are detailed in Appendix G.
We therefore conclude that more attention should be paid to probability calibration between generation steps. This result may also indicate real progress in long-horizon prediction tasks, specifically the maximum horizon one can model accurately. T-mAP is the only metric capable of evaluating the optimal predicted sequence length, whereas the OTD metric does not depend on sequence length as it compares prefixes of fixed size.
7 Related Work
Temporal Point Processes. Traditional MTPP models make strong assumptions about the underlying generative process, resulting in Poisson and Hawkes processes [rizoiu2017hawkes-tutorial]. Over the last decade, there has been a shift towards achieving greater model flexibility by applying neural architectures. Some works have utilized classical RNNs [du2016rmtpp] and transformers [zuo2020thp], while others have proposed architectures with continuous time [mei2017nhp, rubanova2019ode]. In the HoTPP benchmark, we evaluate MTPP models with discrete and continuous time. Unlike other benchmarks, we also evaluate popular approaches from related fields, including next-event prediction models from event sequences [mcdermott2024eventstreamgpt, padhi2021tabgpt] and Next-K models from time series analysis [lim2021timeseriessurvey].
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT [mcdermott2024eventstreamgpt] and TemporAI [saveliev2023temporai] consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick [bacry2017tick] and PyHawkes [^1] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches and are unable to work with modern neural networks. PoPPy [xu2018poppy] and EasyTPP [xue2023easytpp] implement neural methods; however, these benchmarks cannot be extended to use Next-K approaches. PoPPy does not evaluate long-horizon prediction at all. This limitation is partially addressed in EasyTPP, which measures the OTD metric at the end of each sequence. In this work, we address the limitations of the OTD metric and introduce a superior alternative, T-mAP, along with an effective GPU-accelerated implementation.
8 Limitations
The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating modern Transformer approaches[zuo2020thp] or Intensity-free training objective[shchur2019intensityfree].
9 Future Research Directions
In this work, we analyze long-horizon prediction evaluation and propose a new validation process. We strongly encourage future research to develop improved techniques for predicting future events and to establish simple baselines for better-measuring progress in the field. For example, Next-K models, which predict multiple future events without autoregression, show promising results in long-horizon prediction tasks. We suggest further exploration of these models. We also highlight the importance of label distribution estimation and emphasize the need to improve confidence estimation and calibration.
10 Conclusion
In this paper, we propose the HoTPP benchmark to assess the quality of multiple events forecasting. Our extensive evaluations using established datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction does not necessarily correlate with superior performance in horizon prediction. This finding underscores the need for benchmarks like HoTPP, which emphasize the importance of long-horizon accuracy and robustness. By shifting the focus from short-term to long-term predictive capabilities, HoTPP aims to drive the development of more sophisticated and reliable event sequence prediction models. This, in turn, has the potential to significantly enhance the practical applications of sequential event prediction in various domains, fostering innovation and improving decision-making processes across a wide range of industries.
[…]
Appendix A Background on Marked Temporal Point Processes Modeling
[…]
The derivation is provided in [rizoiu2017hawkes-tutorial]. Different TPPs are characterized by their intensity function $\lambda(t_{i})$. In Poisson and non-homogeneous Poisson processes, the intensity function is independent of previous events, meaning event occurrences depend solely on external factors. In contrast, self-exciting processes are characterized by previous events increasing the intensity of future events. A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
$\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$ (5)
where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function. Given the intensity function $\lambda(t_{i})$, predictions are typically made by sampling or expectation estimation. Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach. For details on the implementation of sampling, please refer to [rizoiu2017hawkes-tutorial].
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models, like NHP [mei2017nhp] and Neural ODEs [rubanova2019ode]. The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
Intensity-free modeling. Some methods evaluate the next event time distribution without using intensity functions. For example, intensity-free [shchur2019intensityfree] represents the distribution as a mixture of Gaussians or through normalizing flows. Other approaches do not predict the distribution directly but instead, solve a regression problem using MAE or RMSE loss. However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction. Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution [bishop1994mdn]. In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
Rescoring with HYPRO. HYPRO [xue2022hypro] is an extension applicable to any sequence prediction method capable of sampling. HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample. It is trained with a contrastive loss to distinguish between the generated sequence and the ground truth. HYPRO generates multiple sequences with a background model during inference and selects the best one by maximizing the estimated score. Although HYPRO is intended to improve quality compared to simple sampling, it is unclear whether HYPRO outperforms expectation-based prediction. In our work, we apply HYPRO to the outputs of the RMTPP model.
Appendix B T-mAP Computation and Proofs
Definitions and scope. In this section, we consider weighted bipartite graphs. The first part consists of predictions, and the second consists of ground truth events. We assume that all edges connected to the same prediction have the same weight. A matching is a set of edges $(a_{i},b_{i})$ between predictions and ground truth events. A matching is termed optimal if it (1) has the maximum size among all possible matching and (2) has the minimum total weight among all matches of that size. We denote all optimal matchings in the graph $\mathcal{G}$ as $M(\mathcal{G})$.
Lemma A.
Consider a graph $\mathcal{G}^{\prime}$, constructed from a graph $\mathcal{G}$ by adding one ground truth vertex with corresponding weighted edges. Then any optimal matching $m^{\prime}\in M(\mathcal{G}^{\prime})$ will either have a size greater than the sizes of matchings in $M(\mathcal{G})$ or have a total weight equal to that of matchings in $M(\mathcal{G})$.
Proof.
Figure 6: Illustration of the iterative process in Lemma A.
If the size of $m^{\prime}$ is greater than the size of matchings in $M(\mathcal{G})$ then the lemma holds. The optimal matching in the extended graph $\mathcal{G}^{\prime}$ cannot have a size smaller than the optimal matching in $\mathcal{G}$. Now, consider the case when both matchings have the same size. We will show that $m^{\prime}$ has a total weight equal to that of the matchings in $M(\mathcal{G})$.
Denote $b$ the new ground truth event in graph $\mathcal{G}^{\prime}$. If $m^{\prime}$ does not include vertex $b$, then $m^{\prime}$ is also optimal in $\mathcal{G}$, and the lemma holds. If $m^{\prime}$ includes vertex $b$, then we can walk through the graph with the following process:
  1. 1.
    Start from the new vertex $b$.
  2. 2.
    If we are at a ground truth event $b_{i}$ and there is an edge $(a_{i},b_{i})$ in matching $m^{\prime}$, move to the vertex $a_{i}$.
  3. 3.
    If we are at a prediction $a_{i}$ and there is an edge $(a_{i},b_{i})$ in matching $m$, move to the vertex $b_{i}$.
Example processes are illustrated in Fig. 6. The vertices in each process do not repeat; otherwise, there would be two edges in either $m$ or $m^{\prime}$ with the same vertex, which contradicts the definition of a matching.
If the process finishes at a predicted event, then $m^{\prime}$ has a size greater than the size of $m$, and the lemma holds. If the process finishes at a ground truth event, we can replace a part of matching $m^{\prime}$ traced with a part of matching $m$. The resulting matching $\hat{m}$ will have a size equal to both $m$ and $m^{\prime}$ and will have a total weight equal to both matchings. This proves the lemma. ∎
Theorem 4.1.
For any threshold $h$, there exists an optimal matching in the graph $\mathcal{G}_{h}$ such that it is a subset of an optimal matching in the full graph $\mathcal{G}$:
$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (6)
Proof.
If the threshold is lower than any score $s^{p}_{i}$, then $\mathcal{G}_{h}=\mathcal{G}$, and $m_{h}=m$ satisfies the theorem. Otherwise, some predictions are filtered by the threshold and $\mathcal{G}_{h}\subset\mathcal{G}$.
Without the loss of generality, assume that the threshold is low enough to filter out only one prediction. Otherwise, we can construct a chain of thresholds $h_{1}<h_{2}<\dots<h$, with each subsequent threshold filtering an additional vertex, and apply the theorem iteratively.
Denote $i$ the only vertex present in $\mathcal{G}$ and filtered from $\mathcal{G}_{h}$. If there is no edge containing vertex $i$ in the matching $m(\mathcal{G})$, then $m(\mathcal{G})$ is also the optimal matching in $\mathcal{G}_{h}$, and theorem holds.
Consider the case when matching $m(\mathcal{G})$ contains an edge $(i,j)$. Let $\hat{m}_{h}=m(\mathcal{G})\setminus\{(i,j)\}$, i.e., the matching without edge $(i,j)$. If it is optimal for the graph $\mathcal{G}_{h}$, then it satisfies the theorem. Otherwise, there is an optimal matching $m_{h}$ with either (a) more vertices or (b) a smaller total weight than $\hat{m}_{h}$.
In (a), matching $m_{h}$ has a size greater than the size of $\hat{m}_{h}$. It follows, that $m_{h}$ has the maximum size in the full graph $\mathcal{G}$. If the total weight of $m_{h}$ is larger or equal to the optimal weight in $\mathcal{G}$, then it can not be less than the total weight of $\hat{m}_{h}$, which contradicts our assumption. At the same time, the total weight of $m_{h}$ can not be less than the optimal weight in the full graph $\mathcal{G}$. It follows that case (a) leads to a contradiction.
Consider the case (b) when the size of $m_{h}$ equals $\hat{m}_{h}$. If $m_{h}$ does not include vertex $j$, then both $m_{h}$ and $\hat{m}_{h}$ are optimal. Otherwise, we can remove vertex $j$ from $\mathcal{G}_{h}$ and construct a new optimal matching $m^{\prime}_{h}$ without $j$ by using Lemma A, and again, both matchings $m^{\prime}_{h}$ and $\hat{m}_{h}$ are optimal. This concludes the proof. ∎
Algorithm. Using the theorem, we can introduce an effective algorithm for T-mAP computation. T-mAP is defined on a batch of predictions $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and ground truth sequences $\{\mathcal{S}_{gt}^{i}\}_{i=1}^{n}$. Let $\mathcal{S}^{i,l}_{gt}$ denote a subsequence of the ground truth sequence $\mathcal{S}^{i}_{gt}$ containing all events with label $l$. By definition, multiclass T-mAP is the average of the average precision (AP) values for each label $l$:
$\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$ (7)
Consider AP computation for a particular label $l$. AP is computed as the area under the precision-recall curve:
$\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$ (8)
where $\mathrm{Rec}_{i}$ is the $i$-th recall value in a sorted sequence and $\mathrm{Prec}$ is the corresponding precision. Iteration is done among all distinct recall values ($\mathrm{Rec}_{0}=0$).
We have several correct and incorrect predictions for each threshold $h$ on the predicted label probability. These quantities define precision and recall values along with the total number of ground truth events. A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|<\delta$. Note that each target can be assigned to at most one prediction. Therefore, we define a matching, i.e., the correspondence between predictions and ground truth events. The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 4.2.
As the matching, without loss of generality, can be considered constant for different thresholds, we can split all predictions into two parts: those that were assigned a ground truth and those that were unmatched. The matched set constitutes potential true or false positives, depending on the threshold. The unmatched set is always considered a false positive. Similarly, unmatched ground truth events are always considered false negatives. The resulting algorithm for AP computation for each label $l$ involves the following steps:
  1. 1.
    Compute the optimal matching between predictions and ground truth events.
  2. 2.
    Collect (a) scores of matched predictions, (b) scores of unmatched predictions, and (c) the number of unmatched ground truth events.
  3. 3.
    Assign a positive label to matched predictions and a negative label to unmatched ones.
  4. 4.
    Evaluate maximum recall as the fraction of matched ground truth events.
  5. 5.
    Find the area under the precision-recall curve for the constructed binary classification problem from item 3 and multiply it by the maximum recall value.
We have thus defined all necessary steps for T-mAP evaluation. Its complexity is $\mathcal{O}(LBN^{3})$, where $L$ is the number of classes, $B$ is the number of sequence pairs, and $N$ is the maximum length of predicted and ground truth sequences.
Appendix C HoTPP Benchmark Details
C.1 Methods
We implement popular methods from five groups: intensity-based, intensity-free, reranking, next-k, and statistical baselines.
Intensity-based methods. We implement the RMTPP[du2016rmtpp] and NHP[mei2017nhp] methods.
Intensity-free methods. As mentioned in Appendix A, this group includes intensity-free TPP and RMSE/MAE regression losses. We implement the MAE-CE approach, which uses MAE loss for time interval prediction and cross-entropy loss for labels.
Reranking approaches. We implement HYPRO[xue2022hypro] with the RMTPP backbone.
Next-K models. While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling [lim2021timeseriessurvey]. These methods predict multiple future events at once, avoiding the need for autoregressive inference. The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression. The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length. However, applying a modified autoregressive approach can potentially address this limitation.
In our work, we implement a Next-K variant of the MAE-CE model, which is quite straightforward. In particular, given K predictions and K ground truth events, we just compute MAE and cross-entropy losses between the pairs of corresponding events from two sequences. A bit more complicated is the Next-K variant of the RMTPP approach, as it violates Hawkes’s assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1. So, RMTPP-K, unlike RMTPP, can’t model dependencies between the predicted K events. Despite this, RMTPP-K is still a strong model with top-performing results on the StackOverflow dataset.
In our work, we implement a Next-K variant of the MAE-CE model, which is straightforward. Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences. The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1. Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events. Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
MostPopular baseline. This baseline analyzes the prefix sequence to identify the most frequent label and computes an average time step. The model generates constant output until the required duration (horizon) is reached. While this approach does not yield impressive results, it demonstrates moderate quality, especially in next-event prediction tasks.
Last-K baseline. This method extracts the required number of events from the end of the prefix and shifts timestamps to the current value. Despite its simplicity, this baseline demonstrates OTD quality comparable to neural approaches on small datasets.
C.2 Metrics
MTPP models are typically evaluated by the quality of the next event prediction. These metrics usually include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label quality evaluation. Some works also evaluate test set likelihood as predicted by the model. We do not measure this quantity, as it is intractable for MAE-CE and Next-K models. Previous works have made progress in evaluating long-horizon prediction [mei2019imputing, xue2023easytpp] with optimal transport distance (OTD). We evaluate these metrics, as well as the novel T-mAP metric, which fixes the flaws of previous metrics, as discussed in Section 3. We set the maximum time error $\delta$ for the T-mAP metric twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$.
C.3 Backbones
We use two types of architectures in our experiments. The first one is the GRU network[cho2014gru], which is one of the top-performing neural architectures in the event sequences domain [babaev2022coles]. We also implement a continuous-time LSTM (CT-LSTM) from the NHP method [mei2017nhp]. As CT-LSTM requires a special loss function and increases training time, we apply it only in the NHP method. In other cases, we prefer GRU. Another advantage of GRU is that its output is equal to its hidden state, which simplifies intermediate hidden state estimation for autoregressive inference starting from the middle of a sequence.
C.4 Datasets
For the first time, we combine domains of financial transactions, social networks, and medical records in a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset (AgePred) [babaev2022coles], MIMIC-IV [johnson2020mimic], social networks datasets (Retweet [zhao2015seismic] and StackOverflow [jure2014snap]), and Amazon dataset from the area of recommender systems [jure2014amazon]. These datasets have different underlying processes. For example, social network datasets are subject to cascades [zhao2015seismic]. Medical records show repetitive patterns and transactional data resembles daily activities with some regularity and large uncertainty.
The dataset statistics are presented in Table 2. Transactions have the largest average sequence length and the largest number of classes. MIMIC-IV has the largest number of sequences. Retweet is medium size and Amazon with StackOverflow can be considered small datasets.
Table 2: Datasets statistics
Dataset Sequences Events Mean Length Mean Horizon Length Mean Duration Classes Transactions 50k 43.7M 875 9.0 719 203 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 Retweet 23k 1.3M 56.4 14.7 1805 3 Amazon 9k 403K 43.6 14.8 22.1 16 StackOverflow 2k 138K 64.2 12.0 55.3 22
Transactions [^2] https://huggingface.co/datasets/dllllb/age-group-prediction, Retweet [^3] https://huggingface.co/datasets/easytpp/retweet, Amazon [^4] https://huggingface.co/datasets/easytpp/amazon, and StackOverflow [^5] https://huggingface.co/datasets/easytpp/stackoverflow datasets were obtained from the HuggingFace repository. Transactions data was released in competition and come with a free license [^6] https://www.kaggle.com/competitions/clients-age-group/data. Retweet, Amazon, and StackOverflow come with an Apache 2.0 license. MIMIC-IV is subject to PhysioNet Credentialed Health Data License 1.5.0, which requires ethical use of this dataset. Because of a complex data structure, we implement a custom preprocessing pipeline for the MIMIC-IV dataset.
Notes on MIMIC-IV preprocessing
MIMIC-IV is a publicly available electronic health record database, that includes patient diagnoses, lab measurements, procedures, and treatments. We use data preprocessing from EventStreamGPT to construct intermediate representations in EFGPT format. The following entities are obtained: subjects data frame with time-independent patient records, events data frame with lists of event types happening to subjects at specific timestamps, and measurements data frame containing specific values of time-dependent measurements and linked to events data frame.
We join events and measurements and form sequences of events for subjects. The classification labels can be generally separated into diagnoses in the form of ICD codes and types of events, such as admissions, procedures, measurements, etc. The main issue is the inability to use either diagnoses only or event types only. In the former case, the sequences become too short, while in the latter, the sequences become too repetitive, with periodic events, such as treatment start/finish, dominating the data.
However, using both diagnoses and event types simultaneously presents another problem, as diagnosis designations are distributed sparsely in the constant stream of repeat procedures. This imbalanced class distribution causes poor performance. Therefore, we further filter the data by removing duplicate events between diagnoses. This way, we can retain useful treatment data and preserve class balance.
The final labels are formed by converting ICD-9 and ICD-10 codes to ICD-10 chapters using General Equivalence Mapping (GEM) and adding event types as additional classes. The conversion is needed because the number of ICD codes is too large to use them as classes directly. We also address the issue of multiple diagnoses being designated in a single event by sorting timestamps for reproducibility.
Appendix D Training details
We trained each model with the Adam optimizer, learning rate set to $0.001$, and learning rate scheduler, decreasing learning rate by 20
We used NVIDIA V100 and A100 GPUs for computation. Each method was trained with a single GPU. The training time depends on the dataset and method. It varies from 5 minutes for RMTPP on StackOverflow to 40 minutes for the same method on Transactions or even 15 hours for the NHP method on the Transactions dataset. Multiseed evaluation takes 5 times more time to finish.
Appendix E Hyperparameters
We analyzed the behavior of OTD and T-mAP metrics depending on their parameters. The results are shown in Figure 9. OTD cost has little impact on methods ranking but largely affects the metric scale. The choice of $\delta$ parameter in T-mAP can, in some cases, affect methods ranking, but if it is not too small and too large, the metric demonstrates stable ranking. We set the $\delta$ parameter to about 10-30% of the horizon duration.
Evaluation hyperparameters are presented in Table 3. Dataset-specific training parameters are presented in Table 4.
Table 3: Evaluation hyperparameters.
Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 16 5 1 7 (week) 2 MIMIC-IV 16 5 2 28 (days) 4 Retweet 16 10 15 180 (seconds) 30 Amazon 16 5 1 10 (unk. unit) 2 StackOverflow 16 10 1 10 (minutes) 2
Table 4: Training hyperparameters.
Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens Transactions 30 1200 256 512 512 $\rightarrow$ 256 MIMIC-IV 30 64 16 64 64 Retweet 30 264 16 64 64 Amazon 60 94 32 64 64 StackOverflow 60 101 32 64 64
Appendix F Labels Entropy Degradation
We omitted some datasets in the main part of the paper for simplicity. In Fig. 7 one can see the dependency of the labels distribution entropy depending on step size for all datasets.
Figure 7: Labels distribution entropy depending on the generated sequence position.
Appendix G The Optimal Sequence Length
We omitted some datasets in the main part of the paper for simplicity. In Fig. 8 you can see the dependency of the labels distribution entropy depending on step size for all datasets.
Figure 8: T-mAP dependency on the maximum length of the predicted sequence.
Figure 9: The dependency of metric values on the metric parameter.

### v2

@@
 Abstract
-In sequential event prediction, which finds applications in finance, retail, social networks, and healthcare, a crucial task is forecasting multiple future events within a specified time horizon.
-Traditionally, this has been addressed through autoregressive generation using next-event prediction models, such as Marked Temporal Point Processes.
-However, autoregressive methods use their own output for future predictions, potentially reducing quality as the prediction horizon extends.
-In this paper, we challenge traditional approaches by introducing a novel benchmark, HoTPP, specifically designed to evaluate a model’s ability to predict event sequences over a horizon.
-This benchmark features a new metric inspired by object detection in computer vision, addressing the limitations of existing metrics in assessing models with imprecise time-step predictions.
-Our evaluations on established datasets employing various models demonstrate that high accuracy in next-event prediction does not necessarily translate to superior horizon prediction, and vice versa.
-HoTPP aims to serve as a valuable tool for developing more robust event sequence prediction methods, ultimately paving the way for further advancements in the field.
-The code is available at https://github.com/ivan-chai/hotpp-benchmark.
+Accurately forecasting multiple future events within a given time horizon is crucial for finance, retail, social networks, and healthcare applications.
+Event timing and labels are typically modeled using Marked Temporal Point Processes (MTPP), with evaluations often focused on next-event prediction quality.
+While some studies have extended evaluations to a fixed number of future events, we demonstrate that this approach leads to inaccuracies in handling false positives and false negatives.
+To address these issues, we propose a novel evaluation method inspired by object detection techniques from computer vision.
+Specifically, we introduce Temporal mean Average Precision (T-mAP), a temporal variant of mAP, which overcomes the limitations of existing long-horizon evaluation metrics.
+Our extensive experiments demonstrate that models with strong next-event prediction accuracy can yield poor long-horizon forecasts and vice versa, indicating that specialized methods are needed for each task.
+To support further research, we release HoTPP [^1] https://github.com/ivan-chai/hotpp-benchmark, the first benchmark designed explicitly for evaluating long-horizon MTPP predictions.
+HoTPP includes large-scale datasets with up to 43 million events and provides optimized procedures for both autoregressive and parallel inference, paving the way for future advancements in the field.
 1 Introduction
@@
 Event sequences differ fundamentally from other data types.
-Unlike tabular data [wang2022transtab], ESs include timestamps and possess an inherent order.
-In contrast to time series data [lim2021timeseriessurvey], ESs are characterized by irregular time intervals and structured descriptions of each event.
+Unlike tabular data (Wang & Sun, 2022), ESs include timestamps and possess an inherent order.
+In contrast to time series data (Lim & Zohren, 2021), ESs are characterized by irregular time intervals and structured descriptions of each event.
 These differences necessitate the development of specialized models and evaluation practices.
-The primary task in the domain of ESs is sequence modeling, specifically predicting the next event type and its occurrence time.
-Indeed, the ability to accurately forecast sequential events is vital for applications such as stock price prediction, personalized recommendation systems, and early disease detection.
-The simplified domain, represented as pairs of event types and times, is typically called Marked Temporal Point Processes (MTPP) [rizoiu2017hawkes-tutorial].
-Additionally, structured modeling of dependencies between different data fields [mcdermott2024eventstreamgpt] can be considered an extension of MTPP.
+Sequence modeling is the primary task in Event Sequences (ESs) domain.
+In its simplest form, each event is defined by its type and occurrence time, a framework commonly known as Marked Temporal Point Processes (MTPP) (Rizoiu et al., 2017).
+Some studies extend MTPP to incorporate additional data fields and model complex dependencies between them (McDermott et al., 2024).
+However, the majority of MTPP approaches, as well as their evaluation pipelines, primarily focus on predicting the next event based on historical data.
 In practice, a common question arises: what events will occur, and when, within a specific time horizon?
-Forecasting multiple future events poses unique challenges distinct from traditional next-event prediction tasks.
-The conventional approach to this problem typically involves the autoregressive application of next-event prediction models.
-While these models have demonstrated effectiveness in predicting the immediate next event, their performance tends to decline as the prediction horizon extends.
-This results in increased computational costs and reduced predictive accuracy over longer sequences.
-Despite this, it is commonly assumed that the quality of next-event prediction directly influences the quality of the corresponding autoregressive model.
-In this paper, we challenge this belief and demonstrate that autoregressive models can be suboptimal for long-horizon predictions in some cases.
-To address this, we introduce HoTPP (Horizon Temporal Point Process), a novel benchmark designed to evaluate the effectiveness of event sequence prediction models over horizons.
-Our benchmark leverages a new metric, T-mAP (Temporal mean Average Precision), inspired by object detection methodologies in computer vision.
-We demonstrate that T-mAP offers a more nuanced assessment of a model’s performance in scenarios where precise time step predictions are challenging.
-This innovation addresses the inherent limitations of traditional metrics, which often fail to capture the complexities involved in long-term event forecasting.
-Thus, our main contributions can be summarized as follows:
+Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks.
+For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions.
+However, the challenges of autoregressive prediction in the context of MTPP have not been thoroughly explored.
+Another difficulty lies in evaluation.
+Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su & Hua, 2017).
+Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths.
+Still, the limitations of this metric have not been previously considered.
+In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies.
+Our key contributions are as follows:
   1. 1.
-    We introduce HoTPP, a new open-source benchmark designed to facilitate research on long horizon Event Sequences (ESs) prediction.
-This benchmark integrates datasets and methods from various domains, including financial transactions, social networks, medicine, and recommender systems.
+    We demonstrate that widely used evaluation methods for MTPPs often overlook critical aspects of model performance.
+We demonstrate that simple rule-based baselines can sometimes outperform popular deep learning methods when evaluated using OTD.
   2. 2.
-    We demonstrate that widely used evaluation methods for ESs and MTPPs often overlook critical aspects of model performance.
-Notably, we show that simple statistical baselines can sometimes outperform popular methods in terms of Optimal Transport Distance (OTD).
-To address the shortcomings of OTD, we introduce a new evaluation metric for MTPP called Temporal mAP (T-mAP) (Fig.
-1).
+    We introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision.
+T-mAP evaluates variable length sequences within a specified time horizon, as illustrated in Figure 1.
+Unlike previous approaches, T-mAP accurately accounts for false positives and false negatives while being invariant to linear calibration.
+Additionally, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
   3. 3.
-    We show that the length of the predicted sequence significantly impacts long-horizon prediction performance and must be carefully tuned. We also provide a method for selecting this hyperparameter.
+    Using our established methodology, we demonstrate that high next-event prediction accuracy does not necessarily translate into high-quality long-horizon forecasts; in many cases, our experiments show the opposite.
+This highlights the necessity of developing specialized models for the long-horizon prediction task.
   4. 4.
-    We compare popular approaches from different domains and show that intensity-based MTPP modeling methods can be suboptimal compared to simple approaches.
-We also show that autoregressive approaches can be outperformed by direct prediction of multiple future events in some cases.
+    We release HoTPP, a new open-source benchmark designed to facilitate long-horizon event sequence prediction research.
+HoTPP brings together datasets and methods from various domains, including financial transactions, social networks, healthcare, and recommender systems, greatly expanding the diversity and scale of data compared to prior benchmarks.
+Additionally, we offer an efficient inference algorithm necessary for large-scale evaluations.
 Figure 1: Temporal mAP (T-mAP) evaluation pipeline.
-Unlike previous methods, T-mAP evaluates sequences of variable lengths over the prediction horizon.
-It also enhances performance assessment by analyzing label distributions rather than fixed predictions.
-2 Background
-2.1 Marked Temporal Point Processes
-Marked Temporal Point Process (MTPP) is a stochastic process that consists of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ are event type labels.
-Common MTPP modeling approaches primarily focus on predicting the next event.
-A straightforward solution is to predict time and type independently.
-A more advanced approach involves splitting the original sequence into subsequences, one for each event type, and then independently modeling each subsequence’s time.
-In both approaches, the fundamental problem is modeling the arrival time of the next event.
-The stochastic process that generates these times is typically called a Temporal Point Process [rizoiu2017hawkes-tutorial].
-For the details on MTPP modeling, refer to Appendix A.
-2.2 Evaluation
-MTPPs are typically evaluated based on the quality of next-event predictions.
-Time and type predictions are usually assessed independently.
-The quality of type prediction is characterized by the error rate, while time prediction error is measured using either the Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE).
-Recent studies have advanced the measurement of horizon prediction quality by employing the Optimal Transport Distance (OTD).
-The OTD metric is analogous to the edit distance between the predicted event sequence and the ground truth.
-Below, we provide the formal definition of OTD.
-Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$.
-These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$.
-The prediction is connected to the ground truth event if their types are equal.
-If $M(S_{p},S_{gt})$ is a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$, OTD is computed as follows:
-$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
-where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost.
-It is common to take $C_{ins}=C_{del}$ [mei2019imputing].
-It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
-However, OTD has practical limitations, which are described in the following section.
+Unlike previous methods, T-mAP evaluates sequences of variable lengths within the prediction horizon.
+It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
+2 Related Work
+Marked Temporal Point Processes.
+A Marked Temporal Point Process (MTPP) is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels (Rizoiu et al., 2017).
+Common approaches to MTPP modeling focus on predicting the next event.
+A basic solution involves independently predicting the event time and type.
+A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing.
+Depending on the time-step distribution, the process is called Poisson or Hawkes.
+Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures.
+Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei & Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024).
+In this work, we evaluate MTPP models with both discrete and continuous time architectures.
+Unlike previous benchmarks, we also assess simple rule-based predictors and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021) and Next-K models from time series analysis (Lim & Zohren, 2021).
+For more details on MTPP modeling, refer to Appendix A.
+MTPP Evaluation.
+Previous benchmarks have primarily focused on medical data or traditional MTPP datasets.
+EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev & van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability.
+Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks.
+While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and Next-K approaches.
+Furthermore, PoPPy does not evaluate long-horizon predictions at all.
+EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
+Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su & Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided.
+This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW).
+In our study, we address the limitations of OTD, a variant of the Wasserstein Distance, and propose a new metric called T-mAP.
+Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
 3 Limitations of the Next-Event and OTD Metrics
 Figure 2: Comparison of OTD and T-mAP metrics.
-In example 1.a, prefix evaluation of three events using OTD results in an incorrect false positive.
+In example 1.a prefix evaluation of three events using OTD results in an incorrect, false negative.
 T-mAP addresses this issue by comparing all events within the time horizon, as shown in example 1.b.
@@
 In contrast, T-mAP evaluates the entire distribution, yielding a low score of 0.33 in case 2.a and a high score of 1 in case 2.b.
-A predominant evaluation metric in the MTPP domain is the quality of next-event prediction.
-This metric typically comprises two distinct components: the accuracy (or error rate) of the next label prediction and either Mean Absolute Error (MAE) or Root Mean Square Error (RMSE) for time step prediction.
-However, these metrics do not assess the model’s ability to predict multiple future events.
-For instance, autoregression uses its own outputs as inputs for the next step, which can potentially lead to cumulative prediction errors.
-Consequently, long-horizon evaluation metrics are necessary.
-The OTD metric extends evaluation to sequences by comparing fixed-size prefixes, but this approach limits its flexibility in assessing models with imprecise time step predictions, as illustrated in Fig.
+MTPPs are typically evaluated based on the accuracy of next-event predictions, with time and type predictions assessed independently.
+The quality of type predictions is measured by the error rate, while time prediction error is evaluated using either Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE).
+However, these metrics do not account for the model’s ability to predict multiple future events.
+For example, in autoregressive models, the outputs are fed back as inputs for subsequent steps, which can lead to cumulative prediction errors.
+As a result, long-horizon evaluation metrics are necessary.
+Recent studies have advanced the measurement of horizon prediction quality by employing Optimal Transport Distance (OTD) (Mei et al., 2019).
+This metric is analogous to the edit distance between the predicted event sequence and the ground truth.
+Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$.
+These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$.
+The prediction is connected to the ground truth event if their types are equal.
+Denote $M(S_{p},S_{gt})$ a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$.
+OTD finds the minimum cost among all possible matchings:
+$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
+where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost.
+It is common to take $C_{ins}=C_{del}$ (Mei et al., 2019).
+It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
+OTD is computed between fixed-size prefixes, but this approach limits metric flexibility in assessing models with imprecise time step predictions, as illustrated in Fig.
 2.1.
@@
 This is undesirable as it prevents the inclusion of additional ground truth events in the evaluation.
-Therefore, evaluating a dynamic number of events is essential to properly align with the time horizon in the ground truth.
-Another limitation of the OTD metric is its inflexibility in evaluating predicted distributions of labels, as shown in Fig.
-2.2.
-OTD evaluates only the labels with the highest probability, disregarding the full distribution.
-As a result, it relies on the calibration of model outputs and cannot assess the performance beyond a limited set of popular labels.
-However, models typically predict probabilities for all classes, allowing for an independent assessment of prediction quality for each label.
-Therefore, we aim to develop an evaluation metric that accurately reflects the performance across common and rare classes.
+Therefore, evaluating a dynamic number of events is essential to properly align with the ground truth’s time horizon.
+MTPP evaluation metrics can be categorized into two groups: those considering event indices and those assessing time and label prediction quality independently of actual positions.
+Next-event metrics rely on ordering, even when events share identical timestamps, making the order meaningless.
+While Optimal Transport Distance (OTD) itself is invariant to order, the extraction of sequences for comparison is influenced by event indices.
+For example, when OTD compares prefixes of length $K$, the algorithm must place $K$ correctly predicted events at the beginning of a sequence to minimize the OTD score.
+As a result, OTD, like next-event metrics, can depend on event ordering even when the order cannot be uniquely determined.
+This dependency is non-trivial and difficult to measure or control accurately.
+Another limitation of the OTD metric is its inability to evaluate the full predicted distribution of labels, as shown in Figure 2.2.
+OTD considers only the labels with the highest probability, ignoring the complete distribution.
+This makes it dependent on model calibration and limits the ability to assess performance across a broader range of event types, such as long-tail predictions.
+However, models typically predict probabilities for all classes, allowing for a more comprehensive assessment.
+Therefore, we aim to develop an evaluation metric that accurately captures performance across common and rare classes.
 4 Temporal mAP: Prediction as Detection
-In this section, we introduce a novel metric, Temporal mAP (T-mAP), inspired by object detection methods from computer vision [everingham2010pascalcvdetection], as illustrated in Fig.
+In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration.
+T-mAP is inspired by object detection methods from computer vision (Everingham et al., 2010), as illustrated in Fig.
 1.
 The goal of object detection is to localize objects within an image and identify their types.
-In event sequences, we tackle a similar problem, but we consider the time dimension instead of horizontal and vertical axes.
+In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes.
 Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration.
@@
 $S^{l}_{>}(h)=\{t^{p}_{i}:\exists(t^{p}_{i},s^{p}_{i})\in S^{l}_{p},s^{p}_{i}>h\}.$ (2)
-A prediction $i$ can be matched with ground truth $j$ iff $|t^{p}_{i}-t^{gt}_{j}|<\delta$.
-T-mAP identifies the matching that maximizes precision and recall, matching with maximum cover $c$.
+By definition, a predicted event $i$ can be matched with a ground truth event $j$ iff $|t^{p}_{i}-t^{gt}_{j}|\leq\delta$, meaning the time difference between the predicted and ground truth events is less than or equal to $\delta$.
+T-mAP identifies the matching that maximizes precision and recall, i.e., matching with maximum cover $c$.
 The precision of this matching is calculated as $c/|S^{l}_{>}(h)|$ and the recall as $c/|S^{l}_{gt}|$.
@@
 Note that there are no true negatives, as the model cannot explicitly predict the absence of an event.
-Similarly to binary classification, we can define a precision-recall curve by varying the threshold $h$ and estimating Average Precision (AP), the area under the precision-recall curve.
-T-mAP is the mean AP across all possible event types $l$.
+Similar to binary classification, we can define a precision-recall curve by varying the threshold $h$ and then estimate the Average Precision (AP), the area under the precision-recall curve.
+Finally, T-mAP is defined as the average AP across all classes.
 4.2 Computation
 Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP.
-In this section, we show how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm [jonker1988linsum].
+This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker & Volgenant, 1988).
 The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
@@
 For each pair of sequences $S^{l}_{p}$ and $S^{l}_{gt}$ we define a weighted bipartite graph $\mathcal{G}(S^{l}_{p},S^{l}_{gt})$ with $|S^{l}_{p}|$ vertices in the first and $|S^{l}_{gt}|$ vertices in the second part.
-For each pair of prediction $i$ and ground truth event $j$ with $|t^{p}_{i}-t^{gt}_{j}|<\delta$ we add an edge with weight $-s^{p}_{i}$, as shown in Fig.
+For each pair of prediction $i$ and ground truth event $j$ with $|t^{p}_{i}-t^{gt}_{j}|\leq\delta$ we add an edge with weight $-s^{p}_{i}$, equal to negative logit of the target class, as shown in Fig.
 3.a.
@@
 3.c.
-The proof of the theorem and the complete algorithm for T-mAP evaluation are provided in Appendix B.
+The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in Appendix B.
+4.3 T-mAP Hyperparameters
+T-mAP has two hyperparameters: the maximum allowed time delta $\delta$ and evaluation horizon $H$.
+We set $\delta$ twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$.
+The horizon $H$ must be larger than $\delta$ to evaluate timestamp quality adequately.
+Therefore, depending on the dataset, we select $H$ to be approximately 3-7 times larger than $\delta$, ensuring the horizon captures an average of 6-15 events.
+The empirical study of T-mAP hyperparameters is presented in Figure 4.
+As shown, the chosen $\delta$ parameter is positioned to the right of the initial slope of the parameter-quality curve, ensuring an optimal balance between time granularity and task difficulty.
+Figure 4: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value. Results for the StackOverflow and Amazon datasets are provided in the Appendix E.
 5 HoTPP Benchmark
@@
 Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction.
-HoTPP differs from prior MTPP benchmarks by including simple statistical baselines and next-k models, which simultaneously predict multiple future events.
-The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, and reproducibility.
+HoTPP differs from prior MTPP benchmarks by including simple rule-based baselines and next-k models, simultaneously predicting multiple future events.
+The HoTPP benchmark is designed with a focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
 Simplicity and Extensibility.
-The benchmark code is clearly structured, separating the library, dataset-specific scripts, and configuration files.
-New methods can be implemented at multiple levels: configuration file, model architecture, loss function, metric, and training module.
-Core components can be reused by applying Hydra configuration library [Yadan2019Hydra].
+The benchmark code is organized into a clear structure, separating the core library, dataset-specific scripts, and configuration files.
+New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module.
+Core components are reusable through the Hydra configuration library (Yadan, 2019).
+For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, and ODE.
+Another example is our Next-K implementation, which can be applied to IFTTP or RMTPP.
 Evaluation Stability.
-Unlike previous works [xue2023easytpp, xue2022hypro], we evaluate methods not only at the end of each sequence but also at intermediate points.
-This approach collects a larger set of predictions, reducing the final metric’s STD.
-This is particularly important for datasets with few sequences, such as StackOverflow and Amazon.
+In the MTPP domain, many datasets contain only a few thousand sequences for training and evaluation.
+For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing.
+Previous long-horizon evaluation pipelines (Xue et al., 2022; 2023) made predictions only at the end of each sequence, resulting in a limited number of predictions and reduced evaluation stability.
+To address this limitation, we evaluate long-horizon predictions at multiple intermediate points.
 Reproducibility.
 The HoTPP benchmark ensures reproducibility at multiple levels.
-First, we use the PytorchLightning library [Falcon_PyTorch_Lightning_2019] for training with a specified random seed, and report multi-seed evaluation results.
+First, we use the PytorchLightning library (Falcon & The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results.
 Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly.
 Finally, we specify the environment in a Dockerfile.
-A detailed benchmark description can be found in Appendix C.
-Table 1: Evaluation results. The best result is shown in bold. Mean and STD values of 5 runs with different random seeds are reported.
-Method Mean length Next-item Acc., % Next-item MAE OTD Val / Test T-mAP, % Val / Test Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.09$\pm$ 0.00 / 1.02$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.924$\pm$ 0.000 7.51$\pm$ 0.00 / 7.55$\pm$ 0.00 1.64$\pm$ 0.00 / 1.69$\pm$ 0.00 MAE-CE 14.0 38.18$\pm$ 0.03 0.570$\pm$ 0.001 7.18$\pm$ 0.01 / 7.20$\pm$ 0.01 4.63$\pm$ 0.33 / 4.50$\pm$ 0.27 MAE-CE-K 9.3 37.68$\pm$ 0.01 0.573$\pm$ 0.000 7.10$\pm$ 0.02 / 7.10$\pm$ 0.02 3.34$\pm$ 0.16 / 3.73$\pm$ 0.16 RMTPP 11.5 38.29$\pm$ 0.02 0.658$\pm$ 0.001 6.78$\pm$ 0.01 / 6.80$\pm$ 0.01 6.72$\pm$ 0.35 / 6.49$\pm$ 0.17 RMTPP-K 9.2 37.74$\pm$ 0.03 0.664$\pm$ 0.001 7.11$\pm$ 0.00 / 7.12$\pm$ 0.00 4.79$\pm$ 0.23 / 4.96$\pm$ 0.13 NHP 4.7 30.49$\pm$ x.xx 0.995$\pm$ x.xxx 7.68$\pm$ x.xx / 7.69$\pm$ x.xx 1.62$\pm$ x.xx / 1.51$\pm$ x.xx HYPRO 11.4 38.27$\pm$ x.xx 0.658$\pm$ x.xxx 6.81$\pm$ x.xx / 6.82$\pm$ x.xx 6.72$\pm$ x.xx / 6.51$\pm$ x.xx MIMIC-IV MostPopular 8.3 1.79$\pm$ 0.00 11.23$\pm$ 0.00 19.89$\pm$ 0.00 / 19.88$\pm$ 0.00 0.73$\pm$ 0.00 / 0.75$\pm$ 0.00 Last 5 4.2 0.88$\pm$ 0.00 4.77$\pm$ 0.00 19.77$\pm$ 0.00 / 19.74$\pm$ 0.00 2.31$\pm$ 0.00 / 2.38$\pm$ 0.00 MAE-CE 10.8 57.38$\pm$ 0.05 2.85$\pm$ 0.01 11.69$\pm$ 0.00 / 11.70$\pm$ 0.01 23.48$\pm$ 0.20 / 23.60$\pm$ 0.20 MAE-CE-K 15.3 56.71$\pm$ 0.22 2.89$\pm$ 0.02 14.04$\pm$ 0.24 / 14.01$\pm$ 0.25 21.26$\pm$ 0.16 / 21.03$\pm$ 0.11 RMTPP 15.9 57.35$\pm$ 0.05 2.85$\pm$ 0.01 12.93$\pm$ 0.03 / 12.93$\pm$ 0.03 20.61$\pm$ 0.15 / 20.85$\pm$ 0.23 RMTPP-K 16.0 56.69$\pm$ 0.03 2.79$\pm$ 0.00 13.91$\pm$ 0.04 / 13.88$\pm$ 0.03 23.01$\pm$ 0.04 / 22.85$\pm$ 0.06 NHP 8.5 24.71$\pm$ 0.95 6.72$\pm$ 0.35 18.73$\pm$ 0.19 / 18.69$\pm$ 0.20 8.55$\pm$ 1.76 / 8.57$\pm$ 1.75 HYPRO 16.0 57.18$\pm$ x.xx 2.83$\pm$ x.xx 12.93$\pm$ x.xx / 12.94$\pm$ x.xx 21.47$\pm$ x.xx / 21.50$\pm$ x.xx Retweet MostPopular 15.5 58.50$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 26.51$\pm$ 0.00 / 25.87$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 21.93$\pm$ 0.00 165.5$\pm$ 0.0 / 164.0$\pm$ 0.0 28.54$\pm$ 0.00 / 28.61$\pm$ 0.00 MAE-CE 15.0 59.92$\pm$ 0.04 18.28$\pm$ 0.03 173.6$\pm$ 4.3 / 173.2$\pm$ 4.6 33.10$\pm$ 4.84 / 31.35$\pm$ 4.71 MAE-CE-K 14.8 59.30$\pm$ 0.33 18.25$\pm$ 0.45 169.7$\pm$ 2.2 / 169.3$\pm$ 2.6 36.34$\pm$ 5.06 / 34.09$\pm$ 5.46 RMTPP 12.9 60.06$\pm$ 0.04 18.54$\pm$ 0.09 169.1$\pm$ 2.2 / 167.9$\pm$ 2.3 50.98$\pm$ 0.88 / 48.35$\pm$ 1.01 RMTPP-K 13.6 60.03$\pm$ 0.05 18.42$\pm$ 0.21 166.4$\pm$ 4.7 / 165.8$\pm$ 4.7 49.21$\pm$ 1.06 / 46.31$\pm$ 1.25 NHP 10.7 59.84$\pm$ 0.16 27.37$\pm$ 6.91 194.4$\pm$ 14.6 / 192.1$\pm$ 13.8 39.71$\pm$ 9.02 / 37.52$\pm$ 8.56 HYPRO 13.0 60.12$\pm$ x.xx 18.55$\pm$ x.xx 171.6$\pm$ x.x / 170.6$\pm$ x.x 51.79$\pm$ x.xx / 49.60$\pm$ x.xx Amazon MostPopular 15.7 33.46$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 9.34$\pm$ 0.00 / 9.02$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 0.323$\pm$ 0.000 6.70$\pm$ 0.00 / 6.67$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 MAE-CE 13.2 35.73$\pm$ 0.12 0.242$\pm$ 0.003 6.62$\pm$ 0.06 / 6.55$\pm$ 0.04 21.98$\pm$ 0.36 / 22.58$\pm$ 0.37 MAE-CE-K 13.3 35.10$\pm$ 0.08 0.246$\pm$ 0.000 6.74$\pm$ 0.01 / 6.70$\pm$ 0.01 21.84$\pm$ 0.05 / 22.35$\pm$ 0.07 RMTPP 15.3 35.74$\pm$ 0.04 0.294$\pm$ 0.001 6.68$\pm$ 0.02 / 6.62$\pm$ 0.03 19.92$\pm$ 0.26 / 20.33$\pm$ 0.33 RMTPP-K 15.8 35.10$\pm$ 0.13 0.300$\pm$ 0.001 6.95$\pm$ 0.01 / 6.89$\pm$ 0.01 17.58$\pm$ 0.12 / 17.83$\pm$ 0.13 NHP 14.4 22.38$\pm$ 6.15 0.378$\pm$ 0.073 7.68$\pm$ 0.12 / 7.67$\pm$ 0.14 21.85$\pm$ 0.86 / 21.34$\pm$ 0.86 HYPRO 14.9 35.87$\pm$ x.xx 0.293$\pm$ x.xxx 6.69$\pm$ x.xx / 6.64$\pm$ x.xx 20.40$\pm$ x.xx / 20.96$\pm$ x.xx StackOverflow MostPopular 11.8 42.90$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.33$\pm$ 0.00 / 5.86$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 0.950$\pm$ 0.000 14.87$\pm$ 0.00 / 14.92$\pm$ 0.00 8.61$\pm$ 0.00 / 6.64$\pm$ 0.00 MAE-CE 15.7 45.39$\pm$ 0.11 0.641$\pm$ 0.002 13.58$\pm$ 0.08 / 13.65$\pm$ 0.05 8.82$\pm$ 0.70 / 8.45$\pm$ 0.38 MAE-CE-K 15.3 44.84$\pm$ 0.24 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.52$\pm$ 0.06 12.15$\pm$ 0.99 / 11.18$\pm$ 0.79 RMTPP 12.5 45.44$\pm$ 0.13 0.700$\pm$ 0.007 12.97$\pm$ 0.03 / 13.18$\pm$ 0.05 13.44$\pm$ 0.23 / 12.95$\pm$ 0.11 RMTPP-K 12.0 44.84$\pm$ 0.08 0.687$\pm$ 0.001 12.93$\pm$ 0.01 / 13.13$\pm$ 0.02 14.79$\pm$ 0.45 / 14.19$\pm$ 0.26 NHP 10.4 44.44$\pm$ 0.06 0.758$\pm$ 0.021 13.32$\pm$ 0.16 / 13.59$\pm$ 0.16 12.05$\pm$ 0.64 / 11.35$\pm$ 0.42 HYPRO 11.8 45.42$\pm$ x.xx 0.716$\pm$ x.xxx 13.05$\pm$ x.xx / 13.25$\pm$ x.xx 13.51$\pm$ x.xx / 13.61$\pm$ x.xx
+Computational Efficiency.
+Some methods are particularly slow, especially during autoregressive inference on large datasets.
+Straightforward evaluation can take several hours on a single Nvidia V100 GPU.
+To accelerate computation, we optimized the training and inference pipelines in two ways.
+First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions.
+Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE), achieving up to a 4x performance improvement compared to the official implementations.
+These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV.
+Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
+A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s computational performance in Appendix C.5.
+Table 1: Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
+Method Mean length Next-item Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.86$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.06$\pm$ 0.00 / 0.99$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.87$\pm$ 0.00 0.924$\pm$ 0.000 7.40$\pm$ 0.00 / 7.44$\pm$ 0.00 2.46$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 11.0 34.08$\pm$ 0.04 3.47$\pm$ 0.01 0.693$\pm$ 0.000 6.88$\pm$ 0.01 / 6.90$\pm$ 0.01 5.82$\pm$ 0.27 / 5.88$\pm$ 0.13 IFTPP-K 10.1 33.69$\pm$ 0.03 3.25$\pm$ 0.01 0.698$\pm$ 0.001 7.18$\pm$ 0.00 / 7.19$\pm$ 0.00 4.42$\pm$ 0.15 / 4.43$\pm$ 0.16 RMTPP 7.5 34.15$\pm$ 0.07 3.47$\pm$ 0.02 0.749$\pm$ 0.005 6.86$\pm$ 0.01 / 6.88$\pm$ 0.01 7.08$\pm$ 0.16 / 6.69$\pm$ 0.12 RMTPP-K 7.3 33.63$\pm$ 0.07 3.24$\pm$ 0.02 0.749$\pm$ 0.001 7.10$\pm$ 0.00 / 7.11$\pm$ 0.00 5.82$\pm$ 0.05 / 5.52$\pm$ 0.13 NHP 9.4 35.43$\pm$ 0.07 3.41$\pm$ 0.01 0.696$\pm$ 0.002 6.97$\pm$ 0.01 / 6.98$\pm$ 0.01 5.59$\pm$ 0.07 / 5.61$\pm$ 0.05 AttNHP 7.6 31.12$\pm$ x.xx 1.21$\pm$ x.xx 0.717$\pm$ x.xxx 7.52$\pm$ x.xx / 7.50$\pm$ x.xx 1.71$\pm$ x.xx / 1.48$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 6.9 34.26$\pm$ x.xx 3.46$\pm$ x.xx 0.758$\pm$ x.xxx 7.04$\pm$ x.xx / 7.05$\pm$ x.xx 7.79$\pm$ x.xx / 7.05$\pm$ x.xx MIMIC-IV MostPopular 10.3 4.77$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.82$\pm$ 0.00 / 19.82$\pm$ 0.00 0.55$\pm$ 0.00 / 0.54$\pm$ 0.00 Last 5 4.0 1.02$\pm$ 0.00 2.63$\pm$ 0.00 5.34$\pm$ 0.00 19.75$\pm$ 0.00 / 19.73$\pm$ 0.00 2.41$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 12.7 58.59$\pm$ 0.02 47.32$\pm$ 1.39 3.00$\pm$ 0.01 11.51$\pm$ 0.01 / 11.53$\pm$ 0.01 21.93$\pm$ 0.21 / 21.67$\pm$ 0.21 IFTPP-K 15.1 57.91$\pm$ 0.13 44.60$\pm$ 0.14 3.07$\pm$ 0.02 13.17$\pm$ 0.05 / 13.18$\pm$ 0.05 22.46$\pm$ 0.03 / 22.30$\pm$ 0.03 RMTPP 10.0 58.33$\pm$ 0.02 46.24$\pm$ 1.18 3.89$\pm$ 0.03 13.64$\pm$ 0.03 / 13.71$\pm$ 0.03 21.49$\pm$ 0.31 / 21.08$\pm$ 0.29 RMTPP-K 8.4 57.48$\pm$ 0.06 43.47$\pm$ 0.11 3.62$\pm$ 0.02 14.68$\pm$ 0.01 / 14.72$\pm$ 0.02 20.70$\pm$ 0.16 / 20.39$\pm$ 0.14 NHP 6.1 24.97$\pm$ 0.94 11.12$\pm$ 0.49 6.53$\pm$ 0.77 18.59$\pm$ 0.18 / 18.60$\pm$ 0.19 7.26$\pm$ 1.35 / 7.32$\pm$ 1.33 AttNHP 5.1 42.75$\pm$ 0.52 28.55$\pm$ 1.06 3.08$\pm$ 0.04 14.66$\pm$ 0.08 / 14.68$\pm$ 0.08 22.64$\pm$ 0.41 / 22.46$\pm$ 0.40 ODE 12.5 43.21$\pm$ 2.30 25.34$\pm$ 3.05 2.93$\pm$ 0.03 14.71$\pm$ 0.34 / 14.74$\pm$ 0.34 15.41$\pm$ 0.21 / 15.18$\pm$ 0.15 HYPRO 7.4 58.35$\pm$ x.xx 45.45$\pm$ x.xx 3.95$\pm$ x.xx 14.82$\pm$ x.xx / 14.87$\pm$ x.xx 16.94$\pm$ x.xx / 16.77$\pm$ x.xx Retweet MostPopular 28.0 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 25.15$\pm$ 0.00 / 23.91$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 35.73$\pm$ 0.00 21.87$\pm$ 0.00 152.3$\pm$ 0.0 / 150.3$\pm$ 0.0 29.12$\pm$ 0.00 / 29.24$\pm$ 0.00 IFTPP 26.1 59.95$\pm$ 0.03 46.53$\pm$ 0.04 18.27$\pm$ 0.03 173.3$\pm$ 4.1 / 172.7$\pm$ 4.4 34.90$\pm$ 4.63 / 31.75$\pm$ 4.44 IFTPP-K 14.8 59.55$\pm$ 0.26 45.09$\pm$ 0.62 18.21$\pm$ 0.42 168.6$\pm$ 2.4 / 167.9$\pm$ 2.8 37.11$\pm$ 4.91 / 34.73$\pm$ 5.11 RMTPP 19.0 60.07$\pm$ 0.10 46.81$\pm$ 0.06 18.45$\pm$ 0.16 167.6$\pm$ 3.4 / 166.7$\pm$ 3.3 47.86$\pm$ 0.91 / 44.74$\pm$ 0.94 RMTPP-K 13.7 59.99$\pm$ 0.05 46.34$\pm$ 0.09 18.33$\pm$ 0.13 164.7$\pm$ 2.8 / 163.9$\pm$ 2.8 49.07$\pm$ 1.16 / 46.16$\pm$ 1.32 NHP 18.1 60.08$\pm$ 0.04 46.83$\pm$ 0.11 18.42$\pm$ 0.11 167.0$\pm$ 1.6 / 165.8$\pm$ 1.6 48.31$\pm$ 0.67 / 45.07$\pm$ 0.34 AttNHP 29.5 60.03$\pm$ 0.03 46.74$\pm$ 0.01 18.39$\pm$ 0.05 173.3$\pm$ 1.0 / 171.6$\pm$ 1.0 28.32$\pm$ 1.24 / 25.85$\pm$ 1.08 ODE 19.0 59.95$\pm$ 0.08 46.65$\pm$ 0.04 18.38$\pm$ 0.04 166.5$\pm$ 0.5 / 165.3$\pm$ 0.5 48.70$\pm$ 0.93 / 44.81$\pm$ 0.69 HYPRO 15.0 59.87$\pm$ x.xx 46.69$\pm$ x.xx 18.75$\pm$ x.xx 171.4$\pm$ x.x / 170.7$\pm$ x.x 49.90$\pm$ x.xx / 46.99$\pm$ x.xx Amazon MostPopular 20.6 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 8.59$\pm$ 0.00 / 8.31$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 8.15$\pm$ 0.00 0.321$\pm$ 0.000 6.43$\pm$ 0.00 / 6.41$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 IFTPP 13.7 35.73$\pm$ 0.12 17.14$\pm$ 0.10 0.242$\pm$ 0.003 6.58$\pm$ 0.06 / 6.52$\pm$ 0.05 21.94$\pm$ 0.46 / 22.56$\pm$ 0.52 IFTPP-K 13.3 35.11$\pm$ 0.08 16.48$\pm$ 0.03 0.246$\pm$ 0.000 6.72$\pm$ 0.01 / 6.68$\pm$ 0.01 22.06$\pm$ 0.05 / 22.57$\pm$ 0.07 RMTPP 16.7 35.76$\pm$ 0.06 17.21$\pm$ 0.02 0.294$\pm$ 0.001 6.62$\pm$ 0.02 / 6.57$\pm$ 0.03 19.70$\pm$ 0.31 / 20.06$\pm$ 0.33 RMTPP-K 15.8 35.06$\pm$ 0.14 16.37$\pm$ 0.20 0.300$\pm$ 0.005 6.92$\pm$ 0.01 / 6.87$\pm$ 0.02 17.85$\pm$ 0.29 / 18.12$\pm$ 0.30 NHP 9.9 11.06$\pm$ 3.43 11.22$\pm$ 0.19 0.449$\pm$ 0.014 9.04$\pm$ 0.31 / 9.02$\pm$ 0.35 26.24$\pm$ 0.36 / 26.29$\pm$ 0.55 AttNHP 18.5 31.83$\pm$ 0.32 9.70$\pm$ 0.28 0.461$\pm$ 0.003 7.32$\pm$ 0.06 / 7.30$\pm$ 0.06 14.50$\pm$ 0.58 / 14.62$\pm$ 0.80 ODE 9.6 7.54$\pm$ 0.95 10.14$\pm$ 0.23 0.492$\pm$ 0.018 9.48$\pm$ 0.11 / 9.46$\pm$ 0.08 23.54$\pm$ 0.62 / 22.96$\pm$ 0.61 HYPRO 18.0 35.69$\pm$ x.xx 17.21$\pm$ x.xx 0.295$\pm$ x.xxx 6.63$\pm$ x.xx / 6.61$\pm$ x.xx 20.58$\pm$ x.xx / 20.53$\pm$ x.xx StackOverflow MostPopular 14.0 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.10$\pm$ 0.00 / 5.56$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 5.20$\pm$ 0.00 0.934$\pm$ 0.000 14.52$\pm$ 0.00 / 14.55$\pm$ 0.00 8.67$\pm$ 0.00 / 6.72$\pm$ 0.00 IFTPP 24.1 45.41$\pm$ 0.11 13.00$\pm$ 0.79 0.641$\pm$ 0.002 13.57$\pm$ 0.07 / 13.64$\pm$ 0.05 8.78$\pm$ 0.87 / 8.31$\pm$ 0.50 IFTPP-K 15.3 44.85$\pm$ 0.24 11.16$\pm$ 1.07 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.51$\pm$ 0.06 12.42$\pm$ 0.97 / 11.42$\pm$ 0.78 RMTPP 14.5 45.43$\pm$ 0.13 13.33$\pm$ 0.25 0.701$\pm$ 0.007 12.95$\pm$ 0.02 / 13.17$\pm$ 0.05 13.26$\pm$ 0.29 / 12.72$\pm$ 0.16 RMTPP-K 12.0 44.89$\pm$ 0.09 11.72$\pm$ 0.10 0.689$\pm$ 0.001 12.92$\pm$ 0.01 / 13.13$\pm$ 0.01 14.91$\pm$ 0.19 / 14.30$\pm$ 0.09 NHP 13.0 44.53$\pm$ 0.05 10.86$\pm$ 0.19 0.715$\pm$ 0.004 13.02$\pm$ 0.02 / 13.24$\pm$ 0.02 12.67$\pm$ 0.55 / 11.96$\pm$ 0.40 AttNHP 15.5 45.17$\pm$ 0.10 12.67$\pm$ 0.13 0.705$\pm$ 0.001 13.08$\pm$ 0.03 / 13.30$\pm$ 0.02 11.95$\pm$ 0.19 / 11.13$\pm$ 0.32 ODE 13.9 44.38$\pm$ 0.09 10.12$\pm$ 0.19 0.711$\pm$ 0.004 13.04$\pm$ 0.02 / 13.27$\pm$ 0.03 11.37$\pm$ 0.32 / 10.52$\pm$ 0.23 HYPRO 11.8 45.18$\pm$ x.xx 12.88$\pm$ x.xx 0.715$\pm$ x.xxx 13.04$\pm$ x.xx / 13.26$\pm$ x.xx 15.57$\pm$ x.xx / 14.69$\pm$ x.xx
 6 Experiments
-We conduct experiments on five datasets of varying sizes and origins: Transaction financial dataset [agepred], MIMIC-IV medical dataset [johnson2020mimic], two social networks datasets, Retweet [zhao2015seismic] and StackOverflow [jure2014snap], and Amazon reviews dataset[jure2014amazon].
-A detailed description of these datasets, along with the training and evaluation parameters, is provided in Appendix C.4.
+We conduct experiments on five datasets of varying sizes and origins: the Transactions dataset (AI-Academy, 2021), the MIMIC-IV medical dataset (Johnson et al., 2023), two social networks datasets, Retweet (Zhao et al., 2015) and StackOverflow (Jure, 2014), and the Amazon reviews dataset (Jianmo, 2018).
+Dataset statistics and evaluation parameters are provided in Appendix C.4.
 We evaluate representative modeling methods from different groups:
   1. 1.
-    Statistical baselines. MostPopular Generates a constant output with the most popular label from the prefix and the average time step. Last copies multiple previous events with adjusted timestamps.
+    Rule-based baselines.
+MostPopular Generates a constant output with the most popular label from the prefix and the average time step.
+The Last N baseline outputs N previous events with adjusted timestamps.
   2. 2.
     Intensity-free approaches.
-We implement MAE-CE method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels [padhi2021tabgpt, mcdermott2024eventstreamgpt].
+We implement IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024).
   3. 3.
     Intensity-based approaches.
-We implement RMTPP [du2016rmtpp] method as an example of the TPP approach with a traditional RNN.
-We also add NHP [mei2017nhp], based on a continuous time LSTM architecture.
+We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN.
+We add NHP (Mei & Eisner, 2017), based on a continuous time LSTM architecture.
+We also evaluate an AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019).
   4. 4.
     Next-K approaches.
-We adapt MAE-CE and RMTPP for a direct prediction of multiple future events.
-This approach originates in time series analysis and has not been previously applied in the MTPP domain.
-The resulting methods are called MAE-CE-K and RMTPP-K, respectively.
+We adapt IFTPP and RMTPP to predict multiple future events directly, implementing IFTPP-K and RMTPP-K, respectively.
+These approaches originate in time series analysis and have not been previously applied in the MTPP domain.
   5. 5.
-    Reranking. We also implement HYPRO [xue2022hypro], which generates multiple hypotheses and selects the best sequence using a contrastive approach.
-Details on the methods can be found in Appendix A and Appendix C.1.
-Training specifics are provided in AppendixD.
-Hyperparameters for both methods and metrics are presented in Appendix E.
-The main evaluation results are presented in Table 1.
-Below, we will discuss these results from various perspectives and provide an additional analysis of the methods’ behavior.
-6.1 OTD vs T-mAP
-Evaluation results indicate that the OTD metric can yield low values even for simple statistical baselines.
-For instance, the Last baseline achieves the lowest OTD value on the Retweet dataset and outperforms Next-K methods on the Amazon dataset.
-In contrast, the T-mAP metric clearly differentiates between statistical baselines and deep models, as the former is incapable of predicting label probabilities or confidences.
+    Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses and selects the best sequence using a contrastive approach.
+Additional details on these methods are provided in Appendix A, with training specifics outlined in Appendix D.
+Hyperparameters for both the methods and metrics are presented in Appendix E.
+The main evaluation results are shown in Table 1.
+In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
+6.1 Metrics Comparison
+The evaluation results in Table 1 show that the OTD metric can yield low values even for simple rule-based baselines.
+For instance, the Last baseline achieves the lowest OTD values on the Retweet and Amazon datasets.
+In contrast, the T-mAP metric more clearly distinguishes between rule-based baselines and deep learning models.
+This difference arises from the ability to measure model confidence.
+OTD evaluates predicted labels alone, while T-mAP measures average precision by adjusting logits thresholds.
+Unlike rule-based baselines, deep models provide confidence scores, whereas baselines output only labels, resulting in low T-mAP scores for simple baselines.
+The HYPRO and NHP methods maximize T-mAP, while the lowest OTD values are achieved by statistical baselines, IFTPP, and RMTPP.
+The difference between the two metrics lies in balancing the importance of event order versus time prediction quality.
+OTD evaluates the first 5-10 events, and to minimize OTD, a model must correctly identify the events that constitute the beginning of the sequence.
+IFTPP and RMTPP jointly predict the next event’s time and label, leading to better ordering.
+In contrast, methods like NHP, AttNHP, and ODE predict time steps independently for each event type.
+When the differences between timestamps are small, these methods can result in random ordering, negatively affecting the OTD score.
+T-mAP, on the other hand, is less sensitive to event order as long as the time predictions are accurate.
+The HYPRO method, designed to differentiate between ground truth and predicted sequences using a discriminator model, achieves high T-mAP scores.
+This highlights T-mAP’s ability to reward models that generate realistic sequences.
+We also observed that methods with high T-mAP scores often have low next-event prediction accuracy.
+For instance, on the Retweet, Amazon, and MIMIC-IV datasets, high long-horizon performance is often accompanied by low next-event prediction scores.
+This is particularly true for the HYPRO method, whose loss function is heavily focused on long-horizon quality.
+The contrast on the MIMIC-IV dataset can be attributed to the large number of events with identical timestamps, making precise event ordering more difficult and impacting both next-event and OTD scores.
 6.2 Intensity-Based vs Intensity-Free
-Previous works have compared either intensity-based neural methods [du2016rmtpp, xue2022hypro] or intensity-free approaches [padhi2021tabgpt, mcdermott2024eventstreamgpt].
+Previous works mostly compared either intensity-based neural methods (Du et al., 2016; Xue et al., 2022) or intensity-free approaches (Padhi et al., 2021; McDermott et al., 2024).
+The only exception is IFTPP (Shchur et al., 2019), which did not compare with NHP and included only test set likelihood as an evaluation metric.
 This raises the question: which approach is superior?
-We compare the intensity-free MAE-CE method with intensity-based RMTPP and NHP approaches.
-Our results show that the intensity-free MAE-CE method excels in the next-event MAE prediction, as it optimizes this metric.
+We compare the intensity-free IFTPP method with intensity-based RMTPP, NHP, and ODE approaches.
+Table 1 shows that the intensity-free IFTPP method excels in the next-event MAE prediction, as it optimizes this metric.
 In other scenarios, the results vary between datasets.
-Therefore, we conclude that both approaches warrant attention in future research.
+Therefore, we conclude that there is no preferred solution, and both approaches warrant attention in future research.
 6.3 Autoregression vs Direct Horizon Prediction
-We compared MAE-CE and RMTPP methods with their Next-K counterparts.
-According to our results, autoregression generally performs better on Transactions, MIMIC-IV, and Amazon datasets.
-However, on social network datasets (Retweet and StackOverflow), Next-K models typically outperform autoregression in horizon prediction tasks.
-Despite this, little effort has been made to adapt Next-K models to MTPPs.
-Consequently, we recommend further research into this type of method.
+We compared the IFTPP and RMTPP methods with their Next-K counterparts.
+Our results indicate that autoregressive models generally perform better regarding next-event prediction quality, while Next-K approaches improve T-mAP scores on all datasets except Transactions.
+The OTD-based rankings vary across datasets, with no single approach consistently outperforming the others.
+Despite these findings, little effort has been made to adapt Next-K models to MTPPs, suggesting future research in this area.
 We also observed that the Next-K approaches exhibit some interesting properties compared to autoregression.
-For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Fig.
-4.
+For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Figure 5.
+The likely reason for this behavior is the dependency of the model’s future output on its past errors.
+Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict.
+This suggests that Next-K models have the potential to predict better the distributions of future labels, a factor that should be considered in future research.
 Figure 5: Entropy of label distributions as a function of the position in the generated sequence. Results for StackOverflow and Amazon datasets are provided in Appendix F.
-The likely reason for this behavior is the dependency of the model’s future output on its own past errors.
-Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict.
-This suggests that Next-K models have the potential to better predict the distributions of future labels, a factor that should be considered in future research.
 6.4 The Maximum Sequence Length
@@
 As shown in Fig.
-5, the optimal number of predicted events is usually less then the maximum number of events in the horizon.
-For example, in Transactions dataset, the optimal length ranges from 3 to 5, depending on the method.
-This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to manually limit the number of predictions in the target metric.
+6, the optimal number of predicted events is usually less than the maximum number of events on the horizon.
+For example, in the Transactions dataset, the optimal length ranges from 3 to 5, depending on the method.
+This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to limit the number of predictions manually.
 Figure 6: T-mAP dependency on the maximum length of the predicted sequence. Results for the StackOverflow and Amazon datasets are detailed in Appendix G.
-We therefore conclude that more attention should be paid to probability calibration between generation steps.
-This result may also indicate real progress in long-horizon prediction tasks, specifically the maximum horizon one can model accurately.
-T-mAP is the only metric capable of evaluating the optimal predicted sequence length, whereas the OTD metric does not depend on sequence length as it compares prefixes of fixed size.
-7 Related Work
-Temporal Point Processes.
-Traditional MTPP models make strong assumptions about the underlying generative process, resulting in Poisson and Hawkes processes [rizoiu2017hawkes-tutorial].
-Over the last decade, there has been a shift towards achieving greater model flexibility by applying neural architectures.
-Some works have utilized classical RNNs [du2016rmtpp] and transformers [zuo2020thp], while others have proposed architectures with continuous time [mei2017nhp, rubanova2019ode].
-In the HoTPP benchmark, we evaluate MTPP models with discrete and continuous time.
-Unlike other benchmarks, we also evaluate popular approaches from related fields, including next-event prediction models from event sequences [mcdermott2024eventstreamgpt, padhi2021tabgpt] and Next-K models from time series analysis [lim2021timeseriessurvey].
-MTPP Evaluation.
-Previous benchmarks have primarily focused on medical data or traditional MTPP datasets.
-EventStream-GPT [mcdermott2024eventstreamgpt] and TemporAI [saveliev2023temporai] consider only medical data and do not implement methods from the MTPP field, despite their applicability.
-Early MTPP benchmarks, such as Tick [bacry2017tick] and PyHawkes [^1] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches and are unable to work with modern neural networks.
-PoPPy [xu2018poppy] and EasyTPP [xue2023easytpp] implement neural methods; however, these benchmarks cannot be extended to use Next-K approaches.
-PoPPy does not evaluate long-horizon prediction at all.
-This limitation is partially addressed in EasyTPP, which measures the OTD metric at the end of each sequence.
-In this work, we address the limitations of the OTD metric and introduce a superior alternative, T-mAP, along with an effective GPU-accelerated implementation.
-8 Limitations
+Therefore, we conclude that greater attention should be given to probability calibration between generation steps.
+This finding also highlights progress in long-horizon prediction tasks, particularly determining the maximum horizon that can be accurately modeled.
+T-mAP is the only metric capable of evaluating the optimal predicted sequence length, while the OTD metric remains unaffected by sequence length as it compares fixed-size prefixes.
+7 Limitations and Future Work
+The proposed benchmark has several shortcomings.
 The training process for some methods is relatively slow.
@@
 Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training.
-Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating modern Transformer approaches[zuo2020thp] or Intensity-free training objective[shchur2019intensityfree].
-9 Future Research Directions
-In this work, we analyze long-horizon prediction evaluation and propose a new validation process.
-We strongly encourage future research to develop improved techniques for predicting future events and to establish simple baselines for better-measuring progress in the field.
+Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating hybrid (Deshpande et al., 2021) and diffusion models (Zhou et al., 2023).
+Our benchmark encourages future research to develop improved techniques for predicting future events and establish simple baselines for better measuring progress in the field.
 For example, Next-K models, which predict multiple future events without autoregression, show promising results in long-horizon prediction tasks.
@@
 We also highlight the importance of label distribution estimation and emphasize the need to improve confidence estimation and calibration.
+The T-mAP metric can be applied in domains beyond event sequences, like action recognition (Su & Hua, 2017).
+T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW.
+T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error.
+Our theoretical justification of T-mAP can potentially be adapted to the mAP estimation algorithms in computer vision.
 8 Conclusion
-In this paper, we propose the HoTPP benchmark to assess the quality of multiple events forecasting.
+In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting.
 Our extensive evaluations using established datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction does not necessarily correlate with superior performance in horizon prediction.
@@
 This, in turn, has the potential to significantly enhance the practical applications of sequential event prediction in various domains, fostering innovation and improving decision-making processes across a wide range of industries.
+Appendix A Background on Marked Temporal Point Processes Modeling
+Intensity-based approaches.
+Modeling the probability density function (PDF) $f^{*}(t_{i})=f(t_{i}|t_{1},\dots,t_{i-1})$ for the next event time is typically challenging, as it requires the additional constraint that its integral equals one.
+Instead, the non-negative intensity function $\lambda(t_{i})\geq 0$ is usually modeled.
+The following equation gives the relationship between the PDF and the intensity function:
 […]
-Appendix A Background on Marked Temporal Point Processes Modeling
-[…]
-[…]
-The derivation is provided in [rizoiu2017hawkes-tutorial].
+The derivation is provided in (Rizoiu et al., 2017).
 Different TPPs are characterized by their intensity function $\lambda(t_{i})$.
@@
 Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach.
-For details on the implementation of sampling, please refer to [rizoiu2017hawkes-tutorial].
+For details on the implementation of sampling, please refer to (Rizoiu et al., 2017).
 Recent research has focused on modeling complex intensity functions.
@@
 These approaches differ in the type of neural network used and the model of the intensity function.
-Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models, like NHP [mei2017nhp] and Neural ODEs [rubanova2019ode].
+Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei & Eisner, 2017) and Neural ODEs (Rubanova et al., 2019).
 The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
 Some methods evaluate the next event time distribution without using intensity functions.
-For example, intensity-free [shchur2019intensityfree] represents the distribution as a mixture of Gaussians or through normalizing flows.
-Other approaches do not predict the distribution directly but instead, solve a regression problem using MAE or RMSE loss.
+For example, intensity-free (Shchur et al., 2019) represents the distribution as a mixture of Gaussians or through normalizing flows.
+Instead of predicting the distribution directly, other approaches solve a regression problem using MAE or RMSE loss.
 However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction.
-Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution [bishop1994mdn].
+Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution (Bishop, 1994).
 In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
 Rescoring with HYPRO.
-HYPRO [xue2022hypro] is an extension applicable to any sequence prediction method capable of sampling.
+HYPRO (Xue et al., 2022) is an extension applicable to any sequence prediction method capable of sampling.
 HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample.
@@
 In our work, we apply HYPRO to the outputs of the RMTPP model.
+Next-K models.
+While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling (Lim & Zohren, 2021).
+These methods predict multiple future events at once, avoiding the need for autoregressive inference.
+The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression.
+The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length.
+However, applying a modified autoregressive approach can potentially address this limitation.
+In our work, we implement a Next-K variant of the IFTPP model, which is straightforward.
+Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences.
+The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1.
+Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events.
+Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
 Appendix B T-mAP Computation and Proofs
@@
 Example processes are illustrated in Fig.
-6.
+7.
 The vertices in each process do not repeat; otherwise, there would be two edges in either $m$ or $m^{\prime}$ with the same vertex, which contradicts the definition of a matching.
@@
 These quantities define precision and recall values along with the total number of ground truth events.
-A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|<\delta$.
+A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|\leq\delta$.
 Note that each target can be assigned to at most one prediction.
@@
 Its complexity is $\mathcal{O}(LBN^{3})$, where $L$ is the number of classes, $B$ is the number of sequence pairs, and $N$ is the maximum length of predicted and ground truth sequences.
+Calibration dependency.
+Calibration influences the weights assigned to the edges of the graph $\mathcal{G}$.
+T-mAP computation involves two key steps for each class label: matching and AP estimation.
+While average precision (AP) is invariant to monotonic transformations of predicted class logits, the matching step is only invariant to linear transformations.
+Specifically, the optimal matching seeks to minimize the total weight in the following form:
+$C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$ (9)
+A linear transformation of logits with a positive scaling factor will adjust the total weight accordingly, but the optimal matching will remain unchanged.
+Since the matching is performed independently for each class, we conclude that T-mAP is invariant to linear calibration.
 Appendix C HoTPP Benchmark Details
-C.1 Methods
-We implement popular methods from five groups: intensity-based, intensity-free, reranking, next-k, and statistical baselines.
-Intensity-based methods. We implement the RMTPP[du2016rmtpp] and NHP[mei2017nhp] methods.
-Intensity-free methods.
-As mentioned in Appendix A, this group includes intensity-free TPP and RMSE/MAE regression losses.
-We implement the MAE-CE approach, which uses MAE loss for time interval prediction and cross-entropy loss for labels.
-Reranking approaches. We implement HYPRO[xue2022hypro] with the RMTPP backbone.
-Next-K models.
-While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling [lim2021timeseriessurvey].
-These methods predict multiple future events at once, avoiding the need for autoregressive inference.
-The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression.
-The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length.
-However, applying a modified autoregressive approach can potentially address this limitation.
-In our work, we implement a Next-K variant of the MAE-CE model, which is quite straightforward.
-In particular, given K predictions and K ground truth events, we just compute MAE and cross-entropy losses between the pairs of corresponding events from two sequences.
-A bit more complicated is the Next-K variant of the RMTPP approach, as it violates Hawkes’s assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1.
-So, RMTPP-K, unlike RMTPP, can’t model dependencies between the predicted K events.
-Despite this, RMTPP-K is still a strong model with top-performing results on the StackOverflow dataset.
-In our work, we implement a Next-K variant of the MAE-CE model, which is straightforward.
-Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences.
-The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1.
-Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events.
-Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
-MostPopular baseline.
-This baseline analyzes the prefix sequence to identify the most frequent label and computes an average time step.
-The model generates constant output until the required duration (horizon) is reached.
-While this approach does not yield impressive results, it demonstrates moderate quality, especially in next-event prediction tasks.
-Last-K baseline.
-This method extracts the required number of events from the end of the prefix and shifts timestamps to the current value.
-Despite its simplicity, this baseline demonstrates OTD quality comparable to neural approaches on small datasets.
+C.1 Architecture
+HoTPP incorporates best practices from extensible and reproducible ML pipelines.
+It leverages PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments.
+Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility.
+The overall architecture is illustrated in Figure 8.
+HoTPP supports both discrete-time and continuous-time models, as well as RNN and Transformer architectures.
+Implementing a new method requires only the addition of essential components, while the rest can be configured through Hydra and configuration files.
+Figure 8: The architecture of the HoTPP library.
 C.2 Metrics
-MTPP models are typically evaluated by the quality of the next event prediction.
-These metrics usually include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label quality evaluation.
-Some works also evaluate test set likelihood as predicted by the model.
-We do not measure this quantity, as it is intractable for MAE-CE and Next-K models.
-Previous works have made progress in evaluating long-horizon prediction [mei2019imputing, xue2023easytpp] with optimal transport distance (OTD).
-We evaluate these metrics, as well as the novel T-mAP metric, which fixes the flaws of previous metrics, as discussed in Section 3.
-We set the maximum time error $\delta$ for the T-mAP metric twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$.
+MTPP models are typically evaluated based on the accuracy of next-event predictions.
+Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts, and accuracy or error rate for label predictions.
+Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for IFTPP and Next-K models.
+Previous works have advanced long-horizon evaluation using optimal transport distance (OTD) (Mei et al., 2019; Xue et al., 2023).
+In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.
 C.3 Backbones
-We use two types of architectures in our experiments.
-The first one is the GRU network[cho2014gru], which is one of the top-performing neural architectures in the event sequences domain [babaev2022coles].
-We also implement a continuous-time LSTM (CT-LSTM) from the NHP method [mei2017nhp].
-As CT-LSTM requires a special loss function and increases training time, we apply it only in the NHP method.
-In other cases, we prefer GRU.
-Another advantage of GRU is that its output is equal to its hidden state, which simplifies intermediate hidden state estimation for autoregressive inference starting from the middle of a sequence.
+We use three types of architectures in our experiments.
+The first is the GRU network (Cho et al., 2014), which is one of the top-performing neural architectures for event sequences (Babaev et al., 2022).
+We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei & Eisner, 2017).
+Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases.
+An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence.
+Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
 C.4 Datasets
-For the first time, we combine domains of financial transactions, social networks, and medical records in a single evaluation benchmark.
-Specifically, we provide evaluation results on a transactional dataset (AgePred) [babaev2022coles], MIMIC-IV [johnson2020mimic], social networks datasets (Retweet [zhao2015seismic] and StackOverflow [jure2014snap]), and Amazon dataset from the area of recommender systems [jure2014amazon].
-These datasets have different underlying processes.
-For example, social network datasets are subject to cascades [zhao2015seismic].
-Medical records show repetitive patterns and transactional data resembles daily activities with some regularity and large uncertainty.
+For the first time, we combine domains such as financial transactions, social networks, and medical records into a single evaluation benchmark.
+Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)).
+These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns, and transactional data reflects daily activities, combining regularity with significant uncertainty.
 The dataset statistics are presented in Table 2.
-Transactions have the largest average sequence length and the largest number of classes.
-MIMIC-IV has the largest number of sequences.
-Retweet is medium size and Amazon with StackOverflow can be considered small datasets.
+The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences.
+Retweet is of medium size, whereas Amazon and StackOverflow can be considered small datasets.
 Table 2: Datasets statistics
@@
 Notes on MIMIC-IV preprocessing
-MIMIC-IV is a publicly available electronic health record database, that includes patient diagnoses, lab measurements, procedures, and treatments.
-We use data preprocessing from EventStreamGPT to construct intermediate representations in EFGPT format.
-The following entities are obtained: subjects data frame with time-independent patient records, events data frame with lists of event types happening to subjects at specific timestamps, and measurements data frame containing specific values of time-dependent measurements and linked to events data frame.
-We join events and measurements and form sequences of events for subjects.
-The classification labels can be generally separated into diagnoses in the form of ICD codes and types of events, such as admissions, procedures, measurements, etc.
-The main issue is the inability to use either diagnoses only or event types only.
-In the former case, the sequences become too short, while in the latter, the sequences become too repetitive, with periodic events, such as treatment start/finish, dominating the data.
-However, using both diagnoses and event types simultaneously presents another problem, as diagnosis designations are distributed sparsely in the constant stream of repeat procedures.
-This imbalanced class distribution causes poor performance.
-Therefore, we further filter the data by removing duplicate events between diagnoses.
-This way, we can retain useful treatment data and preserve class balance.
-The final labels are formed by converting ICD-9 and ICD-10 codes to ICD-10 chapters using General Equivalence Mapping (GEM) and adding event types as additional classes.
-The conversion is needed because the number of ICD codes is too large to use them as classes directly.
-We also address the issue of multiple diagnoses being designated in a single event by sorting timestamps for reproducibility.
+MIMIC-IV is a publicly available electronic health record database that includes patient diagnoses, lab measurements, procedures, and treatments.
+We leverage the data preprocessing pipeline from EventStreamGPT to construct intermediate representations in EFGPT format.
+This process yields three key entities: a subjects data frame containing time-independent patient records, an events data frame listing event types occurring to subjects at specific timestamps, and a measurements data frame with time-dependent measurement values linked to the events data frame.
+We combine events and measurements to create sequences of events for each subject.
+The classification labels are generally divided into two categories: diagnoses, represented by ICD codes, and event types, such as admissions, procedures, and measurements.
+A key challenge is that relying solely on diagnoses results in sequences that are too short, while using only event types leads to highly repetitive sequences dominated by periodic events, such as treatment start and finish.
+However, using both diagnoses and event types together introduces another issue: diagnoses are sparsely distributed within a constant stream of repeated procedures, leading to imbalanced class distributions and poor performance.
+To mitigate this, we filter the data by removing duplicate events between diagnoses, allowing us to retain useful treatment data while preserving class balance.
+The final labels are created by converting ICD-9 and ICD-10 codes to ICD-10 chapters using General Equivalence Mapping (GEM) and adding event types as additional classes.
+This conversion is necessary because the number of ICD codes is too large to use them directly as classes.
+Additionally, we address the issue of multiple diagnoses occurring in a single event by sorting timestamps for reproducibility.
+C.5 Performance improvements
+The HoTPP benchmark provides highly optimized training and inference procedures for efficient evaluation on datasets containing up to tens of millions of events.
+First, we implement parallel RNN inference, which reuses computations when inference is initiated from multiple starting points within a sequence.
+Additionally, we optimize the code and apply PyTorch Script to the CT-LSTM model from NHP and the ODE model.
+CT-LSTM from EasyTPP and the ODE model from the original repository serve as baselines.
+The timing results are presented in Table 3.
+Experiments were conducted using a synthetic batch with a size of 64, a sequence length of 100, and an embedding dimension of 64.
+Evaluation was performed on a single Nvidia RTX 4060 GPU.
+The results show that HoTPP is 17 times faster at RNN autoregressive inference compared to simple prefix extension.
+HoTPP also accelerates CT-LSTM and ODE by 4 to 8 times.
+These optimizations significantly extend the applicability of the implemented methods to larger-scale datasets [^8] Experiments are documented in the “notebooks” folder of the HoTPP repository.
+Additionally, HoTPP offers a GPU implementation of the Hungarian algorithm, which also finds applications in computer vision.
+Table 3: HoTPP computation speed improvements in terms of seconds per batch.
+Implementation Study
+RNN autoreg. inference CT-LSTM inference CT-LSTM train ODE inference ODE train
+Baseline 2.44 0.0158 0.0519 0.08 0.162
+HoTPP 0.14 0.0025 0.0276 0.01 0.048
 Appendix D Training details
-We trained each model with the Adam optimizer, learning rate set to $0.001$, and learning rate scheduler, decreasing learning rate by 20
-We used NVIDIA V100 and A100 GPUs for computation.
-Each method was trained with a single GPU.
-The training time depends on the dataset and method.
-It varies from 5 minutes for RMTPP on StackOverflow to 40 minutes for the same method on Transactions or even 15 hours for the NHP method on the Transactions dataset.
-Multiseed evaluation takes 5 times more time to finish.
+We trained each model using the Adam optimizer with a learning rate of 0.001 and a scheduler that reduces the learning rate by 20% after every 5 epochs.
+Gradient clipping was applied, with a maximum L2-norm set to 1.
+We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060.
+Each method was trained on a single GPU.
+The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset, and up to 15 hours for NHP on Transactions.
+Multi-seed evaluations took approximately five times longer to complete.
 Appendix E Hyperparameters
@@
 The results are shown in Figure 9.
-OTD cost has little impact on methods ranking but largely affects the metric scale.
-The choice of $\delta$ parameter in T-mAP can, in some cases, affect methods ranking, but if it is not too small and too large, the metric demonstrates stable ranking.
-We set the $\delta$ parameter to about 10-30% of the horizon duration.
-Evaluation hyperparameters are presented in Table 3. Dataset-specific training parameters are presented in Table 4.
+The OTD cost has little effect on the ranking of methods but significantly influenced the metric scale.
+The choice of the $\delta$ parameter in T-mAP can, in some cases, affect method rankings, but as long as $\delta$ is not set too small or too large, the metric demonstrates stable rankings.
+We set the $\delta$ parameter to approximately 10-30% of the horizon duration.
+Evaluation hyperparameters are presented in Table 4. Dataset-specific training parameters are listed in Table 5.
 Table 4: Evaluation hyperparameters.
-Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 16 5 1 7 (week) 2 MIMIC-IV 16 5 2 28 (days) 4 Retweet 16 10 15 180 (seconds) 30 Amazon 16 5 1 10 (unk.
-unit) 2 StackOverflow 16 10 1 10 (minutes) 2
+Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 32 5 1 7 (week) 2 MIMIC-IV 32 5 2 28 (days) 4 Retweet 32 10 15 180 (seconds) 30 Amazon 32 5 1 10 (unk.
+unit) 2 StackOverflow 32 10 1 10 (minutes) 2
 Table 5: Training hyperparameters.
@@
 Size Hidden Size Head hiddens Transactions 30 1200 256 512 512 $\rightarrow$ 256 MIMIC-IV 30 64 16 64 64 Retweet 30 264 16 64 64 Amazon 60 94 32 64 64 StackOverflow 60 101 32 64 64
+Figure 9: The dependency of metric values on the metric parameter.
 Appendix F Labels Entropy Degradation
-We omitted some datasets in the main part of the paper for simplicity. In Fig. 7 one can see the dependency of the labels distribution entropy depending on step size for all datasets.
-Figure 7: Labels distribution entropy depending on the generated sequence position.
+For simplicity, some datasets were omitted in the main part of the paper. In Figure 10, we show the dependency of label distribution entropy on step size for all datasets.
+Figure 10: Entropy of label distributions as a function of the position in the generated sequence.
 Appendix G The Optimal Sequence Length
-We omitted some datasets in the main part of the paper for simplicity. In Fig. 8 you can see the dependency of the labels distribution entropy depending on step size for all datasets.
+For simplicity, some datasets were omitted from the main part of the paper. Figure 11 illustrates the relationship between label distribution entropy and step size across all datasets.
 Figure 11: T-mAP dependency on the maximum length of the predicted sequence.
-Figure 9: The dependency of metric values on the metric parameter.

### v3

@@
 HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?
-Abstract
-Accurately forecasting multiple future events within a given time horizon is crucial for finance, retail, social networks, and healthcare applications.
-Event timing and labels are typically modeled using Marked Temporal Point Processes (MTPP), with evaluations often focused on next-event prediction quality.
-While some studies have extended evaluations to a fixed number of future events, we demonstrate that this approach leads to inaccuracies in handling false positives and false negatives.
-To address these issues, we propose a novel evaluation method inspired by object detection techniques from computer vision.
-Specifically, we introduce Temporal mean Average Precision (T-mAP), a temporal variant of mAP, which overcomes the limitations of existing long-horizon evaluation metrics.
-Our extensive experiments demonstrate that models with strong next-event prediction accuracy can yield poor long-horizon forecasts and vice versa, indicating that specialized methods are needed for each task.
-To support further research, we release HoTPP [^1] https://github.com/ivan-chai/hotpp-benchmark, the first benchmark designed explicitly for evaluating long-horizon MTPP predictions.
-HoTPP includes large-scale datasets with up to 43 million events and provides optimized procedures for both autoregressive and parallel inference, paving the way for future advancements in the field.
-1 Introduction
-The world is full of events.
+Abstract.
+Forecasting multiple future events within a given time horizon is essential for applications in finance, retail, social networks, and healthcare.
+Marked Temporal Point Processes (MTPP) provide a principled framework to model both the timing and labels of events.
+However, most existing research focuses on predicting only the next event, leaving long-horizon forecasting largely underexplored.
+To address this gap, we introduce HoTPP, the first benchmark specifically designed to rigorously evaluate long-horizon predictions.
+We identify shortcomings in widely used evaluation metrics, propose a theoretically grounded T-mAP metric, present strong statistical baselines, and offer efficient implementations of popular models.
+Our empirical results demonstrate that modern MTPP approaches often underperform simple statistical baselines.
+Furthermore, we analyze the diversity of predicted sequences and find that most methods exhibit mode collapse.
+Finally, we analyze the impact of autoregression and intensity-based losses on prediction quality, and outline promising directions for future research.
+The HoTPP source code, hyperparameters, and full evaluation results are available on GitHub [^1] https://github.com/ivan-chai/hotpp-benchmark.
+copyright: acmlicensedjournalyear: 2025
+1. Introduction
 Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information.
@@
 Event sequences differ fundamentally from other data types.
-Unlike tabular data (Wang & Sun, 2022), ESs include timestamps and possess an inherent order.
-In contrast to time series data (Lim & Zohren, 2021), ESs are characterized by irregular time intervals and structured descriptions of each event.
+Unlike tabular data (Wang and Sun, 2022), ESs include timestamps and possess an inherent order.
+In contrast to time series data (Lim and Zohren, 2021), ESs are characterized by irregular time intervals and additional data fields, such as event types.
 These differences necessitate the development of specialized models and evaluation practices.
-Sequence modeling is the primary task in Event Sequences (ESs) domain.
-In its simplest form, each event is defined by its type and occurrence time, a framework commonly known as Marked Temporal Point Processes (MTPP) (Rizoiu et al., 2017).
-Some studies extend MTPP to incorporate additional data fields and model complex dependencies between them (McDermott et al., 2024).
-However, the majority of MTPP approaches, as well as their evaluation pipelines, primarily focus on predicting the next event based on historical data.
+Irregularly spaced timestamps are commonly modeled via Temporal Point Processes (TPP) (Rizoiu et al., 2017).
+Assigning categorical labels to these events leads to the Marked Temporal Point Process (MTPP) formalism.
+Since most event-sequence data can be viewed as MTPPs augmented with additional fields (McDermott et al., 2024), accurately modeling event times and labels remains a central challenge in the broader event-stream (ES) domain.
 In practice, a common question arises: what events will occur, and when, within a specific time horizon?
-Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks.
+Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks (Xue et al., 2022; Zhou et al., 2025).
 For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions.
@@
 Another difficulty lies in evaluation.
-Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su & Hua, 2017).
-Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths.
-Still, the limitations of this metric have not been previously considered.
+Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su and Hua, 2017).
+Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths (Mei et al., 2019; Xue et al., 2022, 2023).
+However, the limitations of this metric have not been previously considered.
+Figure 1.
+T-mAP evaluation pipeline.
+Unlike previous approaches, T-mAP evaluates sequences of variable lengths within the prediction horizon.
+It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
 In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies.
-Our key contributions are as follows:
-  1. 1.
-    We demonstrate that widely used evaluation methods for MTPPs often overlook critical aspects of model performance.
-We demonstrate that simple rule-based baselines can sometimes outperform popular deep learning methods when evaluated using OTD.
-  2. 2.
-    We introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision.
-T-mAP evaluates variable length sequences within a specified time horizon, as illustrated in Figure 1.
-Unlike previous approaches, T-mAP accurately accounts for false positives and false negatives while being invariant to linear calibration.
-Additionally, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
-  3. 3.
-    Using our established methodology, we demonstrate that high next-event prediction accuracy does not necessarily translate into high-quality long-horizon forecasts; in many cases, our experiments show the opposite.
-This highlights the necessity of developing specialized models for the long-horizon prediction task.
-  4. 4.
-    We release HoTPP, a new open-source benchmark designed to facilitate long-horizon event sequence prediction research.
-HoTPP brings together datasets and methods from various domains, including financial transactions, social networks, healthcare, and recommender systems, greatly expanding the diversity and scale of data compared to prior benchmarks.
-Additionally, we offer an efficient inference algorithm necessary for large-scale evaluations.
-Figure 1: Temporal mAP (T-mAP) evaluation pipeline.
-Unlike previous methods, T-mAP evaluates sequences of variable lengths within the prediction horizon.
-It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
-2 Related Work
+Our key contributions are summarized as follows:
+  1. (1)
+    We introduce HoTPP, the first open-source benchmark specifically designed for long-horizon event forecasting.
+By considering datasets and methods spanning finance, social networks, healthcare, and recommender systems, HoTPP significantly broadens the diversity and scale of data compared to previous benchmarks (Xue et al., 2023; McDermott et al., 2024; Xu, 2018).
+Moreover, we provide optimized training and inference algorithms to support large-scale studies.
+  2. (2)
+    We demonstrate that widely used evaluation metrics for MTPPs overlook critical aspects of model performance.
+In particular, they are sensitive to calibration and incorrectly count false positives and false negatives.
+To address these issues, we introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision (Lin et al., 2014).
+Furthermore, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
+  3. (3)
+    Guided by our methodology, we evaluate the forecasting performance of a diverse set of approaches.
+Our findings indicate that methods excelling at next-event prediction often underperform in long-horizon scenarios, and vice versa.
+Moreover, in 6 out of 10 long-horizon tasks, simple statistical baselines either outperform or match modern MTPP models, highlighting the need for more effective solutions.
+  4. (4)
+    Finally, we quantify the entropy of predicted labels and demonstrate that most modern approaches tend to predict only the most frequent labels, suffering from the problem known as mode collapse.
+This issue severely restricts practical applicability and motivates further research into more accurate prediction strategies.
+2. Related Work
 Marked Temporal Point Processes.
@@
 A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing.
-Depending on the time-step distribution, the process is called Poisson or Hawkes.
+Depending on the form of time-step distribution, the process is called Poisson or Hawkes.
 Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures.
-Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei & Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024).
+Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei and Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024).
 In this work, we evaluate MTPP models with both discrete and continuous time architectures.
-Unlike previous benchmarks, we also assess simple rule-based predictors and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021) and Next-K models from time series analysis (Lim & Zohren, 2021).
+Unlike previous benchmarks, we also assess simple rule-based baselines and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021).
 For more details on MTPP modeling, refer to Appendix A.
@@
 Previous benchmarks have primarily focused on medical data or traditional MTPP datasets.
-EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev & van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability.
+EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev and van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability.
 Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks.
-While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and Next-K approaches.
+While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and horizon prediction approaches, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025).
 Furthermore, PoPPy does not evaluate long-horizon predictions at all.
 EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
-Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su & Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided.
+Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su and Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided.
 This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW).
@@
 Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
-3 Limitations of the Next-Event and OTD Metrics
-Figure 2: Comparison of OTD and T-mAP metrics.
-In example 1.a prefix evaluation of three events using OTD results in an incorrect, false negative.
-T-mAP addresses this issue by comparing all events within the time horizon, as shown in example 1.b.
-Additionally, OTD evaluates only the label with the maximum probability.
-From the OTD perspective, cases 2.a and 2.b have the same quality.
-In contrast, T-mAP evaluates the entire distribution, yielding a low score of 0.33 in case 2.a and a high score of 1 in case 2.b.
+Figure 2.
+Comparison of the OTD and T-mAP metrics.
+In the example 1.a, OTD’s prefix-based evaluation of three events erroneously produces a false negative, whereas T-mAP in example 1.b compares all events within the specified time horizon for a more accurate assessment.
+Moreover, OTD evaluates only the label with the highest probability, treating cases 2.a and 2.b as equivalent.
+By contrast, T-mAP considers the entire distribution, yielding a low score (0.33) for 2.a and a perfect score (1) for 2.b.
+3. T-mAP Metric
+In this section, we show that widely used evaluation metrics for long-horizon prediction often miscount false positives (FP) and false negatives (FN).
+We then introduce Temporal mean Average Precision (T-mAP), a novel, theoretically grounded metric that addresses these issues.
+In addition, T-mAP is invariant to linear calibration of model outputs, simplifying and stabilizing comparisons across different methods.
+Additional T-mAP experiments on sequences with highly irregular time steps are presented in Appendix H.
+Long-tail performance measurement is discussed in Appendix I.
+3.1. Limitations of the Next-Event and OTD Metrics
 MTPPs are typically evaluated based on the accuracy of next-event predictions, with time and type predictions assessed independently.
@@
 As a result, long-horizon evaluation metrics are necessary.
-Recent studies have advanced the measurement of horizon prediction quality by employing Optimal Transport Distance (OTD) (Mei et al., 2019).
+Recent studies have advanced the measurement of horizon prediction quality by employing OTD (Mei et al., 2019).
 This metric is analogous to the edit distance between the predicted event sequence and the ground truth.
@@
 OTD finds the minimum cost among all possible matchings:
-$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
+(1) $\mathrm{OTD}(S_{p},S_{gt})=\\ \min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$
 where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost.
@@
 It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
-OTD is computed between fixed-size prefixes, but this approach limits metric flexibility in assessing models with imprecise time step predictions, as illustrated in Fig.
+OTD is computed over fixed-size prefixes, limiting its flexibility when models produce imprecise time-step predictions, as shown in Fig.
 2.1.
-If the model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives.
-This outcome is inaccurate because allowing the model to predict more events would alter the evaluation.
-Conversely, if the time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of false positives.
-This is undesirable as it prevents the inclusion of additional ground truth events in the evaluation.
-Therefore, evaluating a dynamic number of events is essential to properly align with the ground truth’s time horizon.
-MTPP evaluation metrics can be categorized into two groups: those considering event indices and those assessing time and label prediction quality independently of actual positions.
-Next-event metrics rely on ordering, even when events share identical timestamps, making the order meaningless.
-While Optimal Transport Distance (OTD) itself is invariant to order, the extraction of sequences for comparison is influenced by event indices.
-For example, when OTD compares prefixes of length $K$, the algorithm must place $K$ correctly predicted events at the beginning of a sequence to minimize the OTD score.
-As a result, OTD, like next-event metrics, can depend on event ordering even when the order cannot be uniquely determined.
-This dependency is non-trivial and difficult to measure or control accurately.
-Another limitation of the OTD metric is its inability to evaluate the full predicted distribution of labels, as shown in Figure 2.2.
+When a model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives.
+This outcome is misleading because considering a longer predicted sequence would alter the evaluation.
+Conversely, if the predicted time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of incorrect false positives.
+Therefore, evaluating a dynamic number of events is essential to align with the ground truth’s time horizon properly.
+Figure 3. T-mAP computation pipeline.
+Metrics for MTPP evaluation can be categorized into two groups: those relying on event indices and those that measure time and label prediction quality independently of event sequential positions.
+Next-event metrics depend on ordering, even if multiple events share the same timestamp, which makes the order meaningless.
+Although OTD is formally invariant to order, extracting sequences for comparison still depends on event indices.
+For instance, when OTD compares prefixes of length $K$, the first $K$ predicted events must align with the ground truth to minimize the OTD score.
+As a result, OTD, like next-event metrics, can be influenced by ordering even when it cannot be uniquely determined.
+This dependency is non-trivial and cannot be easily measured or controlled.
+Another limitation of the OTD metric is its inability to evaluate the full predicted distribution of labels, as shown in Fig.
+2.2.
 OTD considers only the labels with the highest probability, ignoring the complete distribution.
@@
 Therefore, we aim to develop an evaluation metric that accurately captures performance across common and rare classes.
-4 Temporal mAP: Prediction as Detection
+3.2. Temporal mAP: Definition and Computation
 In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration.
@@
 1.
-The goal of object detection is to localize objects within an image and identify their types.
+Object detection aims to localize objects within an image and identify their types.
 In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes.
@@
 These adjustments result in significant differences in formulation and computation, detailed below.
-4.1 Definition
+3.2.1. Definition
 T-mAP is parameterized by the horizon length $T$ and the maximum allowed time error $\delta$.
@@
 For any threshold value $h$, we can select a subset of the predicted sequence $S^{l}_{p}$ with scores exceeding the threshold:
-$S^{l}_{>}(h)=\{t^{p}_{i}:\exists(t^{p}_{i},s^{p}_{i})\in S^{l}_{p},s^{p}_{i}>h\}.$ (2)
+(2) $S^{l}_{>}(h)=\{t^{p}_{i}:\exists(t^{p}_{i},s^{p}_{i})\in S^{l}_{p},s^{p}_{i}>h\}.$
 By definition, a predicted event $i$ can be matched with a ground truth event $j$ iff $|t^{p}_{i}-t^{gt}_{j}|\leq\delta$, meaning the time difference between the predicted and ground truth events is less than or equal to $\delta$.
@@
 Finally, T-mAP is defined as the average AP across all classes.
-4.2 Computation
+Figure 4. T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value.
+3.2.2. Computation
 Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP.
-This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker & Volgenant, 1988).
+This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker and Volgenant, 1988).
 The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
-Figure 3: T-mAP computation pipeline.
 For each pair of sequences $S^{l}_{p}$ and $S^{l}_{gt}$ we define a weighted bipartite graph $\mathcal{G}(S^{l}_{p},S^{l}_{gt})$ with $|S^{l}_{p}|$ vertices in the first and $|S^{l}_{gt}|$ vertices in the second part.
@@
 For any threshold $h$ there exists an optimal matching in the graph $\mathcal{G}_{h}$, that is a subset of an optimal matching in the full graph $\mathcal{G}$:
-$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (3)
+(3) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
 According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig.
@@
 The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in Appendix B.
-4.3 T-mAP Hyperparameters
+3.2.3. T-mAP Hyperparameters
 T-mAP has two hyperparameters: the maximum allowed time delta $\delta$ and evaluation horizon $H$.
@@
 Therefore, depending on the dataset, we select $H$ to be approximately 3-7 times larger than $\delta$, ensuring the horizon captures an average of 6-15 events.
-The empirical study of T-mAP hyperparameters is presented in Figure 4.
+The empirical study of T-mAP hyperparameters is presented in Fig.
+4.
 As shown, the chosen $\delta$ parameter is positioned to the right of the initial slope of the parameter-quality curve, ensuring an optimal balance between time granularity and task difficulty.
-Figure 4: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value. Results for the StackOverflow and Amazon datasets are provided in the Appendix E.
-5 HoTPP Benchmark
+4. HoTPP Benchmark
 The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox.
 Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction.
-HoTPP differs from prior MTPP benchmarks by including simple rule-based baselines and next-k models, simultaneously predicting multiple future events.
-The HoTPP benchmark is designed with a focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
+HoTPP also differs from prior MTPP benchmarks by including simple rule-based baselines and long-horizon prediction methods that simultaneously predict multiple future events, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025).
+The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
 Simplicity and Extensibility.
@@
 Core components are reusable through the Hydra configuration library (Yadan, 2019).
-For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, and ODE.
-Another example is our Next-K implementation, which can be applied to IFTTP or RMTPP.
+For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, or ODE.
+Furthermore, each method’s backbone architecture can be easily switched between RNNs and Transformers.
+Methods such as IFTPP, RMTPP, and Diffusion are compatible with most HuggingFace models (Wolf et al., 2020).
 Evaluation Stability.
@@
 For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing.
-Previous long-horizon evaluation pipelines (Xue et al., 2022; 2023) made predictions only at the end of each sequence, resulting in a limited number of predictions and reduced evaluation stability.
-To address this limitation, we evaluate long-horizon predictions at multiple intermediate points.
+This possesses a unique challenge in the long-horizon forecasting task.
+Previous long-horizon evaluation pipelines (Xue et al., 2022, 2023) made predictions only at the end of each sequence, resulting in a limited number of measurements and a low evaluation stability.
+To address this limitation, we evaluate long-horizon predictions at multiple intermediate points, decreasing the variance of evaluation metrics.
+Table 1. Datasets statistics
+Dataset # Sequences # Events Mean Length Mean Horizon Mean Duration # Classes OTD T-mAP Length Steps Cost $\delta$ Horizon Transactions 50k 43.7M 875 9.0 719 203 5 1 2 7 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 5 2 4 28 Retweet 23k 1.3M 56.4 14.7 1805 3 10 15 30 180 Amazon 9k 403K 43.6 14.8 22.1 16 5 1 2 10 StackOverflow 2k 138K 64.2 12.0 55.3 22 10 1 2 10
 Reproducibility.
 The HoTPP benchmark ensures reproducibility at multiple levels.
-First, we use the PytorchLightning library (Falcon & The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results.
+First, we use the PytorchLightning library (Falcon and The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results.
 Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly.
 Finally, we specify the environment in a Dockerfile.
+Table 2. Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
+Method Mean length Next-event Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 5.0 32.69$\pm$ 0.00 1.09$\pm$ 0.00 0.649$\pm$ 0.000 7.03$\pm$ 0.00 / 7.05$\pm$ 0.00 2.86$\pm$ 0.00 / 3.10$\pm$ 0.00 HistoryDensity 154.2 32.78$\pm$ 0.00 0.85$\pm$ 0.00 1.063$\pm$ 0.000 7.42$\pm$ 0.00 / 7.44$\pm$ 0.00 7.92$\pm$ 0.00 / 7.51$\pm$ 0.00 IFTPP 13.3 38.27$\pm$ 0.03 4.65$\pm$ 0.03 0.637$\pm$ 0.001 6.79$\pm$ 0.01 / 6.81$\pm$ 0.01 6.20$\pm$ 0.10 / 6.08$\pm$ 0.14 IFTPP-T 12.8 37.68$\pm$ 0.03 4.30$\pm$ 0.03 0.640$\pm$ 0.001 6.84$\pm$ 0.02 / 6.85$\pm$ 0.02 5.80$\pm$ 0.27 / 5.55$\pm$ 0.19 RMTPP 8.7 38.28$\pm$ 0.04 4.64$\pm$ 0.02 0.693$\pm$ 0.005 6.68$\pm$ 0.01 / 6.70$\pm$ 0.01 7.41$\pm$ 0.35 / 7.12$\pm$ 0.10 NHP 9.4 35.47$\pm$ 0.12 3.36$\pm$ 0.02 0.693$\pm$ 0.005 6.99$\pm$ 0.03 / 6.99$\pm$ 0.03 5.66$\pm$ 0.17 / 5.48$\pm$ 0.12 AttNHP 6.8 30.68$\pm$ x.xx 1.09$\pm$ x.xx 0.749$\pm$ x.xxx 7.47$\pm$ x.xx / 7.48$\pm$ x.xx 1.50$\pm$ x.xx / 1.49$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 7.3 22.50$\pm$ x.xx 4.62$\pm$ x.xx 0.694$\pm$ x.xxx 7.62$\pm$ x.xx / 7.62$\pm$ x.xx 8.20$\pm$ x.xx / 7.54$\pm$ x.xx Diffusion 10.4 36.05$\pm$ 0.07 3.66$\pm$ 0.03 0.650$\pm$ 0.002 6.86$\pm$ 0.01 / 6.88$\pm$ 0.00 5.95$\pm$ 0.05 / 6.04$\pm$ 0.08 MIMIC-IV MostPopular 2.7 4.80$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.75$\pm$ 0.00 / 19.75$\pm$ 0.00 1.63$\pm$ 0.00 / 1.56$\pm$ 0.00 HistoryDensity 160.0 0.74$\pm$ 0.00 2.63$\pm$ 0.00 3.23$\pm$ 0.00 18.94$\pm$ 0.00 / 18.92$\pm$ 0.00 4.25$\pm$ 0.00 / 4.19$\pm$ 0.00 IFTPP 12.4 58.52$\pm$ 0.06 46.04$\pm$ 0.26 3.00$\pm$ 0.02 11.53$\pm$ 0.03 / 11.55$\pm$ 0.03 22.56$\pm$ 0.13 / 22.31$\pm$ 0.15 IFTPP-T 11.7 58.57$\pm$ 0.04 47.96$\pm$ 1.51 3.03$\pm$ 0.01 11.46$\pm$ 0.02 / 11.48$\pm$ 0.02 24.07$\pm$ 0.31 / 23.68$\pm$ 0.29 RMTPP 9.8 58.24$\pm$ 0.14 45.62$\pm$ 1.28 3.93$\pm$ 0.03 13.69$\pm$ 0.05 / 13.76$\pm$ 0.06 22.00$\pm$ 0.28 / 21.62$\pm$ 0.29 NHP 3.0 23.19$\pm$ 0.61 11.67$\pm$ 0.39 4.71$\pm$ 0.31 17.92$\pm$ 0.34 / 17.92$\pm$ 0.33 9.76$\pm$ 0.41 / 9.80$\pm$ 0.38 AttNHP 6.2 44.99$\pm$ 1.11 31.88$\pm$ 0.64 2.95$\pm$ 0.20 14.30$\pm$ 0.16 / 14.31$\pm$ 0.16 19.98$\pm$ 0.79 / 19.74$\pm$ 0.71 ODE 11.6 44.16$\pm$ 0.76 27.48$\pm$ 1.04 2.91$\pm$ 0.01 14.36$\pm$ 0.20 / 14.39$\pm$ 0.20 16.37$\pm$ 0.75 / 16.19$\pm$ 0.74 HYPRO 7.4 45.69$\pm$ x.xx 45.54$\pm$ x.xx 3.98$\pm$ x.xx 14.68$\pm$ x.xx / 14.75$\pm$ x.xx 16.41$\pm$ x.xx / 16.13$\pm$ x.xx Diffusion 15.4 46.22$\pm$ 5.31 33.67$\pm$ 3.66 2.99$\pm$ 0.01 13.25$\pm$ 0.15 / 13.28$\pm$ 0.14 22.90$\pm$ 0.10 / 22.82$\pm$ 0.12 Retweet MostPopular 9.9 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 141.7$\pm$ 0.0 / 139.1$\pm$ 0.0 29.72$\pm$ 0.00 / 30.02$\pm$ 0.00 HistoryDensity 14.9 58.62$\pm$ 0.00 39.86$\pm$ 0.00 20.68$\pm$ 0.00 178.5$\pm$ 0.0 / 177.0$\pm$ 0.0 44.48$\pm$ 0.00 / 44.60$\pm$ 0.00 IFTPP 23.1 59.75$\pm$ 0.25 46.23$\pm$ 0.31 18.27$\pm$ 0.05 165.8$\pm$ 0.2 / 164.8$\pm$ 0.3 43.93$\pm$ 0.69 / 40.59$\pm$ 0.84 IFTPP-T 23.2 59.88$\pm$ 0.06 46.07$\pm$ 0.11 18.47$\pm$ 0.06 167.4$\pm$ 1.3 / 166.1$\pm$ 1.1 43.73$\pm$ 1.88 / 40.05$\pm$ 1.73 RMTPP 16.6 60.01$\pm$ 0.06 46.81$\pm$ 0.03 18.60$\pm$ 0.04 169.2$\pm$ 1.0 / 168.0$\pm$ 1.0 49.03$\pm$ 0.15 / 45.60$\pm$ 0.28 NHP 18.1 60.05$\pm$ 0.08 46.84$\pm$ 0.13 18.39$\pm$ 0.12 167.0$\pm$ 1.7 / 165.7$\pm$ 1.7 49.25$\pm$ 0.37 / 45.64$\pm$ 0.26 AttNHP 29.8 60.06$\pm$ 0.03 46.66$\pm$ 0.02 18.32$\pm$ 0.02 174.7$\pm$ 0.7 / 172.9$\pm$ 0.7 26.74$\pm$ 1.16 / 24.24$\pm$ 0.94 ODE 17.7 59.93$\pm$ 0.06 46.36$\pm$ 0.30 18.53$\pm$ 0.09 168.2$\pm$ 1.1 / 166.7$\pm$ 1.0 48.95$\pm$ 0.30 / 45.45$\pm$ 0.38 HYPRO 16.2 50.84$\pm$ x.xx 46.84$\pm$ x.xx 18.61$\pm$ x.xx 161.8$\pm$ x.x / 160.4$\pm$ x.x 50.58$\pm$ x.xx / 46.87$\pm$ x.xx Diffusion 14.7 57.72$\pm$ 0.77 42.96$\pm$ 0.67 18.60$\pm$ 0.06 159.0$\pm$ 1.4 / 158.0$\pm$ 1.1 54.24$\pm$ 0.89 / 52.24$\pm$ 0.68 Amazon MostPopular 5.0 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 6.33$\pm$ 0.00 / 6.30$\pm$ 0.00 10.10$\pm$ 0.00 / 9.72$\pm$ 0.00 HistoryDensity 71.8 32.68$\pm$ 0.00 9.63$\pm$ 0.00 0.325$\pm$ 0.000 7.29$\pm$ 0.00 / 7.27$\pm$ 0.00 28.25$\pm$ 0.00 / 27.53$\pm$ 0.00 IFTPP 12.8 35.54$\pm$ 0.08 16.72$\pm$ 0.04 0.250$\pm$ 0.001 6.70$\pm$ 0.01 / 6.64$\pm$ 0.01 23.05$\pm$ 0.48 / 23.79$\pm$ 0.39 IFTPP-T 13.1 35.27$\pm$ 0.04 16.46$\pm$ 0.07 0.245$\pm$ 0.001 6.64$\pm$ 0.02 / 6.59$\pm$ 0.02 22.81$\pm$ 0.10 / 23.42$\pm$ 0.13 RMTPP 16.4 35.85$\pm$ 0.03 17.21$\pm$ 0.04 0.292$\pm$ 0.002 6.67$\pm$ 0.03 / 6.62$\pm$ 0.03 19.65$\pm$ 0.23 / 19.99$\pm$ 0.18 NHP 9.9 16.58$\pm$ 3.26 11.19$\pm$ 0.18 0.438$\pm$ 0.009 8.54$\pm$ 0.28 / 8.52$\pm$ 0.30 27.19$\pm$ 0.35 / 26.95$\pm$ 0.37 AttNHP 12.4 22.21$\pm$ 5.64 8.22$\pm$ 0.39 0.526$\pm$ 0.075 7.63$\pm$ 0.12 / 7.61$\pm$ 0.14 19.29$\pm$ 1.45 / 18.59$\pm$ 1.52 ODE 9.4 10.44$\pm$ 2.84 10.28$\pm$ 0.57 0.473$\pm$ 0.031 9.23$\pm$ 0.17 / 9.23$\pm$ 0.18 23.73$\pm$ 0.36 / 23.13$\pm$ 0.46 HYPRO 17.8 21.47$\pm$ x.xx 17.14$\pm$ x.xx 0.288$\pm$ x.xxx 6.95$\pm$ x.xx / 6.95$\pm$ x.xx 20.47$\pm$ x.xx / 20.28$\pm$ x.xx Diffusion 13.3 30.57$\pm$ 1.28 13.36$\pm$ 0.49 0.247$\pm$ 0.000 6.56$\pm$ 0.03 / 6.52$\pm$ 0.04 30.08$\pm$ 0.25 / 30.29$\pm$ 0.32 StackOverflow MostPopular 9.8 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.602$\pm$ 0.000 13.12$\pm$ 0.00 / 13.15$\pm$ 0.00 8.97$\pm$ 0.00 / 8.16$\pm$ 0.00 HistoryDensity 74.0 42.51$\pm$ 0.00 5.61$\pm$ 0.00 0.669$\pm$ 0.000 17.30$\pm$ 0.00 / 17.29$\pm$ 0.00 16.23$\pm$ 0.00 / 14.90$\pm$ 0.00 IFTPP 24.1 45.31$\pm$ 0.10 11.59$\pm$ 0.70 0.646$\pm$ 0.003 13.56$\pm$ 0.19 / 13.66$\pm$ 0.19 9.68$\pm$ 0.55 / 8.87$\pm$ 0.59 IFTPP-T 23.0 45.40$\pm$ 0.04 12.78$\pm$ 0.29 0.639$\pm$ 0.001 13.51$\pm$ 0.07 / 13.61$\pm$ 0.05 9.62$\pm$ 0.15 / 8.81$\pm$ 0.20 RMTPP 12.7 45.30$\pm$ 0.07 12.61$\pm$ 0.40 0.717$\pm$ 0.006 13.00$\pm$ 0.03 / 13.25$\pm$ 0.04 13.81$\pm$ 0.30 / 13.35$\pm$ 0.28 NHP 11.9 44.12$\pm$ 0.79 9.90$\pm$ 1.16 0.735$\pm$ 0.021 13.27$\pm$ 0.31 / 13.46$\pm$ 0.27 13.12$\pm$ 0.26 / 12.41$\pm$ 0.31 AttNHP 15.6 45.17$\pm$ 0.16 12.56$\pm$ 0.15 0.703$\pm$ 0.000 13.10$\pm$ 0.03 / 13.31$\pm$ 0.01 11.95$\pm$ 0.38 / 11.15$\pm$ 0.29 ODE 13.3 44.19$\pm$ 0.13 9.08$\pm$ 0.27 0.728$\pm$ 0.007 13.12$\pm$ 0.05 / 13.35$\pm$ 0.05 11.48$\pm$ 0.29 / 10.75$\pm$ 0.21 HYPRO 11.8 30.54$\pm$ 0.70 12.61$\pm$ 0.40 0.718$\pm$ 0.006 14.00$\pm$ 0.19 / 14.12$\pm$ 0.17 15.25$\pm$ 0.28 / 14.84$\pm$ 0.23 Diffusion 14.4 37.03$\pm$ 1.24 7.48$\pm$ 0.16 0.668$\pm$ 0.006 12.87$\pm$ 0.17 / 13.01$\pm$ 0.16 15.80$\pm$ 0.85 / 15.07$\pm$ 0.74
 Computational Efficiency.
@@
 Straightforward evaluation can take several hours on a single Nvidia V100 GPU.
-To accelerate computation, we optimized the training and inference pipelines in two ways.
+We optimized the training and inference pipelines in two ways to accelerate computation.
 First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions.
-Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE), achieving up to a 4x performance improvement compared to the official implementations.
+Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE) (Rizoiu et al., 2017), achieving up to a 4x performance improvement compared to the official implementations.
 These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV.
 Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
-A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s computational performance in Appendix C.5.
-Table 1: Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
-Method Mean length Next-item Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.86$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.06$\pm$ 0.00 / 0.99$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.87$\pm$ 0.00 0.924$\pm$ 0.000 7.40$\pm$ 0.00 / 7.44$\pm$ 0.00 2.46$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 11.0 34.08$\pm$ 0.04 3.47$\pm$ 0.01 0.693$\pm$ 0.000 6.88$\pm$ 0.01 / 6.90$\pm$ 0.01 5.82$\pm$ 0.27 / 5.88$\pm$ 0.13 IFTPP-K 10.1 33.69$\pm$ 0.03 3.25$\pm$ 0.01 0.698$\pm$ 0.001 7.18$\pm$ 0.00 / 7.19$\pm$ 0.00 4.42$\pm$ 0.15 / 4.43$\pm$ 0.16 RMTPP 7.5 34.15$\pm$ 0.07 3.47$\pm$ 0.02 0.749$\pm$ 0.005 6.86$\pm$ 0.01 / 6.88$\pm$ 0.01 7.08$\pm$ 0.16 / 6.69$\pm$ 0.12 RMTPP-K 7.3 33.63$\pm$ 0.07 3.24$\pm$ 0.02 0.749$\pm$ 0.001 7.10$\pm$ 0.00 / 7.11$\pm$ 0.00 5.82$\pm$ 0.05 / 5.52$\pm$ 0.13 NHP 9.4 35.43$\pm$ 0.07 3.41$\pm$ 0.01 0.696$\pm$ 0.002 6.97$\pm$ 0.01 / 6.98$\pm$ 0.01 5.59$\pm$ 0.07 / 5.61$\pm$ 0.05 AttNHP 7.6 31.12$\pm$ x.xx 1.21$\pm$ x.xx 0.717$\pm$ x.xxx 7.52$\pm$ x.xx / 7.50$\pm$ x.xx 1.71$\pm$ x.xx / 1.48$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 6.9 34.26$\pm$ x.xx 3.46$\pm$ x.xx 0.758$\pm$ x.xxx 7.04$\pm$ x.xx / 7.05$\pm$ x.xx 7.79$\pm$ x.xx / 7.05$\pm$ x.xx MIMIC-IV MostPopular 10.3 4.77$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.82$\pm$ 0.00 / 19.82$\pm$ 0.00 0.55$\pm$ 0.00 / 0.54$\pm$ 0.00 Last 5 4.0 1.02$\pm$ 0.00 2.63$\pm$ 0.00 5.34$\pm$ 0.00 19.75$\pm$ 0.00 / 19.73$\pm$ 0.00 2.41$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 12.7 58.59$\pm$ 0.02 47.32$\pm$ 1.39 3.00$\pm$ 0.01 11.51$\pm$ 0.01 / 11.53$\pm$ 0.01 21.93$\pm$ 0.21 / 21.67$\pm$ 0.21 IFTPP-K 15.1 57.91$\pm$ 0.13 44.60$\pm$ 0.14 3.07$\pm$ 0.02 13.17$\pm$ 0.05 / 13.18$\pm$ 0.05 22.46$\pm$ 0.03 / 22.30$\pm$ 0.03 RMTPP 10.0 58.33$\pm$ 0.02 46.24$\pm$ 1.18 3.89$\pm$ 0.03 13.64$\pm$ 0.03 / 13.71$\pm$ 0.03 21.49$\pm$ 0.31 / 21.08$\pm$ 0.29 RMTPP-K 8.4 57.48$\pm$ 0.06 43.47$\pm$ 0.11 3.62$\pm$ 0.02 14.68$\pm$ 0.01 / 14.72$\pm$ 0.02 20.70$\pm$ 0.16 / 20.39$\pm$ 0.14 NHP 6.1 24.97$\pm$ 0.94 11.12$\pm$ 0.49 6.53$\pm$ 0.77 18.59$\pm$ 0.18 / 18.60$\pm$ 0.19 7.26$\pm$ 1.35 / 7.32$\pm$ 1.33 AttNHP 5.1 42.75$\pm$ 0.52 28.55$\pm$ 1.06 3.08$\pm$ 0.04 14.66$\pm$ 0.08 / 14.68$\pm$ 0.08 22.64$\pm$ 0.41 / 22.46$\pm$ 0.40 ODE 12.5 43.21$\pm$ 2.30 25.34$\pm$ 3.05 2.93$\pm$ 0.03 14.71$\pm$ 0.34 / 14.74$\pm$ 0.34 15.41$\pm$ 0.21 / 15.18$\pm$ 0.15 HYPRO 7.4 58.35$\pm$ x.xx 45.45$\pm$ x.xx 3.95$\pm$ x.xx 14.82$\pm$ x.xx / 14.87$\pm$ x.xx 16.94$\pm$ x.xx / 16.77$\pm$ x.xx Retweet MostPopular 28.0 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 25.15$\pm$ 0.00 / 23.91$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 35.73$\pm$ 0.00 21.87$\pm$ 0.00 152.3$\pm$ 0.0 / 150.3$\pm$ 0.0 29.12$\pm$ 0.00 / 29.24$\pm$ 0.00 IFTPP 26.1 59.95$\pm$ 0.03 46.53$\pm$ 0.04 18.27$\pm$ 0.03 173.3$\pm$ 4.1 / 172.7$\pm$ 4.4 34.90$\pm$ 4.63 / 31.75$\pm$ 4.44 IFTPP-K 14.8 59.55$\pm$ 0.26 45.09$\pm$ 0.62 18.21$\pm$ 0.42 168.6$\pm$ 2.4 / 167.9$\pm$ 2.8 37.11$\pm$ 4.91 / 34.73$\pm$ 5.11 RMTPP 19.0 60.07$\pm$ 0.10 46.81$\pm$ 0.06 18.45$\pm$ 0.16 167.6$\pm$ 3.4 / 166.7$\pm$ 3.3 47.86$\pm$ 0.91 / 44.74$\pm$ 0.94 RMTPP-K 13.7 59.99$\pm$ 0.05 46.34$\pm$ 0.09 18.33$\pm$ 0.13 164.7$\pm$ 2.8 / 163.9$\pm$ 2.8 49.07$\pm$ 1.16 / 46.16$\pm$ 1.32 NHP 18.1 60.08$\pm$ 0.04 46.83$\pm$ 0.11 18.42$\pm$ 0.11 167.0$\pm$ 1.6 / 165.8$\pm$ 1.6 48.31$\pm$ 0.67 / 45.07$\pm$ 0.34 AttNHP 29.5 60.03$\pm$ 0.03 46.74$\pm$ 0.01 18.39$\pm$ 0.05 173.3$\pm$ 1.0 / 171.6$\pm$ 1.0 28.32$\pm$ 1.24 / 25.85$\pm$ 1.08 ODE 19.0 59.95$\pm$ 0.08 46.65$\pm$ 0.04 18.38$\pm$ 0.04 166.5$\pm$ 0.5 / 165.3$\pm$ 0.5 48.70$\pm$ 0.93 / 44.81$\pm$ 0.69 HYPRO 15.0 59.87$\pm$ x.xx 46.69$\pm$ x.xx 18.75$\pm$ x.xx 171.4$\pm$ x.x / 170.7$\pm$ x.x 49.90$\pm$ x.xx / 46.99$\pm$ x.xx Amazon MostPopular 20.6 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 8.59$\pm$ 0.00 / 8.31$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 8.15$\pm$ 0.00 0.321$\pm$ 0.000 6.43$\pm$ 0.00 / 6.41$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 IFTPP 13.7 35.73$\pm$ 0.12 17.14$\pm$ 0.10 0.242$\pm$ 0.003 6.58$\pm$ 0.06 / 6.52$\pm$ 0.05 21.94$\pm$ 0.46 / 22.56$\pm$ 0.52 IFTPP-K 13.3 35.11$\pm$ 0.08 16.48$\pm$ 0.03 0.246$\pm$ 0.000 6.72$\pm$ 0.01 / 6.68$\pm$ 0.01 22.06$\pm$ 0.05 / 22.57$\pm$ 0.07 RMTPP 16.7 35.76$\pm$ 0.06 17.21$\pm$ 0.02 0.294$\pm$ 0.001 6.62$\pm$ 0.02 / 6.57$\pm$ 0.03 19.70$\pm$ 0.31 / 20.06$\pm$ 0.33 RMTPP-K 15.8 35.06$\pm$ 0.14 16.37$\pm$ 0.20 0.300$\pm$ 0.005 6.92$\pm$ 0.01 / 6.87$\pm$ 0.02 17.85$\pm$ 0.29 / 18.12$\pm$ 0.30 NHP 9.9 11.06$\pm$ 3.43 11.22$\pm$ 0.19 0.449$\pm$ 0.014 9.04$\pm$ 0.31 / 9.02$\pm$ 0.35 26.24$\pm$ 0.36 / 26.29$\pm$ 0.55 AttNHP 18.5 31.83$\pm$ 0.32 9.70$\pm$ 0.28 0.461$\pm$ 0.003 7.32$\pm$ 0.06 / 7.30$\pm$ 0.06 14.50$\pm$ 0.58 / 14.62$\pm$ 0.80 ODE 9.6 7.54$\pm$ 0.95 10.14$\pm$ 0.23 0.492$\pm$ 0.018 9.48$\pm$ 0.11 / 9.46$\pm$ 0.08 23.54$\pm$ 0.62 / 22.96$\pm$ 0.61 HYPRO 18.0 35.69$\pm$ x.xx 17.21$\pm$ x.xx 0.295$\pm$ x.xxx 6.63$\pm$ x.xx / 6.61$\pm$ x.xx 20.58$\pm$ x.xx / 20.53$\pm$ x.xx StackOverflow MostPopular 14.0 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.10$\pm$ 0.00 / 5.56$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 5.20$\pm$ 0.00 0.934$\pm$ 0.000 14.52$\pm$ 0.00 / 14.55$\pm$ 0.00 8.67$\pm$ 0.00 / 6.72$\pm$ 0.00 IFTPP 24.1 45.41$\pm$ 0.11 13.00$\pm$ 0.79 0.641$\pm$ 0.002 13.57$\pm$ 0.07 / 13.64$\pm$ 0.05 8.78$\pm$ 0.87 / 8.31$\pm$ 0.50 IFTPP-K 15.3 44.85$\pm$ 0.24 11.16$\pm$ 1.07 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.51$\pm$ 0.06 12.42$\pm$ 0.97 / 11.42$\pm$ 0.78 RMTPP 14.5 45.43$\pm$ 0.13 13.33$\pm$ 0.25 0.701$\pm$ 0.007 12.95$\pm$ 0.02 / 13.17$\pm$ 0.05 13.26$\pm$ 0.29 / 12.72$\pm$ 0.16 RMTPP-K 12.0 44.89$\pm$ 0.09 11.72$\pm$ 0.10 0.689$\pm$ 0.001 12.92$\pm$ 0.01 / 13.13$\pm$ 0.01 14.91$\pm$ 0.19 / 14.30$\pm$ 0.09 NHP 13.0 44.53$\pm$ 0.05 10.86$\pm$ 0.19 0.715$\pm$ 0.004 13.02$\pm$ 0.02 / 13.24$\pm$ 0.02 12.67$\pm$ 0.55 / 11.96$\pm$ 0.40 AttNHP 15.5 45.17$\pm$ 0.10 12.67$\pm$ 0.13 0.705$\pm$ 0.001 13.08$\pm$ 0.03 / 13.30$\pm$ 0.02 11.95$\pm$ 0.19 / 11.13$\pm$ 0.32 ODE 13.9 44.38$\pm$ 0.09 10.12$\pm$ 0.19 0.711$\pm$ 0.004 13.04$\pm$ 0.02 / 13.27$\pm$ 0.03 11.37$\pm$ 0.32 / 10.52$\pm$ 0.23 HYPRO 11.8 45.18$\pm$ x.xx 12.88$\pm$ x.xx 0.715$\pm$ x.xxx 13.04$\pm$ x.xx / 13.26$\pm$ x.xx 15.57$\pm$ x.xx / 14.69$\pm$ x.xx
-6 Experiments
-We conduct experiments on five datasets of varying sizes and origins: the Transactions dataset (AI-Academy, 2021), the MIMIC-IV medical dataset (Johnson et al., 2023), two social networks datasets, Retweet (Zhao et al., 2015) and StackOverflow (Jure, 2014), and the Amazon reviews dataset (Jianmo, 2018).
-Dataset statistics and evaluation parameters are provided in Appendix C.4.
-We evaluate representative modeling methods from different groups:
-  1. 1.
-    Rule-based baselines.
-MostPopular Generates a constant output with the most popular label from the prefix and the average time step.
-The Last N baseline outputs N previous events with adjusted timestamps.
-  2. 2.
+A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s performance optimizations outlined in Appendix C.5.
+4.1. Datasets
+We for the first time combine datasets from a wide range of domains such as financial transactions, social networks, and medical records into a single evaluation benchmark.
+Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)).
+These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns and transactional data reflects daily activities, combining regularity with significant uncertainty.
+The dataset statistics are presented in Table 1.
+The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences.
+Retweet is medium size, whereas Amazon and StackOverflow can be considered small datasets.
+Further datasets details are reported in Appendix C.4 and Appendix F.
+4.2. Methods
+We implement representatives from different groups, listed below.
+Statistical baselines.
+MostPopular generates K future events with average time steps.
+The predicted labels approximate distribution of labels in historical data.
+The HistoryDistribution baseline resembles MostPopular, but estimates types densities w.r.t.
+time features.
 Intensity-free approaches.
-We implement IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024).
-  3. 3.
+We implement the IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024).
+IFTPP uses GRU reccurrent network as a backbone (Cho et al., 2014).
+We also evaluate IFTPP-T model with a transformer backbone (Radford et al., 2019).
 Intensity-based approaches.
 We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN.
-We add NHP (Mei & Eisner, 2017), based on a continuous time LSTM architecture.
-We also evaluate an AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019).
-  4. 4.
-    Next-K approaches.
-We adapt IFTPP and RMTPP to predict multiple future events directly, implementing IFTPP-K and RMTPP-K, respectively.
-These approaches originate in time series analysis and have not been previously applied in the MTPP domain.
-  5. 5.
-    Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses and selects the best sequence using a contrastive approach.
-Additional details on these methods are provided in Appendix A, with training specifics outlined in Appendix D.
-Hyperparameters for both the methods and metrics are presented in Appendix E.
-The main evaluation results are shown in Table 1.
-In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
-6.1 Metrics Comparison
-The evaluation results in Table 1 show that the OTD metric can yield low values even for simple rule-based baselines.
-For instance, the Last baseline achieves the lowest OTD values on the Retweet and Amazon datasets.
-In contrast, the T-mAP metric more clearly distinguishes between rule-based baselines and deep learning models.
-This difference arises from the ability to measure model confidence.
-OTD evaluates predicted labels alone, while T-mAP measures average precision by adjusting logits thresholds.
-Unlike rule-based baselines, deep models provide confidence scores, whereas baselines output only labels, resulting in low T-mAP scores for simple baselines.
-The HYPRO and NHP methods maximize T-mAP, while the lowest OTD values are achieved by statistical baselines, IFTPP, and RMTPP.
-The difference between the two metrics lies in balancing the importance of event order versus time prediction quality.
-OTD evaluates the first 5-10 events, and to minimize OTD, a model must correctly identify the events that constitute the beginning of the sequence.
-IFTPP and RMTPP jointly predict the next event’s time and label, leading to better ordering.
-In contrast, methods like NHP, AttNHP, and ODE predict time steps independently for each event type.
-When the differences between timestamps are small, these methods can result in random ordering, negatively affecting the OTD score.
-T-mAP, on the other hand, is less sensitive to event order as long as the time predictions are accurate.
-The HYPRO method, designed to differentiate between ground truth and predicted sequences using a discriminator model, achieves high T-mAP scores.
-This highlights T-mAP’s ability to reward models that generate realistic sequences.
-We also observed that methods with high T-mAP scores often have low next-event prediction accuracy.
-For instance, on the Retweet, Amazon, and MIMIC-IV datasets, high long-horizon performance is often accompanied by low next-event prediction scores.
-This is particularly true for the HYPRO method, whose loss function is heavily focused on long-horizon quality.
-The contrast on the MIMIC-IV dataset can be attributed to the large number of events with identical timestamps, making precise event ordering more difficult and impacting both next-event and OTD scores.
-6.2 Intensity-Based vs Intensity-Free
-Previous works mostly compared either intensity-based neural methods (Du et al., 2016; Xue et al., 2022) or intensity-free approaches (Padhi et al., 2021; McDermott et al., 2024).
-The only exception is IFTPP (Shchur et al., 2019), which did not compare with NHP and included only test set likelihood as an evaluation metric.
-This raises the question: which approach is superior?
-We compare the intensity-free IFTPP method with intensity-based RMTPP, NHP, and ODE approaches.
-Table 1 shows that the intensity-free IFTPP method excels in the next-event MAE prediction, as it optimizes this metric.
-In other scenarios, the results vary between datasets.
-Therefore, we conclude that there is no preferred solution, and both approaches warrant attention in future research.
-6.3 Autoregression vs Direct Horizon Prediction
-We compared the IFTPP and RMTPP methods with their Next-K counterparts.
-Our results indicate that autoregressive models generally perform better regarding next-event prediction quality, while Next-K approaches improve T-mAP scores on all datasets except Transactions.
-The OTD-based rankings vary across datasets, with no single approach consistently outperforming the others.
-Despite these findings, little effort has been made to adapt Next-K models to MTPPs, suggesting future research in this area.
-We also observed that the Next-K approaches exhibit some interesting properties compared to autoregression.
-For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Figure 5.
-The likely reason for this behavior is the dependency of the model’s future output on its past errors.
-Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict.
-This suggests that Next-K models have the potential to predict better the distributions of future labels, a factor that should be considered in future research.
-Figure 5: Entropy of label distributions as a function of the position in the generated sequence. Results for StackOverflow and Amazon datasets are provided in Appendix F.
-6.4 The Maximum Sequence Length
-We observed that long-horizon prediction quality largely depends on the maximum number of allowed predictions in autoregression and Next-K models.
-As shown in Fig.
-6, the optimal number of predicted events is usually less than the maximum number of events on the horizon.
-For example, in the Transactions dataset, the optimal length ranges from 3 to 5, depending on the method.
-This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to limit the number of predictions manually.
-Figure 6: T-mAP dependency on the maximum length of the predicted sequence. Results for the StackOverflow and Amazon datasets are detailed in Appendix G.
-Therefore, we conclude that greater attention should be given to probability calibration between generation steps.
-This finding also highlights progress in long-horizon prediction tasks, particularly determining the maximum horizon that can be accurately modeled.
-T-mAP is the only metric capable of evaluating the optimal predicted sequence length, while the OTD metric remains unaffected by sequence length as it compares fixed-size prefixes.
-7 Limitations and Future Work
+We add NHP (Mei and Eisner, 2017), based on a continuous time LSTM architecture.
+We also evaluate the AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019), which applies ordinary differential equations for temporal dynamics modeling.
+Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses with RMTPP and selects the best sequence using a contrastive approach.
+Diffusion.
+We include the Diffusion approach (Zhou et al., 2025), which extends previous research in language modeling (Li et al., 2022).
+Diffusion model iteratively refines predicted sequence, simultaneously generating all events within the specified horizon.
+Notably, HoTPP provides the first public implementation of this approach.
+We also explored ContiFormer (Chen et al., 2023) and NJDTPP (Zhang et al., 2024), but excluded them from our final comparison for the reasons detailed in Appendix A.1.
+Additional details on implemented methods are provided in Appendix A, with training specifics outlined in Appendix D and hyperparameters listed in Appendix E.
+5. Experiments
+The main evaluation results are presented in Table 2. In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
+5.1. Long-horizon Prediction Quality
+According to Table 2, autoregressive approaches such as IFTPP, RMTPP, and NHP excel at next-event prediction (MAE and accuracy), while horizon prediction methods like HYPRO and Diffusion perform better in terms of OTD and T-mAP.
+These findings indicate that next-event and horizon prediction are distinct tasks requiring specialized models, and none of the existing approaches achieve top performance in both domains.
+Our statistical baselines rank first in 3 out of 10 long-horizon prediction tasks, and in three additional tasks they perform on par with the top methods without a statistically significant difference.
+These findings highlight the need for greater focus on long-horizon forecasting and point to opportunities for future improvements.
+Intensity-based approaches such as RMTPP and NHP generally outperform the autoregressive, intensity-free strategy.
+Meanwhile, Diffusion relies on intensity-free losses, suggesting that integrating intensity-based loss functions into Diffusion could potentially enhance long-horizon prediction.
+Notably, Transformer-based methods rarely offer advantages over RNN architectures.
+For instance, IFTPP-T significantly surpasses the RNN-based IFTPP only on MIMIC-IV, while being outperformed in most other scenarios.
+Consequently, the potential of Transformer architectures in the MTPP domain remains insufficiently explored.
+Figure 5. Comparison of predicted label entropies, with the dashed horizontal line indicating the ground truth entropy.
+5.2. Metrics Comparison
+Table 2 shows that the OTD metric can yield low values even for simple rule-based baselines.
+For example, the MostPopular baseline achieves the lowest OTD scores on the Retweet and Amazon datasets.
+In contrast, T-mAP more clearly distinguishes between rule-based baselines and deep learning models by measuring predicted labels probabilities rather than hard labels.
+The difference between OTD and T-mAP also emerges when comparing autoregressive methods that use the NHP loss (NHP, ODE) with those that jointly model timestamp and label for the next event (IFTPP, RMTPP).
+For instance, NHP performs poorly on OTD yet ranks among the top autoregressive approaches by T-mAP.
+This discrepancy arises from how the two metrics balance event order and time prediction quality.
+OTD considers only the first 5–10 events, so the model must accurately predict these initial events.
+Because IFTPP and RMTPP jointly predict the next event’s time and label, they generally preserve ordering more effectively.
+In contrast, NHP, AttNHP, and ODE predict times separately for each event type, causing random ordering when timestamps are close and thereby lowering the OTD score.
+We conclude that T-mAP is less sensitive to event order as long as the predicted times are accurate.
+The HYPRO and Diffusion methods, specifically designed for long-horizon prediction, achieve high T-mAP scores but often underperform autoregressive approaches in terms of OTD.
+This highlights T-mAP’s ability to reward models that produce realistic sequences.
+We also observe that methods with high T-mAP scores frequently have low next-event prediction accuracy.
+For example, on the Retweet, Amazon, and StackOverflow datasets, high long-horizon performance often coincides with low next-event prediction scores.
+This is especially true for HYPRO and Diffusion, whose loss functions heavily emphasize long-horizon quality.
+Meanwhile, the strong correlation of next-event prediction quality and long-horizon performance on MIMIC-IV can be attributed to the large number of events with identical timestamps, which makes precise event ordering crucial and favors autoregressive approaches like IFTPP.
+Table 3. Examples of predicted sequences (labels only).
+Method Transactions MIMIC-IV Retweet Amazon StackOverflow IFTPP 6, 6, 6, 6, 6, 6 14, 12, 4, 6, 11, 10, 1, 3, 5, 8, 9, 16 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 11, 0, 11, 0, 0, 0, 0, 0, 0, 0, 0 7, 17, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 RMTPP 6, 6, 6, 6, 6, 6 14, 4, 6, 11 0, 0, 0, 0, 0, 0, 0 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11 17, 3, 3, 3, 3, 3 NHP 6, 6, 6, 6, 6 4, 6 1, 1, 1, 1, 1, 1, 1, 1 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 AttNHP 1, 1, 1, 1, 1 4, 6, 10, 1 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 3, 0, 4, 1, 1, 1, 0, 0, 0, 0, 0, 0 3, 3, 3, 3, 3, 3, 3 ODE 1, 6, 6, 6, 6, 6 4, 6, 1, 3, 5, 2, 7, 4, 6 0, 0, 0, 0, 0, 0, 0 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 HYPRO 3, 6, 6, 1, 6, 1, 6 24, 1, 3, 5, 17, 8, 2, 20 0, 0, 1, 0, 0, 1, 0 1, 11, 0, 0, 11, 11, 11, 5, 1, 0, 5, 11 17, 5, 3, 5, 13, 3, 5 Diffusion 6, 6, 6, 6, 6, 6, 6 14, 4, 6, 1, 1, 3, 3, 3, 8, 3, 2, 2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 0, 0, 8, 5, 11, 11, 5, 5, 0, 5, 5, 11 3, 3, 3, 3, 3, 3, 3, 3
+5.3. Mode Collapse
+We found that most models tend to predict a single label or repetitive patterns in long-horizon tasks, as shown in Table 3.
+To quantify this effect, we measured the entropies of predicted labels (Fig.5).
+Autoregressive methods generally exhibit low prediction diversity across all datasets except MIMIC-IV, where strong dependencies enable the model to learn typical sequences.
+Meanwhile, HYPRO, which samples from RMTPP with an increased temperature, substantially enhances label diversity.
+Diffusion similarly boosts diversity except on the Transactions dataset.
+Overall, most methods, except HYPRO, experience mode collapse in long-horizon prediction, although horizon-based approaches partly mitigate this issue.
+Sampling with an increased temperature can potentially solve this issue by offering a trade-off between predictions quality and diversity.
+Additional examples of predicted sequences are presented in Appendix G.
+6. Limitations and Future Work
 The proposed benchmark has several shortcomings.
@@
 Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training.
-Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating hybrid (Deshpande et al., 2021) and diffusion models (Zhou et al., 2023).
-Our benchmark encourages future research to develop improved techniques for predicting future events and establish simple baselines for better measuring progress in the field.
-For example, Next-K models, which predict multiple future events without autoregression, show promising results in long-horizon prediction tasks.
-We suggest further exploration of these models.
-We also highlight the importance of label distribution estimation and emphasize the need to improve confidence estimation and calibration.
-The T-mAP metric can be applied in domains beyond event sequences, like action recognition (Su & Hua, 2017).
+Additionally, the list of implemented methods can be extended, for example, by incorporating multi-task models (Deshpande et al., 2021).
+Our benchmark encourages future research to develop methods capable of solving next-event and long-horizon prediction tasks simultaneously, as discussed in Section 5.1.
+One promising direction involves adapting concepts from our statistical baselines for deep models, thereby surpassing baseline performance.
+A key question in this effort is the correct alignment between predicted and ground truth events during training, as discussed in the context of T-mAP (Section 3).
+Future work could also target the reduction of mode collapse and the enhancement of predictions diversity, as highlighted in Section 5.3.
+On the other hand, the T-mAP metric can be applied in domains beyond event sequences, such as action recognition (Su and Hua, 2017).
 T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW.
 T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error.
-Our theoretical justification of T-mAP can potentially be adapted to the mAP estimation algorithms in computer vision.
-8 Conclusion
+Our theoretical justification of T-mAP can potentially be adapted for mAP estimation algorithms in computer vision.
+7. Conclusion
 In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting.
-Our extensive evaluations using established datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction does not necessarily correlate with superior performance in horizon prediction.
-This finding underscores the need for benchmarks like HoTPP, which emphasize the importance of long-horizon accuracy and robustness.
-By shifting the focus from short-term to long-term predictive capabilities, HoTPP aims to drive the development of more sophisticated and reliable event sequence prediction models.
-This, in turn, has the potential to significantly enhance the practical applications of sequential event prediction in various domains, fostering innovation and improving decision-making processes across a wide range of industries.
-Appendix A Background on Marked Temporal Point Processes Modeling
+HoTPP provides optimized implementations of the state-of-the-art methods ranging from statistical baselines to ODE and Diffusion (Sec.4), along with diverse metrics for both next-event and long-horizon prediction (Sec.3).
+Our extensive evaluations using diverse datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction typically does not correlate with superior performance in horizon prediction (Sec.
+5.1).
+Moreover, most existing methods suffer from mode collapse when forecasting longer sequences (Sec.
+5.3).
+These findings underscore the need for developing specialized models.
+By emphasizing long-term predictive capabilities, HoTPP aims to stimulate the development of more robust and reliable event-sequence models.
+This advance has the potential to substantially enhance practical applications in various industries, driving both innovation and more informed decision-making.
+Appendix A Modeling Marked Temporal Point Processes
 Intensity-based approaches.
@@
 A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
-$\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$ (5)
+(5) $\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$
 where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function.
@@
 These approaches differ in the type of neural network used and the model of the intensity function.
-Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei & Eisner, 2017) and Neural ODEs (Rubanova et al., 2019).
+Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei and Eisner, 2017) and Neural ODEs (Rubanova et al., 2019).
 The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
 In our work, we apply HYPRO to the outputs of the RMTPP model.
-Next-K models.
-While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling (Lim & Zohren, 2021).
-These methods predict multiple future events at once, avoiding the need for autoregressive inference.
-The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression.
-The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length.
-However, applying a modified autoregressive approach can potentially address this limitation.
-In our work, we implement a Next-K variant of the IFTPP model, which is straightforward.
-Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences.
-The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1.
-Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events.
-Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
+Diffusion.
+Recently, diffusion-based approaches have been applied to NLP tasks (Li et al., 2022).
+This idea has been extended to event sequences by adapting best practices from NLP (Zhou et al., 2025).
+In such methods, events are first mapped to continuous embeddings, where the diffusion process operates, then mapped back to event space.
+The process begins with embeddings sampled from Gaussian noise and iteratively refines them using historical event embeddings.
+Finally, the refined embeddings are projected to event times and labels through a softmax-based layer.
+Training involves a combination of MSE loss in the continuous space, MAE loss for time reconstruction, and cross-entropy loss for label prediction.
+In HoTPP, we provide the first public implementation of the approach proposed by (Zhou et al., 2025).
+MostPopular Baseline.
+This baseline estimates the average inter-event time and the empirical label distribution from historical data.
+For next-event prediction, it outputs the average inter-event duration and selects the label with the highest probability.
+For long-horizon forecasting, it generates events at fixed time intervals and assigns labels such that the predicted distribution aligns with the estimated historical distribution, thus minimizing the discrepancy between predicted and observed label frequencies.
+Because it assigns hard labels, this method is particularly designed for next-event prediction and minimizing the OTD metric.
+HistoryDensity Baseline.
+Similar to MostPopular, this approach estimates label distributions from historical data.
+Unlike MostPopular, it focuses on label densities with respect to time rather than normalizing over the total number of events.
+Consequently, for long-horizon forecasting, it assigns soft probabilities to each predicted event, thereby maximizing the T-mAP score.
+A.1. Methods Excluded from Comparison
+We attempted to incorporate ContiFormer (Chen et al., 2023) into our benchmark but encountered multiple challenges.
+The official implementation provides scripts only for training on a toy spiral dataset, making it impossible to reproduce the method’s performance on real-world data.
+Moreover, training even on this toy example required approximately three hours on an NVIDIA RTX 4060, suggesting that ContiFormer cannot be efficiently scaled to large datasets such as Transactions.
+In addition, contrary to the claims in the paper, ODE (Rubanova et al., 2019) and continuous-time transformers (Yang et al., 2022) were already discussed in earlier work.
+Therefore, we excluded ContiFormer from our benchmark in favor of ODE and AttNHP.
+We also attempted to include NJDTPP (Zhang et al., 2024) in our benchmark, but encountered several challenges.
+Even on StackOverflow, our smallest dataset, the inference time spanned two hours for a single test set prediction.
+Moreover, after running the official implementation on StackOverflow, we observed accuracy values that were inconsistent with the paper and fell below most baseline results.
+We also identified data leakage during inference, where future event timestamps were used to predict the subsequent event.
+Consequently, we chose not to include NJDTPP in our comparisons.
 Appendix B T-mAP Computation and Proofs
@@
 Proof.
-Figure 7: Illustration of the iterative process in Lemma A.
-If the size of $m^{\prime}$ is greater than the size of matchings in $M(\mathcal{G})$ then the lemma holds.
+Figure 6. Illustration of the iterative process in Lemma A.
+If the size of $m^{\prime}$ is greater than the size of matchings in $M(\mathcal{G})$, then the lemma holds.
 The optimal matching in the extended graph $\mathcal{G}^{\prime}$ cannot have a size smaller than the optimal matching in $\mathcal{G}$.
@@
 If $m^{\prime}$ includes vertex $b$, then we can walk through the graph with the following process:
-  1. 1.
+  1. (1)
     Start from the new vertex $b$.
-  2. 2.
+  2. (2)
     If we are at a ground truth event $b_{i}$ and there is an edge $(a_{i},b_{i})$ in matching $m^{\prime}$, move to the vertex $a_{i}$.
-  3. 3.
+  3. (3)
     If we are at a prediction $a_{i}$ and there is an edge $(a_{i},b_{i})$ in matching $m$, move to the vertex $b_{i}$.
 Example processes are illustrated in Fig.
-7.
+6.
 The vertices in each process do not repeat; otherwise, there would be two edges in either $m$ or $m^{\prime}$ with the same vertex, which contradicts the definition of a matching.
@@
 If the process finishes at a ground truth event, we can replace a part of matching $m^{\prime}$ traced with a part of matching $m$.
-The resulting matching $\hat{m}$ will have a size equal to both $m$ and $m^{\prime}$ and will have a total weight equal to both matchings.
+The resulting matching $\hat{m}$ will have a size equal to $m$ and $m^{\prime}$ and a total weight equal to both matchings.
 This proves the lemma.
@@
 For any threshold $h$, there exists an optimal matching in the graph $\mathcal{G}_{h}$ such that it is a subset of an optimal matching in the full graph $\mathcal{G}$:
-$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (6)
+(6) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
 Proof.
@@
 Otherwise, we can construct a chain of thresholds $h_{1}<h_{2}<\dots<h$, with each subsequent threshold filtering an additional vertex, and apply the theorem iteratively.
-Denote $i$ the only vertex present in $\mathcal{G}$ and filtered from $\mathcal{G}_{h}$.
-If there is no edge containing vertex $i$ in the matching $m(\mathcal{G})$, then $m(\mathcal{G})$ is also the optimal matching in $\mathcal{G}_{h}$, and theorem holds.
+Denote $i$ to be the only vertex present in $\mathcal{G}$ and filtered from $\mathcal{G}_{h}$.
+If there is no edge containing vertex $i$ in the matching $m(\mathcal{G})$, then $m(\mathcal{G})$ is also the optimal matching in $\mathcal{G}_{h}$ and theorem holds.
 Consider the case when matching $m(\mathcal{G})$ contains an edge $(i,j)$.
 Let $\hat{m}_{h}=m(\mathcal{G})\setminus\{(i,j)\}$, i.e., the matching without edge $(i,j)$.
-If it is optimal for the graph $\mathcal{G}_{h}$, then it satisfies the theorem.
+If it is optimal for the graph $\mathcal{G}_{h}$, then it will satisfy the theorem.
 Otherwise, there is an optimal matching $m_{h}$ with either (a) more vertices or (b) a smaller total weight than $\hat{m}_{h}$.
 In (a), matching $m_{h}$ has a size greater than the size of $\hat{m}_{h}$.
-It follows, that $m_{h}$ has the maximum size in the full graph $\mathcal{G}$.
+It follows that $m_{h}$ has the maximum size in the complete graph $\mathcal{G}$.
 If the total weight of $m_{h}$ is larger or equal to the optimal weight in $\mathcal{G}$, then it can not be less than the total weight of $\hat{m}_{h}$, which contradicts our assumption.
-At the same time, the total weight of $m_{h}$ can not be less than the optimal weight in the full graph $\mathcal{G}$.
+At the same time, the total weight of $m_{h}$ can not be less than the optimal weight in the complete graph $\mathcal{G}$.
 It follows that case (a) leads to a contradiction.
@@
 By definition, multiclass T-mAP is the average of the average precision (AP) values for each label $l$:
-$\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$ (7)
+(7) $\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$
 Consider AP computation for a particular label $l$. AP is computed as the area under the precision-recall curve:
-$\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$ (8)
+(8) $\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$
 where $\mathrm{Rec}_{i}$ is the $i$-th recall value in a sorted sequence and $\mathrm{Prec}$ is the corresponding precision. Iteration is done among all distinct recall values ($\mathrm{Rec}_{0}=0$).
 We have several correct and incorrect predictions for each threshold $h$ on the predicted label probability.
-These quantities define precision and recall values along with the total number of ground truth events.
+These quantities define precision, recall, and the total number of ground truth events.
 A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|\leq\delta$.
@@
 Therefore, we define a matching, i.e., the correspondence between predictions and ground truth events.
-The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 4.2.
+The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 3.2.2.
 As the matching, without loss of generality, can be considered constant for different thresholds, we can split all predictions into two parts: those that were assigned a ground truth and those that were unmatched.
@@
 The resulting algorithm for AP computation for each label $l$ involves the following steps:
-  1. 1.
+  1. (1)
     Compute the optimal matching between predictions and ground truth events.
-  2. 2.
+  2. (2)
     Collect (a) scores of matched predictions, (b) scores of unmatched predictions, and (c) the number of unmatched ground truth events.
-  3. 3.
+  3. (3)
     Assign a positive label to matched predictions and a negative label to unmatched ones.
-  4. 4.
+  4. (4)
     Evaluate maximum recall as the fraction of matched ground truth events.
-  5. 5.
+  5. (5)
     Find the area under the precision-recall curve for the constructed binary classification problem from item 3 and multiply it by the maximum recall value.
@@
 Specifically, the optimal matching seeks to minimize the total weight in the following form:
-$C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$ (9)
+(9) $C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$
 A linear transformation of logits with a positive scaling factor will adjust the total weight accordingly, but the optimal matching will remain unchanged.
@@
 Appendix C HoTPP Benchmark Details
-C.1 Architecture
+C.1. Architecture
 HoTPP incorporates best practices from extensible and reproducible ML pipelines.
-It leverages PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments.
+It leverages PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments.
 Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility.
-The overall architecture is illustrated in Figure 8.
-HoTPP supports both discrete-time and continuous-time models, as well as RNN and Transformer architectures.
-Implementing a new method requires only the addition of essential components, while the rest can be configured through Hydra and configuration files.
-Figure 8: The architecture of the HoTPP library.
-C.2 Metrics
+The overall architecture is illustrated in Fig.
+7.
+HoTPP supports both discrete-time and continuous-time models as well as RNN and Transformer architectures.
+Implementing a new method requires only adding essential components, while the rest can be configured through Hydra and configuration files.
+Figure 7. The architecture of the HoTPP library.
+C.2. Metrics
 MTPP models are typically evaluated based on the accuracy of next-event predictions.
-Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts, and accuracy or error rate for label predictions.
-Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for IFTPP and Next-K models.
-Previous works have advanced long-horizon evaluation using optimal transport distance (OTD) (Mei et al., 2019; Xue et al., 2023).
-In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.
-C.3 Backbones
+Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label predictions.
+Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for the IFTPP and Diffusion models.
+Previous works have advanced long-horizon evaluation using OTD (Mei et al., 2019; Xue et al., 2023).
+In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.1.
+C.3. Backbones
 We use three types of architectures in our experiments.
-The first is the GRU network (Cho et al., 2014), which is one of the top-performing neural architectures for event sequences (Babaev et al., 2022).
-We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei & Eisner, 2017).
+The first is the GRU network (Cho et al., 2014), one of the top-performing neural architectures for event sequences (Babaev et al., 2022).
+We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei and Eisner, 2017).
 Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases.
@@
 Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
-C.4 Datasets
-For the first time, we combine domains such as financial transactions, social networks, and medical records into a single evaluation benchmark.
-Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)).
-These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns, and transactional data reflects daily activities, combining regularity with significant uncertainty.
-The dataset statistics are presented in Table 2.
-The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences.
-Retweet is of medium size, whereas Amazon and StackOverflow can be considered small datasets.
-Table 2: Datasets statistics
-Dataset Sequences Events Mean Length Mean Horizon Length Mean Duration Classes Transactions 50k 43.7M 875 9.0 719 203 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 Retweet 23k 1.3M 56.4 14.7 1805 3 Amazon 9k 403K 43.6 14.8 22.1 16 StackOverflow 2k 138K 64.2 12.0 55.3 22
+C.4. Datasets
 Transactions [^3] https://huggingface.co/datasets/dllllb/age-group-prediction, Retweet [^4] https://huggingface.co/datasets/easytpp/retweet, Amazon [^5] https://huggingface.co/datasets/easytpp/amazon, and StackOverflow [^6] https://huggingface.co/datasets/easytpp/stackoverflow datasets were obtained from the HuggingFace repository.
-Transactions data was released in competition and come with a free license [^7] https://www.kaggle.com/competitions/clients-age-group/data.
+Transactions data was released in competition and came with a free license [^7] https://www.kaggle.com/competitions/clients-age-group/data.
 Retweet, Amazon, and StackOverflow come with an Apache 2.0 license.
@@
 Additionally, we address the issue of multiple diagnoses occurring in a single event by sorting timestamps for reproducibility.
-C.5 Performance improvements
-The HoTPP benchmark provides highly optimized training and inference procedures for efficient evaluation on datasets containing up to tens of millions of events.
+C.5. Performance improvements
+The HoTPP benchmark provides highly optimized training and inference procedures for the efficient evaluation of datasets containing up to tens of millions of events.
 First, we implement parallel RNN inference, which reuses computations when inference is initiated from multiple starting points within a sequence.
@@
 Additionally, HoTPP offers a GPU implementation of the Hungarian algorithm, which also finds applications in computer vision.
-Table 3: HoTPP computation speed improvements in terms of seconds per batch.
+Table 4. HoTPP computation speed improvements in terms of seconds per batch.
 Implementation Study
@@
 HoTPP 0.14 0.0025 0.0276 0.01 0.048
+Impr. 17x 6x 2x 8x 3x
 Appendix D Training details
-We trained each model using the Adam optimizer with a learning rate of 0.001 and a scheduler that reduces the learning rate by 20% after every 5 epochs.
-Gradient clipping was applied, with a maximum L2-norm set to 1.
+We trained each model using the Adam optimizer, which has a learning rate 0.001, and a scheduler that reduces the learning rate by 20% after every five epochs.
+Gradient clipping was applied, and the maximum L2-norm was set to 1.
 We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060.
 Each method was trained on a single GPU.
-The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset, and up to 15 hours for NHP on Transactions.
+The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset and up to 15 hours for NHP on Transactions.
 Multi-seed evaluations took approximately five times longer to complete.
 Appendix E Hyperparameters
-We analyzed the behavior of OTD and T-mAP metrics depending on their parameters.
-The results are shown in Figure 9.
-The OTD cost has little effect on the ranking of methods but significantly influenced the metric scale.
-The choice of the $\delta$ parameter in T-mAP can, in some cases, affect method rankings, but as long as $\delta$ is not set too small or too large, the metric demonstrates stable rankings.
-We set the $\delta$ parameter to approximately 10-30% of the horizon duration.
-Evaluation hyperparameters are presented in Table 4. Dataset-specific training parameters are listed in Table 5.
-Table 4: Evaluation hyperparameters.
-Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 32 5 1 7 (week) 2 MIMIC-IV 32 5 2 28 (days) 4 Retweet 32 10 15 180 (seconds) 30 Amazon 32 5 1 10 (unk.
-unit) 2 StackOverflow 32 10 1 10 (minutes) 2
-Table 5: Training hyperparameters.
-Dataset Num epochs Max Seq.
-Len.
-Label Emb.
-Size Hidden Size Head hiddens Transactions 30 1200 256 512 512 $\rightarrow$ 256 MIMIC-IV 30 64 16 64 64 Retweet 30 264 16 64 64 Amazon 60 94 32 64 64 StackOverflow 60 101 32 64 64
-Figure 9: The dependency of metric values on the metric parameter.
-Appendix F Labels Entropy Degradation
-For simplicity, some datasets were omitted in the main part of the paper. In Figure 10, we show the dependency of label distribution entropy on step size for all datasets.
-Figure 10: Entropy of label distributions as a function of the position in the generated sequence.
-Appendix G The Optimal Sequence Length
-For simplicity, some datasets were omitted from the main part of the paper. Figure 11 illustrates the relationship between label distribution entropy and step size across all datasets.
-Figure 11: T-mAP dependency on the maximum length of the predicted sequence.
+Dataset-specific training parameters are listed in Table 5.
+Table 5. Training hyperparameters.
+Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens
+Transactions 30 1200 256 512 512 $\rightarrow$ 256
+MIMIC-IV 30 64 16 64 64
+Retweet 30 264 16 64 64
+Amazon 60 94 32 64 64
+StackOverflow 60 101 32 64 64
+Appendix F Datasets Analysis
+HoTPP incorporates datasets from various domains, including finance, social networks, and healthcare. Below, we provide additional details highlighting the key differences among these datasets.
+Ordering Sensitivity.
+While most datasets—such as StackOverflow, Amazon, Retweet, and Transactions—exhibit a relatively balanced distribution of timestamps, we observed that in MIMIC-IV, the proportion of zero time steps exceeds 50%, as shown in Table 6.
+This results in a significant number of events sharing the same timestamp.
+Consequently, the actual order of events within the dataset may hold greater significance during evaluation than the precise prediction of timestamps.
+This characteristic can lead to instability in timestamp-based metrics, such as OTD and T-mAP, compared to index-based metrics like next-event quality, pairwise MAE and accuracy.
+The low next-event accuracy of methods based on NHP loss (NHP, AttNHP, and ODE) on MIMIC-IV can be attributed to their independent modeling of each event class, which results in a random ordering of events with identical timestamps.
+Table 6. Time step percentiles.
+Dataset 1% 5% 10% 50% 90% 95% 99%
+Transactions 0.0 0.0 0.0 1.0 2.0 2.0 5.0
+MIMIC-IV 0.0 0.0 0.0 0.0 6.9 63.8 768.0
+Retweet 0.0 0.0 1.0 8.0 85.0 151.0 377.0
+Amazon 0.01 0.01 0.01 0.73 0.79 0.79 0.80
+StackOverflow 0.0003 0.01 0.05 0.51 2.16 2.97 5.08
+Discrete Timestamps.
+Some datasets feature continuous timestamps (e.g., StackOverflow, Amazon, MIMIC-IV), while others round timestamps to a specific precision (e.g., Retweet, Transactions).
+Modeling discrete timestamps presents a unique challenge, as density estimation methods like NHP can produce infinite PDF values.
+We introduce small Gaussian noise to discrete timestamps to address this issue, effectively smoothing the distribution.
+The degree of smoothing was carefully tuned for each dataset individually.
+The exact values are provided in the configuration files included with the source code.
+Appendix G Qualitative Analysis of Predictions
+Table 7 presents example predictions from various methods on the Transactions, MIMIC-IV, and Retweet datasets.
+The table focuses exclusively on predicted label sequences.
+It can be seen that all methods, except HYPRO, exhibit issues with constant or repetitive outputs.
+We believe this behavior stems from a bias in the predictions toward the most frequent labels, especially in scenarios with high uncertainty.
+For the MIMIC-IV dataset, the prediction patterns differ significantly, as most events follow a predefined sequence, such as admission, a standard set of laboratory tests, and diagnosis.
+In this case, the uncertainty is lower, enabling the methods to generate sequences with more diverse event types.
+Addressing the mentioned limitation in future research could lead to the development of methods capable of producing more varied and realistic sequences in high-uncertainty scenarios.
+Table 7. Example predictions (labels only).
+Method Seq. ID Transactions MIMIC-IV Retweet
+IFTPP 0 3, 1, 3, 1, 3, 1, 3, 3, 3 10, 27, 23, 22, 1, 27, 3, 28, 26, 25, 23, 30 1, 1
+1 6, 3, 6, 6, 6, 6, 6 11, 10 0, 0, 0, 0, 0, 0, 0, 0, 0
+2 3, 1, 3, 1, 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
+RMTPP 0 3, 1, 3, 1, 3, 1 10, 27, 23, 22, 1, 30, 27, 5, 8, 9, 16, 15 1, 1, 1, 1
+1 6, 6, 6, 6, 6, 6 11, 10, 1 0, 0, 0, 0, 0, 0, 0, 0
+2 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
+NHP 0 1, 1, 1, 1, 1, 3, 1, 1, 1, 1 6 1, 1, 1, 1
+1 1, 1, 6, 6, 6, 6, 6 6, 6, 6 1, 1, 1, 1, 1, 1, 1
+2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 6, 6 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
+ODE 0 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 10, 1, 3, 5, 2, 4, 6, 1 1, 1, 1, 1
+1 6, 6, 6, 6, 6 1, 3, 5, 2, 4, 6, 1, 3, 5, 2, 4, 6 0, 0, 0, 0, 0, 0, 0
+2 1, 1, 1, 1, 1, 1, 1 2, 4, 6, 1, 3, 5, 2, 4, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
+HYPRO 0 3, 1, 16, 3, 12, 1 23, 27, 22, 1, 28, 26, 25, 23 0, 0, 1, 0
+1 1, 32, 6, 6, 1, 6 1, 28, 25, 23, 3, 26, 22, 5 0, 0, 0, 1, 1, 0, 1
+2 3, 5, 1, 3, 5 15, 13, 2, 7, 4, 1, 3, 5 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0
+Diffusion 0 1, 3, 1, 1, 1, 3, 18, 1 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0
+1 6, 6, 6, 6, 6, 6, 6 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 0, 0, 1, 1, 1, 1, 0, 1, 0
+2 8, 1, 3, 8, 3, 8, 1, 3, 8, 1, 1, 8 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1
+Appendix H T-mAP for Highly Irregular Sequences
+This section analyzes a toy dataset containing highly irregular event sequences.
+This dataset includes a single label, which aims to predict timestamps.
+Most time intervals in the dataset are zero, with only 5% of intervals equal to one.
+We compare three simple baselines:
+  • ZeroStep, which predicts events with timestamps identical to the last observed event (zero intervals);
+  • UnitStep, which predicts events with a unit time step (the largest time step in the dataset);
+  • MeanStep predicts events using the average time step computed from historical data.
+Evaluation results are shown in Fig.
+8.
+The results indicate that the MAE and OTD metrics assign the lowest error to the ZeroStep baseline, which simply repeats the last event without accounting for the dataset’s irregularity.
+In contrast, T-mAP identifies the MeanStep baseline as the most effective, as it is the only method that analyzes historical data and incorporates timestamp statistics (mean interval) into its predictions.
+These findings suggest that T-mAP is a more appropriate metric for assessing the ability of methods to predict irregular events.
+Figure 8. Comparison of simple baselines on the Toy dataset with highly irregular time intervals. MAE and OTD metrics represent error values, while T-mAP measures model quality.
+Appendix I Long-tail Prediction
+In this section, we assess the capability of various evaluation metrics to capture long-tail prediction quality, specifically the ability of models to predict rare classes.
+Unlike OTD, T-mAP evaluates each class independently, allowing for different aggregation strategies.
+The standard T-mAP computes a macro average, where the quality for each event class contributes equally to the final score.
+Additionally, the HoTPP benchmark includes a weighted variant of T-mAP, where the weights are proportional to class frequencies.
+Fig.
+9 compares the performance of IFTPP and RMTPP on the Transactions dataset, which includes 203 classes.
+The results show that all metrics, except macro T-mAP, remain unaffected as the dataset size increases beyond 60 classes.
+In contrast, macro T-mAP effectively evaluates the ability of models to predict across all available classes, highlighting its suitability for long-tail prediction tasks.
+Figure 9. Comparison of various metrics on the Transactions dataset across subsets with fewer event classes.

### v4

@@
 HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?
-Abstract.
+Abstract
 Forecasting multiple future events within a given time horizon is essential for applications in finance, retail, social networks, and healthcare.
@@
 Finally, we analyze the impact of autoregression and intensity-based losses on prediction quality, and outline promising directions for future research.
-The HoTPP source code, hyperparameters, and full evaluation results are available on GitHub [^1] https://github.com/ivan-chai/hotpp-benchmark.
-copyright: acmlicensedjournalyear: 2025
-1. Introduction
+The HoTPP source code, hyperparameters, and full evaluation results are available at https://github.com/ivan-chai/hotpp-benchmark.
+1 Introduction
 Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information.
@@
 Event sequences differ fundamentally from other data types.
-Unlike tabular data (Wang and Sun, 2022), ESs include timestamps and possess an inherent order.
-In contrast to time series data (Lim and Zohren, 2021), ESs are characterized by irregular time intervals and additional data fields, such as event types.
+Unlike tabular data [29], ESs include timestamps and possess an inherent order.
+In contrast to time series data [16], ESs are characterized by irregular time intervals and additional data fields, such as event types.
 These differences necessitate the development of specialized models and evaluation practices.
-Irregularly spaced timestamps are commonly modeled via Temporal Point Processes (TPP) (Rizoiu et al., 2017).
+Irregularly spaced timestamps are commonly modeled via Temporal Point Processes (TPP) [24].
 Assigning categorical labels to these events leads to the Marked Temporal Point Process (MTPP) formalism.
-Since most event-sequence data can be viewed as MTPPs augmented with additional fields (McDermott et al., 2024), accurately modeling event times and labels remains a central challenge in the broader event-stream (ES) domain.
+Since most event-sequence data can be viewed as MTPPs augmented with additional fields [18], accurately modeling event times and labels remains a central challenge in the broader event-stream (ES) domain.
 In practice, a common question arises: what events will occur, and when, within a specific time horizon?
-Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks (Xue et al., 2022; Zhou et al., 2025).
+Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks [34, 40].
 For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions.
@@
 Another difficulty lies in evaluation.
-Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su and Hua, 2017).
-Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths (Mei et al., 2019; Xue et al., 2022, 2023).
+Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints [28].
+Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths [20, 34, 33].
 However, the limitations of this metric have not been previously considered.
-Figure 1.
-T-mAP evaluation pipeline.
+Figure 1: T-mAP evaluation pipeline.
 Unlike previous approaches, T-mAP evaluates sequences of variable lengths within the prediction horizon.
@@
 In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies.
+To this end, we present HoTPP, a novel benchmark illustrated in Figure 1.
 Our key contributions are summarized as follows:
-  1. (1)
-    We introduce HoTPP, the first open-source benchmark specifically designed for long-horizon event forecasting.
-By considering datasets and methods spanning finance, social networks, healthcare, and recommender systems, HoTPP significantly broadens the diversity and scale of data compared to previous benchmarks (Xue et al., 2023; McDermott et al., 2024; Xu, 2018).
+  1. 1.
+    We introduce the first open-source benchmark specifically designed for long-horizon event forecasting.
+By considering datasets and methods spanning finance, social networks, healthcare, and recommender systems, HoTPP significantly broadens the diversity and scale of data compared to previous benchmarks [33, 18, 32].
 Moreover, we provide optimized training and inference algorithms to support large-scale studies.
-  2. (2)
+  2. 2.
     We demonstrate that widely used evaluation metrics for MTPPs overlook critical aspects of model performance.
 In particular, they are sensitive to calibration and incorrectly count false positives and false negatives.
-To address these issues, we introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision (Lin et al., 2014).
+To address these issues, we introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision [17].
 Furthermore, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
-  3. (3)
+  3. 3.
     Guided by our methodology, we evaluate the forecasting performance of a diverse set of approaches.
@@
 Moreover, in 6 out of 10 long-horizon tasks, simple statistical baselines either outperform or match modern MTPP models, highlighting the need for more effective solutions.
-  4. (4)
+  4. 4.
     Finally, we quantify the entropy of predicted labels and demonstrate that most modern approaches tend to predict only the most frequent labels, suffering from the problem known as mode collapse.
 This issue severely restricts practical applicability and motivates further research into more accurate prediction strategies.
-2. Related Work
+2 Related Work
 Marked Temporal Point Processes.
-A Marked Temporal Point Process (MTPP) is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels (Rizoiu et al., 2017).
+MTPP is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels [24].
 Common approaches to MTPP modeling focus on predicting the next event.
@@
 Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures.
-Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei and Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024).
+Several works have employed classical RNNs [7, 31, 21] and transformers [41, 37], while others have proposed architectures with continuous time [19, 25, 14].
 In this work, we evaluate MTPP models with both discrete and continuous time architectures.
-Unlike previous benchmarks, we also assess simple rule-based baselines and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021).
-For more details on MTPP modeling, refer to Appendix A.
+Unlike previous benchmarks, we also assess simple rule-based baselines and popular methods from related fields, including GPT-like prediction models for event sequences [18, 22].
+For more details on MTPP modeling, refer to A.
 MTPP Evaluation.
 Previous benchmarks have primarily focused on medical data or traditional MTPP datasets.
-EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev and van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability.
-Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks.
-While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and horizon prediction approaches, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025).
+EventStream-GPT [18] and TemporAI [26] consider only medical data and do not implement methods from the MTPP field, despite their applicability.
+Early MTPP benchmarks, such as Tick [2] and PyHawkes [^1] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks.
+While PoPPy [32] and EasyTPP [33] include neural methods, they do not consider rule-based and horizon prediction approaches, such as HYPRO [34] and Diffusion [40].
 Furthermore, PoPPy does not evaluate long-horizon predictions at all.
 EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
-Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su and Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided.
+Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation [28], showing that DTW’s strict ordering constraints are impractical and should be avoided.
 This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW).
@@
 Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
-Figure 2.
-Comparison of the OTD and T-mAP metrics.
-In the example 1.a, OTD’s prefix-based evaluation of three events erroneously produces a false negative, whereas T-mAP in example 1.b compares all events within the specified time horizon for a more accurate assessment.
-Moreover, OTD evaluates only the label with the highest probability, treating cases 2.a and 2.b as equivalent.
-By contrast, T-mAP considers the entire distribution, yielding a low score (0.33) for 2.a and a perfect score (1) for 2.b.
-3. T-mAP Metric
+3 T-mAP Metric
 In this section, we show that widely used evaluation metrics for long-horizon prediction often miscount false positives (FP) and false negatives (FN).
@@
 In addition, T-mAP is invariant to linear calibration of model outputs, simplifying and stabilizing comparisons across different methods.
-Additional T-mAP experiments on sequences with highly irregular time steps are presented in Appendix H.
-Long-tail performance measurement is discussed in Appendix I.
-3.1. Limitations of the Next-Event and OTD Metrics
+A general overview of the T-mAP evaluation pipeline is presented in Fig.
+1.
+3.1 Limitations of the Next-Event and OTD Metrics
 MTPPs are typically evaluated based on the accuracy of next-event predictions, with time and type predictions assessed independently.
@@
 As a result, long-horizon evaluation metrics are necessary.
-Recent studies have advanced the measurement of horizon prediction quality by employing OTD (Mei et al., 2019).
+Figure 2: Comparison of the OTD and T-mAP metrics.
+In the example 1.a, OTD’s prefix-based evaluation of three events erroneously produces a false negative, whereas T-mAP in example 1.b compares all events within the specified time horizon for a more accurate assessment.
+Moreover, OTD evaluates only the label with the highest probability, treating cases 2.a and 2.b as equivalent.
+By contrast, T-mAP considers the entire distribution, yielding a low score (0.33) for 2.a and a perfect score (1) for 2.b.
+Recent studies have advanced the measurement of horizon prediction quality by employing OTD [20].
 This metric is analogous to the edit distance between the predicted event sequence and the ground truth.
@@
 OTD finds the minimum cost among all possible matchings:
-(1) $\mathrm{OTD}(S_{p},S_{gt})=\\ \min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$
+$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
 where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost.
-It is common to take $C_{ins}=C_{del}$ (Mei et al., 2019).
+It is common to take $C_{ins}=C_{del}$ [20].
 It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
@@
 Therefore, evaluating a dynamic number of events is essential to align with the ground truth’s time horizon properly.
-Figure 3. T-mAP computation pipeline.
 Metrics for MTPP evaluation can be categorized into two groups: those relying on event indices and those that measure time and label prediction quality independently of event sequential positions.
@@
 Therefore, we aim to develop an evaluation metric that accurately captures performance across common and rare classes.
-3.2. Temporal mAP: Definition and Computation
+3.2 Temporal mAP: Definition and Computation
 In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration.
-T-mAP is inspired by object detection methods from computer vision (Everingham et al., 2010), as illustrated in Fig.
-1.
+T-mAP is inspired by object detection methods from computer vision [8].
 Object detection aims to localize objects within an image and identify their types.
@@
 These adjustments result in significant differences in formulation and computation, detailed below.
-3.2.1. Definition
+3.2.1 Definition
 T-mAP is parameterized by the horizon length $T$ and the maximum allowed time error $\delta$.
@@
 For any threshold value $h$, we can select a subset of the predicted sequence $S^{l}_{p}$ with scores exceeding the threshold:
-(2) $S^{l}_{>}(h)=\{t^{p}_{i}:\exists(t^{p}_{i},s^{p}_{i})\in S^{l}_{p},s^{p}_{i}>h\}.$
+$S^{l}_{>}(h)=\{t^{p}_{i}:\exists(t^{p}_{i},s^{p}_{i})\in S^{l}_{p},s^{p}_{i}>h\}.$ (2)
 By definition, a predicted event $i$ can be matched with a ground truth event $j$ iff $|t^{p}_{i}-t^{gt}_{j}|\leq\delta$, meaning the time difference between the predicted and ground truth events is less than or equal to $\delta$.
@@
 Finally, T-mAP is defined as the average AP across all classes.
-Figure 4. T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value.
-3.2.2. Computation
+3.2.2 Computation
 Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP.
-This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker and Volgenant, 1988).
+This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm [12].
 The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
+Figure 3: T-mAP computation pipeline.
 For each pair of sequences $S^{l}_{p}$ and $S^{l}_{gt}$ we define a weighted bipartite graph $\mathcal{G}(S^{l}_{p},S^{l}_{gt})$ with $|S^{l}_{p}|$ vertices in the first and $|S^{l}_{gt}|$ vertices in the second part.
@@
 For any threshold $h$ there exists an optimal matching in the graph $\mathcal{G}_{h}$, that is a subset of an optimal matching in the full graph $\mathcal{G}$:
-(3) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
+$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (3)
 According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig.
 3.c.
-The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in Appendix B.
-3.2.3. T-mAP Hyperparameters
+The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in B.
+3.2.3 T-mAP Hyperparameters
 T-mAP has two hyperparameters: the maximum allowed time delta $\delta$ and evaluation horizon $H$.
@@
 As shown, the chosen $\delta$ parameter is positioned to the right of the initial slope of the parameter-quality curve, ensuring an optimal balance between time granularity and task difficulty.
-4. HoTPP Benchmark
+Figure 4: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value. The results for all 5 datasets are presented in C.
+3.2.4 Additional T-mAP Studies
+We analyze simple synthetic datasets to further investigate the properties of the proposed T-mAP metric.
+In D, we demonstrate that T-mAP more effectively distinguishes trivial baselines from statistical methods in scenarios characterized by highly irregular time steps.
+Additionally, the results in E show that, unlike OTD, T-mAP captures model performance on rare (long-tail) classes.
+4 HoTPP Benchmark
 The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox.
 Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction.
-HoTPP also differs from prior MTPP benchmarks by including simple rule-based baselines and long-horizon prediction methods that simultaneously predict multiple future events, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025).
+HoTPP also differs from prior MTPP benchmarks by including simple rule-based baselines and long-horizon prediction methods that simultaneously predict multiple future events, such as HYPRO [34] and Diffusion [40].
 The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
@@
 New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module.
-Core components are reusable through the Hydra configuration library (Yadan, 2019).
+Core components are reusable through the Hydra configuration library [35].
 For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, or ODE.
 Furthermore, each method’s backbone architecture can be easily switched between RNNs and Transformers.
-Methods such as IFTPP, RMTPP, and Diffusion are compatible with most HuggingFace models (Wolf et al., 2020).
+Methods such as IFTPP, RMTPP, and Diffusion are compatible with most HuggingFace models [30].
 Evaluation Stability.
@@
 This possesses a unique challenge in the long-horizon forecasting task.
-Previous long-horizon evaluation pipelines (Xue et al., 2022, 2023) made predictions only at the end of each sequence, resulting in a limited number of measurements and a low evaluation stability.
+Previous long-horizon evaluation pipelines [34, 33] made predictions only at the end of each sequence, resulting in a limited number of measurements and a low evaluation stability.
 To address this limitation, we evaluate long-horizon predictions at multiple intermediate points, decreasing the variance of evaluation metrics.
-Table 1. Datasets statistics
-Dataset # Sequences # Events Mean Length Mean Horizon Mean Duration # Classes OTD T-mAP Length Steps Cost $\delta$ Horizon Transactions 50k 43.7M 875 9.0 719 203 5 1 2 7 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 5 2 4 28 Retweet 23k 1.3M 56.4 14.7 1805 3 10 15 30 180 Amazon 9k 403K 43.6 14.8 22.1 16 5 1 2 10 StackOverflow 2k 138K 64.2 12.0 55.3 22 10 1 2 10
 Reproducibility.
 The HoTPP benchmark ensures reproducibility at multiple levels.
-First, we use the PytorchLightning library (Falcon and The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results.
+First, we use the PytorchLightning library [9] for training with a specified random seed and report multi-seed evaluation results.
 Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly.
 Finally, we specify the environment in a Dockerfile.
-Table 2. Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
-Method Mean length Next-event Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 5.0 32.69$\pm$ 0.00 1.09$\pm$ 0.00 0.649$\pm$ 0.000 7.03$\pm$ 0.00 / 7.05$\pm$ 0.00 2.86$\pm$ 0.00 / 3.10$\pm$ 0.00 HistoryDensity 154.2 32.78$\pm$ 0.00 0.85$\pm$ 0.00 1.063$\pm$ 0.000 7.42$\pm$ 0.00 / 7.44$\pm$ 0.00 7.92$\pm$ 0.00 / 7.51$\pm$ 0.00 IFTPP 13.3 38.27$\pm$ 0.03 4.65$\pm$ 0.03 0.637$\pm$ 0.001 6.79$\pm$ 0.01 / 6.81$\pm$ 0.01 6.20$\pm$ 0.10 / 6.08$\pm$ 0.14 IFTPP-T 12.8 37.68$\pm$ 0.03 4.30$\pm$ 0.03 0.640$\pm$ 0.001 6.84$\pm$ 0.02 / 6.85$\pm$ 0.02 5.80$\pm$ 0.27 / 5.55$\pm$ 0.19 RMTPP 8.7 38.28$\pm$ 0.04 4.64$\pm$ 0.02 0.693$\pm$ 0.005 6.68$\pm$ 0.01 / 6.70$\pm$ 0.01 7.41$\pm$ 0.35 / 7.12$\pm$ 0.10 NHP 9.4 35.47$\pm$ 0.12 3.36$\pm$ 0.02 0.693$\pm$ 0.005 6.99$\pm$ 0.03 / 6.99$\pm$ 0.03 5.66$\pm$ 0.17 / 5.48$\pm$ 0.12 AttNHP 6.8 30.68$\pm$ x.xx 1.09$\pm$ x.xx 0.749$\pm$ x.xxx 7.47$\pm$ x.xx / 7.48$\pm$ x.xx 1.50$\pm$ x.xx / 1.49$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 7.3 22.50$\pm$ x.xx 4.62$\pm$ x.xx 0.694$\pm$ x.xxx 7.62$\pm$ x.xx / 7.62$\pm$ x.xx 8.20$\pm$ x.xx / 7.54$\pm$ x.xx Diffusion 10.4 36.05$\pm$ 0.07 3.66$\pm$ 0.03 0.650$\pm$ 0.002 6.86$\pm$ 0.01 / 6.88$\pm$ 0.00 5.95$\pm$ 0.05 / 6.04$\pm$ 0.08 MIMIC-IV MostPopular 2.7 4.80$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.75$\pm$ 0.00 / 19.75$\pm$ 0.00 1.63$\pm$ 0.00 / 1.56$\pm$ 0.00 HistoryDensity 160.0 0.74$\pm$ 0.00 2.63$\pm$ 0.00 3.23$\pm$ 0.00 18.94$\pm$ 0.00 / 18.92$\pm$ 0.00 4.25$\pm$ 0.00 / 4.19$\pm$ 0.00 IFTPP 12.4 58.52$\pm$ 0.06 46.04$\pm$ 0.26 3.00$\pm$ 0.02 11.53$\pm$ 0.03 / 11.55$\pm$ 0.03 22.56$\pm$ 0.13 / 22.31$\pm$ 0.15 IFTPP-T 11.7 58.57$\pm$ 0.04 47.96$\pm$ 1.51 3.03$\pm$ 0.01 11.46$\pm$ 0.02 / 11.48$\pm$ 0.02 24.07$\pm$ 0.31 / 23.68$\pm$ 0.29 RMTPP 9.8 58.24$\pm$ 0.14 45.62$\pm$ 1.28 3.93$\pm$ 0.03 13.69$\pm$ 0.05 / 13.76$\pm$ 0.06 22.00$\pm$ 0.28 / 21.62$\pm$ 0.29 NHP 3.0 23.19$\pm$ 0.61 11.67$\pm$ 0.39 4.71$\pm$ 0.31 17.92$\pm$ 0.34 / 17.92$\pm$ 0.33 9.76$\pm$ 0.41 / 9.80$\pm$ 0.38 AttNHP 6.2 44.99$\pm$ 1.11 31.88$\pm$ 0.64 2.95$\pm$ 0.20 14.30$\pm$ 0.16 / 14.31$\pm$ 0.16 19.98$\pm$ 0.79 / 19.74$\pm$ 0.71 ODE 11.6 44.16$\pm$ 0.76 27.48$\pm$ 1.04 2.91$\pm$ 0.01 14.36$\pm$ 0.20 / 14.39$\pm$ 0.20 16.37$\pm$ 0.75 / 16.19$\pm$ 0.74 HYPRO 7.4 45.69$\pm$ x.xx 45.54$\pm$ x.xx 3.98$\pm$ x.xx 14.68$\pm$ x.xx / 14.75$\pm$ x.xx 16.41$\pm$ x.xx / 16.13$\pm$ x.xx Diffusion 15.4 46.22$\pm$ 5.31 33.67$\pm$ 3.66 2.99$\pm$ 0.01 13.25$\pm$ 0.15 / 13.28$\pm$ 0.14 22.90$\pm$ 0.10 / 22.82$\pm$ 0.12 Retweet MostPopular 9.9 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 141.7$\pm$ 0.0 / 139.1$\pm$ 0.0 29.72$\pm$ 0.00 / 30.02$\pm$ 0.00 HistoryDensity 14.9 58.62$\pm$ 0.00 39.86$\pm$ 0.00 20.68$\pm$ 0.00 178.5$\pm$ 0.0 / 177.0$\pm$ 0.0 44.48$\pm$ 0.00 / 44.60$\pm$ 0.00 IFTPP 23.1 59.75$\pm$ 0.25 46.23$\pm$ 0.31 18.27$\pm$ 0.05 165.8$\pm$ 0.2 / 164.8$\pm$ 0.3 43.93$\pm$ 0.69 / 40.59$\pm$ 0.84 IFTPP-T 23.2 59.88$\pm$ 0.06 46.07$\pm$ 0.11 18.47$\pm$ 0.06 167.4$\pm$ 1.3 / 166.1$\pm$ 1.1 43.73$\pm$ 1.88 / 40.05$\pm$ 1.73 RMTPP 16.6 60.01$\pm$ 0.06 46.81$\pm$ 0.03 18.60$\pm$ 0.04 169.2$\pm$ 1.0 / 168.0$\pm$ 1.0 49.03$\pm$ 0.15 / 45.60$\pm$ 0.28 NHP 18.1 60.05$\pm$ 0.08 46.84$\pm$ 0.13 18.39$\pm$ 0.12 167.0$\pm$ 1.7 / 165.7$\pm$ 1.7 49.25$\pm$ 0.37 / 45.64$\pm$ 0.26 AttNHP 29.8 60.06$\pm$ 0.03 46.66$\pm$ 0.02 18.32$\pm$ 0.02 174.7$\pm$ 0.7 / 172.9$\pm$ 0.7 26.74$\pm$ 1.16 / 24.24$\pm$ 0.94 ODE 17.7 59.93$\pm$ 0.06 46.36$\pm$ 0.30 18.53$\pm$ 0.09 168.2$\pm$ 1.1 / 166.7$\pm$ 1.0 48.95$\pm$ 0.30 / 45.45$\pm$ 0.38 HYPRO 16.2 50.84$\pm$ x.xx 46.84$\pm$ x.xx 18.61$\pm$ x.xx 161.8$\pm$ x.x / 160.4$\pm$ x.x 50.58$\pm$ x.xx / 46.87$\pm$ x.xx Diffusion 14.7 57.72$\pm$ 0.77 42.96$\pm$ 0.67 18.60$\pm$ 0.06 159.0$\pm$ 1.4 / 158.0$\pm$ 1.1 54.24$\pm$ 0.89 / 52.24$\pm$ 0.68 Amazon MostPopular 5.0 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 6.33$\pm$ 0.00 / 6.30$\pm$ 0.00 10.10$\pm$ 0.00 / 9.72$\pm$ 0.00 HistoryDensity 71.8 32.68$\pm$ 0.00 9.63$\pm$ 0.00 0.325$\pm$ 0.000 7.29$\pm$ 0.00 / 7.27$\pm$ 0.00 28.25$\pm$ 0.00 / 27.53$\pm$ 0.00 IFTPP 12.8 35.54$\pm$ 0.08 16.72$\pm$ 0.04 0.250$\pm$ 0.001 6.70$\pm$ 0.01 / 6.64$\pm$ 0.01 23.05$\pm$ 0.48 / 23.79$\pm$ 0.39 IFTPP-T 13.1 35.27$\pm$ 0.04 16.46$\pm$ 0.07 0.245$\pm$ 0.001 6.64$\pm$ 0.02 / 6.59$\pm$ 0.02 22.81$\pm$ 0.10 / 23.42$\pm$ 0.13 RMTPP 16.4 35.85$\pm$ 0.03 17.21$\pm$ 0.04 0.292$\pm$ 0.002 6.67$\pm$ 0.03 / 6.62$\pm$ 0.03 19.65$\pm$ 0.23 / 19.99$\pm$ 0.18 NHP 9.9 16.58$\pm$ 3.26 11.19$\pm$ 0.18 0.438$\pm$ 0.009 8.54$\pm$ 0.28 / 8.52$\pm$ 0.30 27.19$\pm$ 0.35 / 26.95$\pm$ 0.37 AttNHP 12.4 22.21$\pm$ 5.64 8.22$\pm$ 0.39 0.526$\pm$ 0.075 7.63$\pm$ 0.12 / 7.61$\pm$ 0.14 19.29$\pm$ 1.45 / 18.59$\pm$ 1.52 ODE 9.4 10.44$\pm$ 2.84 10.28$\pm$ 0.57 0.473$\pm$ 0.031 9.23$\pm$ 0.17 / 9.23$\pm$ 0.18 23.73$\pm$ 0.36 / 23.13$\pm$ 0.46 HYPRO 17.8 21.47$\pm$ x.xx 17.14$\pm$ x.xx 0.288$\pm$ x.xxx 6.95$\pm$ x.xx / 6.95$\pm$ x.xx 20.47$\pm$ x.xx / 20.28$\pm$ x.xx Diffusion 13.3 30.57$\pm$ 1.28 13.36$\pm$ 0.49 0.247$\pm$ 0.000 6.56$\pm$ 0.03 / 6.52$\pm$ 0.04 30.08$\pm$ 0.25 / 30.29$\pm$ 0.32 StackOverflow MostPopular 9.8 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.602$\pm$ 0.000 13.12$\pm$ 0.00 / 13.15$\pm$ 0.00 8.97$\pm$ 0.00 / 8.16$\pm$ 0.00 HistoryDensity 74.0 42.51$\pm$ 0.00 5.61$\pm$ 0.00 0.669$\pm$ 0.000 17.30$\pm$ 0.00 / 17.29$\pm$ 0.00 16.23$\pm$ 0.00 / 14.90$\pm$ 0.00 IFTPP 24.1 45.31$\pm$ 0.10 11.59$\pm$ 0.70 0.646$\pm$ 0.003 13.56$\pm$ 0.19 / 13.66$\pm$ 0.19 9.68$\pm$ 0.55 / 8.87$\pm$ 0.59 IFTPP-T 23.0 45.40$\pm$ 0.04 12.78$\pm$ 0.29 0.639$\pm$ 0.001 13.51$\pm$ 0.07 / 13.61$\pm$ 0.05 9.62$\pm$ 0.15 / 8.81$\pm$ 0.20 RMTPP 12.7 45.30$\pm$ 0.07 12.61$\pm$ 0.40 0.717$\pm$ 0.006 13.00$\pm$ 0.03 / 13.25$\pm$ 0.04 13.81$\pm$ 0.30 / 13.35$\pm$ 0.28 NHP 11.9 44.12$\pm$ 0.79 9.90$\pm$ 1.16 0.735$\pm$ 0.021 13.27$\pm$ 0.31 / 13.46$\pm$ 0.27 13.12$\pm$ 0.26 / 12.41$\pm$ 0.31 AttNHP 15.6 45.17$\pm$ 0.16 12.56$\pm$ 0.15 0.703$\pm$ 0.000 13.10$\pm$ 0.03 / 13.31$\pm$ 0.01 11.95$\pm$ 0.38 / 11.15$\pm$ 0.29 ODE 13.3 44.19$\pm$ 0.13 9.08$\pm$ 0.27 0.728$\pm$ 0.007 13.12$\pm$ 0.05 / 13.35$\pm$ 0.05 11.48$\pm$ 0.29 / 10.75$\pm$ 0.21 HYPRO 11.8 30.54$\pm$ 0.70 12.61$\pm$ 0.40 0.718$\pm$ 0.006 14.00$\pm$ 0.19 / 14.12$\pm$ 0.17 15.25$\pm$ 0.28 / 14.84$\pm$ 0.23 Diffusion 14.4 37.03$\pm$ 1.24 7.48$\pm$ 0.16 0.668$\pm$ 0.006 12.87$\pm$ 0.17 / 13.01$\pm$ 0.16 15.80$\pm$ 0.85 / 15.07$\pm$ 0.74
 Computational Efficiency.
@@
 First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions.
-Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE) (Rizoiu et al., 2017), achieving up to a 4x performance improvement compared to the official implementations.
+Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE) [24], achieving up to a 4x performance improvement compared to the official implementations.
 These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV.
 Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
-A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s performance optimizations outlined in Appendix C.5.
-4.1. Datasets
+A detailed description of the benchmark can be found in F, with further details on HoTPP’s performance optimizations outlined in G.
+Table 1: Datasets statistics
+Dataset # Sequences # Events Mean Length Mean Horizon # Classes OTD T-mAP Length Steps Cost $\delta$ Horizon Transactions 50k 43.7M 875 9.0 203 5 1 2 7 MIMIC-IV 120k 2.4M 19.7 6.6 34 5 2 4 28 Retweet 23k 1.3M 56.4 14.7 3 10 15 30 180 Amazon 9k 403K 43.6 14.8 16 5 1 2 10 StackOverflow 2k 138K 64.2 12.0 22 10 1 2 10
+4.1 Datasets
 We for the first time combine datasets from a wide range of domains such as financial transactions, social networks, and medical records into a single evaluation benchmark.
-Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)).
-These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns and transactional data reflects daily activities, combining regularity with significant uncertainty.
+Specifically, we provide evaluation results on a transactional dataset [1], the MIMIC-IV medical dataset [11], and social network datasets (Retweet [39], StackOverflow [13], and Amazon [10]).
+These datasets represent diverse underlying processes: social network data is influenced by cascades [39], medical records exhibit repetitive patterns and transactional data reflects daily activities, combining regularity with significant uncertainty.
 The dataset statistics are presented in Table 1.
@@
 Retweet is medium size, whereas Amazon and StackOverflow can be considered small datasets.
-Further datasets details are reported in Appendix C.4 and Appendix F.
-4.2. Methods
+Further datasets details are reported in F and H.
+4.2 Methods
 We implement representatives from different groups, listed below.
@@
   • Intensity-free approaches.
-We implement the IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024).
-IFTPP uses GRU reccurrent network as a backbone (Cho et al., 2014).
-We also evaluate IFTPP-T model with a transformer backbone (Radford et al., 2019).
+We implement the IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels [27, 22, 18].
+IFTPP uses GRU reccurrent network as a backbone [5].
+We also evaluate IFTPP-T model with a transformer backbone [23].
   • Intensity-based approaches.
-We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN.
-We add NHP (Mei and Eisner, 2017), based on a continuous time LSTM architecture.
-We also evaluate the AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019), which applies ordinary differential equations for temporal dynamics modeling.
-Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses with RMTPP and selects the best sequence using a contrastive approach.
+We implement RMTPP [7] as an example of the TPP approach with a traditional RNN.
+We add NHP [19], based on a continuous time LSTM architecture.
+We also evaluate the AttNHP approach that utilizes a continuous time transformer model [36] and ODE [25], which applies ordinary differential equations for temporal dynamics modeling.
+  • Reranking. We implement HYPRO [34], which generates multiple hypotheses with RMTPP and selects the best sequence using a contrastive approach.
   • Diffusion.
-We include the Diffusion approach (Zhou et al., 2025), which extends previous research in language modeling (Li et al., 2022).
+We include the Diffusion approach [40], which extends previous research in language modeling [15].
 Diffusion model iteratively refines predicted sequence, simultaneously generating all events within the specified horizon.
 Notably, HoTPP provides the first public implementation of this approach.
-We also explored ContiFormer (Chen et al., 2023) and NJDTPP (Zhang et al., 2024), but excluded them from our final comparison for the reasons detailed in Appendix A.1.
-Additional details on implemented methods are provided in Appendix A, with training specifics outlined in Appendix D and hyperparameters listed in Appendix E.
-5. Experiments
+We also explored ContiFormer [4] and NJDTPP [38], but excluded them from our final comparison for the reasons detailed in F.
+Additional details on implemented methods are provided in A, with training specifics outlined in I and hyperparameters listed in J.
+Table 2: Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
+Method Mean length Next-event Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 5.0 32.69$\pm$ 0.00 1.09$\pm$ 0.00 0.649$\pm$ 0.000 7.03$\pm$ 0.00 / 7.05$\pm$ 0.00 2.86$\pm$ 0.00 / 3.10$\pm$ 0.00 HistoryDensity 154.2 32.78$\pm$ 0.00 0.85$\pm$ 0.00 1.063$\pm$ 0.000 7.42$\pm$ 0.00 / 7.44$\pm$ 0.00 7.92$\pm$ 0.00 / 7.51$\pm$ 0.00 IFTPP 13.3 38.27$\pm$ 0.03 4.65$\pm$ 0.03 0.637$\pm$ 0.001 6.79$\pm$ 0.01 / 6.81$\pm$ 0.01 6.20$\pm$ 0.10 / 6.08$\pm$ 0.14 IFTPP-T 12.8 37.68$\pm$ 0.03 4.30$\pm$ 0.03 0.640$\pm$ 0.001 6.84$\pm$ 0.02 / 6.85$\pm$ 0.02 5.80$\pm$ 0.27 / 5.55$\pm$ 0.19 RMTPP 8.7 38.28$\pm$ 0.04 4.64$\pm$ 0.02 0.693$\pm$ 0.005 6.68$\pm$ 0.01 / 6.70$\pm$ 0.01 7.41$\pm$ 0.35 / 7.12$\pm$ 0.10 NHP 9.4 35.47$\pm$ 0.12 3.36$\pm$ 0.02 0.693$\pm$ 0.005 6.99$\pm$ 0.03 / 6.99$\pm$ 0.03 5.66$\pm$ 0.17 / 5.48$\pm$ 0.12 AttNHP 6.8 30.68$\pm$ x.xx 1.09$\pm$ x.xx 0.749$\pm$ x.xxx 7.47$\pm$ x.xx / 7.48$\pm$ x.xx 1.50$\pm$ x.xx / 1.49$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 7.3 22.50$\pm$ x.xx 4.62$\pm$ x.xx 0.694$\pm$ x.xxx 7.62$\pm$ x.xx / 7.62$\pm$ x.xx 8.20$\pm$ x.xx / 7.54$\pm$ x.xx Diffusion 10.4 36.05$\pm$ 0.07 3.66$\pm$ 0.03 0.650$\pm$ 0.002 6.86$\pm$ 0.01 / 6.88$\pm$ 0.00 5.95$\pm$ 0.05 / 6.04$\pm$ 0.08 MIMIC-IV MostPopular 2.7 4.80$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.75$\pm$ 0.00 / 19.75$\pm$ 0.00 1.63$\pm$ 0.00 / 1.56$\pm$ 0.00 HistoryDensity 160.0 0.74$\pm$ 0.00 2.63$\pm$ 0.00 3.23$\pm$ 0.00 18.94$\pm$ 0.00 / 18.92$\pm$ 0.00 4.25$\pm$ 0.00 / 4.19$\pm$ 0.00 IFTPP 12.4 58.52$\pm$ 0.06 46.04$\pm$ 0.26 3.00$\pm$ 0.02 11.53$\pm$ 0.03 / 11.55$\pm$ 0.03 22.56$\pm$ 0.13 / 22.31$\pm$ 0.15 IFTPP-T 11.7 58.57$\pm$ 0.04 47.96$\pm$ 1.51 3.03$\pm$ 0.01 11.46$\pm$ 0.02 / 11.48$\pm$ 0.02 24.07$\pm$ 0.31 / 23.68$\pm$ 0.29 RMTPP 9.8 58.24$\pm$ 0.14 45.62$\pm$ 1.28 3.93$\pm$ 0.03 13.69$\pm$ 0.05 / 13.76$\pm$ 0.06 22.00$\pm$ 0.28 / 21.62$\pm$ 0.29 NHP 3.0 23.19$\pm$ 0.61 11.67$\pm$ 0.39 4.71$\pm$ 0.31 17.92$\pm$ 0.34 / 17.92$\pm$ 0.33 9.76$\pm$ 0.41 / 9.80$\pm$ 0.38 AttNHP 6.2 44.99$\pm$ 1.11 31.88$\pm$ 0.64 2.95$\pm$ 0.20 14.30$\pm$ 0.16 / 14.31$\pm$ 0.16 19.98$\pm$ 0.79 / 19.74$\pm$ 0.71 ODE 11.6 44.16$\pm$ 0.76 27.48$\pm$ 1.04 2.91$\pm$ 0.01 14.36$\pm$ 0.20 / 14.39$\pm$ 0.20 16.37$\pm$ 0.75 / 16.19$\pm$ 0.74 HYPRO 7.4 45.69$\pm$ x.xx 45.54$\pm$ x.xx 3.98$\pm$ x.xx 14.68$\pm$ x.xx / 14.75$\pm$ x.xx 16.41$\pm$ x.xx / 16.13$\pm$ x.xx Diffusion 15.4 46.22$\pm$ 5.31 33.67$\pm$ 3.66 2.99$\pm$ 0.01 13.25$\pm$ 0.15 / 13.28$\pm$ 0.14 22.90$\pm$ 0.10 / 22.82$\pm$ 0.12 Retweet MostPopular 9.9 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 141.7$\pm$ 0.0 / 139.1$\pm$ 0.0 29.72$\pm$ 0.00 / 30.02$\pm$ 0.00 HistoryDensity 14.9 58.62$\pm$ 0.00 39.86$\pm$ 0.00 20.68$\pm$ 0.00 178.5$\pm$ 0.0 / 177.0$\pm$ 0.0 44.48$\pm$ 0.00 / 44.60$\pm$ 0.00 IFTPP 23.1 59.75$\pm$ 0.25 46.23$\pm$ 0.31 18.27$\pm$ 0.05 165.8$\pm$ 0.2 / 164.8$\pm$ 0.3 43.93$\pm$ 0.69 / 40.59$\pm$ 0.84 IFTPP-T 23.2 59.88$\pm$ 0.06 46.07$\pm$ 0.11 18.47$\pm$ 0.06 167.4$\pm$ 1.3 / 166.1$\pm$ 1.1 43.73$\pm$ 1.88 / 40.05$\pm$ 1.73 RMTPP 16.6 60.01$\pm$ 0.06 46.81$\pm$ 0.03 18.60$\pm$ 0.04 169.2$\pm$ 1.0 / 168.0$\pm$ 1.0 49.03$\pm$ 0.15 / 45.60$\pm$ 0.28 NHP 18.1 60.05$\pm$ 0.08 46.84$\pm$ 0.13 18.39$\pm$ 0.12 167.0$\pm$ 1.7 / 165.7$\pm$ 1.7 49.25$\pm$ 0.37 / 45.64$\pm$ 0.26 AttNHP 29.8 60.06$\pm$ 0.03 46.66$\pm$ 0.02 18.32$\pm$ 0.02 174.7$\pm$ 0.7 / 172.9$\pm$ 0.7 26.74$\pm$ 1.16 / 24.24$\pm$ 0.94 ODE 17.7 59.93$\pm$ 0.06 46.36$\pm$ 0.30 18.53$\pm$ 0.09 168.2$\pm$ 1.1 / 166.7$\pm$ 1.0 48.95$\pm$ 0.30 / 45.45$\pm$ 0.38 HYPRO 16.2 50.84$\pm$ x.xx 46.84$\pm$ x.xx 18.61$\pm$ x.xx 161.8$\pm$ x.x / 160.4$\pm$ x.x 50.58$\pm$ x.xx / 46.87$\pm$ x.xx Diffusion 14.7 57.72$\pm$ 0.77 42.96$\pm$ 0.67 18.60$\pm$ 0.06 159.0$\pm$ 1.4 / 158.0$\pm$ 1.1 54.24$\pm$ 0.89 / 52.24$\pm$ 0.68 Amazon MostPopular 5.0 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 6.33$\pm$ 0.00 / 6.30$\pm$ 0.00 10.10$\pm$ 0.00 / 9.72$\pm$ 0.00 HistoryDensity 71.8 32.68$\pm$ 0.00 9.63$\pm$ 0.00 0.325$\pm$ 0.000 7.29$\pm$ 0.00 / 7.27$\pm$ 0.00 28.25$\pm$ 0.00 / 27.53$\pm$ 0.00 IFTPP 12.8 35.54$\pm$ 0.08 16.72$\pm$ 0.04 0.250$\pm$ 0.001 6.70$\pm$ 0.01 / 6.64$\pm$ 0.01 23.05$\pm$ 0.48 / 23.79$\pm$ 0.39 IFTPP-T 13.1 35.27$\pm$ 0.04 16.46$\pm$ 0.07 0.245$\pm$ 0.001 6.64$\pm$ 0.02 / 6.59$\pm$ 0.02 22.81$\pm$ 0.10 / 23.42$\pm$ 0.13 RMTPP 16.4 35.85$\pm$ 0.03 17.21$\pm$ 0.04 0.292$\pm$ 0.002 6.67$\pm$ 0.03 / 6.62$\pm$ 0.03 19.65$\pm$ 0.23 / 19.99$\pm$ 0.18 NHP 9.9 16.58$\pm$ 3.26 11.19$\pm$ 0.18 0.438$\pm$ 0.009 8.54$\pm$ 0.28 / 8.52$\pm$ 0.30 27.19$\pm$ 0.35 / 26.95$\pm$ 0.37 AttNHP 12.4 22.21$\pm$ 5.64 8.22$\pm$ 0.39 0.526$\pm$ 0.075 7.63$\pm$ 0.12 / 7.61$\pm$ 0.14 19.29$\pm$ 1.45 / 18.59$\pm$ 1.52 ODE 9.4 10.44$\pm$ 2.84 10.28$\pm$ 0.57 0.473$\pm$ 0.031 9.23$\pm$ 0.17 / 9.23$\pm$ 0.18 23.73$\pm$ 0.36 / 23.13$\pm$ 0.46 HYPRO 17.8 21.47$\pm$ x.xx 17.14$\pm$ x.xx 0.288$\pm$ x.xxx 6.95$\pm$ x.xx / 6.95$\pm$ x.xx 20.47$\pm$ x.xx / 20.28$\pm$ x.xx Diffusion 13.3 30.57$\pm$ 1.28 13.36$\pm$ 0.49 0.247$\pm$ 0.000 6.56$\pm$ 0.03 / 6.52$\pm$ 0.04 30.08$\pm$ 0.25 / 30.29$\pm$ 0.32 StackOverflow MostPopular 9.8 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.602$\pm$ 0.000 13.12$\pm$ 0.00 / 13.15$\pm$ 0.00 8.97$\pm$ 0.00 / 8.16$\pm$ 0.00 HistoryDensity 74.0 42.51$\pm$ 0.00 5.61$\pm$ 0.00 0.669$\pm$ 0.000 17.30$\pm$ 0.00 / 17.29$\pm$ 0.00 16.23$\pm$ 0.00 / 14.90$\pm$ 0.00 IFTPP 24.1 45.31$\pm$ 0.10 11.59$\pm$ 0.70 0.646$\pm$ 0.003 13.56$\pm$ 0.19 / 13.66$\pm$ 0.19 9.68$\pm$ 0.55 / 8.87$\pm$ 0.59 IFTPP-T 23.0 45.40$\pm$ 0.04 12.78$\pm$ 0.29 0.639$\pm$ 0.001 13.51$\pm$ 0.07 / 13.61$\pm$ 0.05 9.62$\pm$ 0.15 / 8.81$\pm$ 0.20 RMTPP 12.7 45.30$\pm$ 0.07 12.61$\pm$ 0.40 0.717$\pm$ 0.006 13.00$\pm$ 0.03 / 13.25$\pm$ 0.04 13.81$\pm$ 0.30 / 13.35$\pm$ 0.28 NHP 11.9 44.12$\pm$ 0.79 9.90$\pm$ 1.16 0.735$\pm$ 0.021 13.27$\pm$ 0.31 / 13.46$\pm$ 0.27 13.12$\pm$ 0.26 / 12.41$\pm$ 0.31 AttNHP 15.6 45.17$\pm$ 0.16 12.56$\pm$ 0.15 0.703$\pm$ 0.000 13.10$\pm$ 0.03 / 13.31$\pm$ 0.01 11.95$\pm$ 0.38 / 11.15$\pm$ 0.29 ODE 13.3 44.19$\pm$ 0.13 9.08$\pm$ 0.27 0.728$\pm$ 0.007 13.12$\pm$ 0.05 / 13.35$\pm$ 0.05 11.48$\pm$ 0.29 / 10.75$\pm$ 0.21 HYPRO 11.8 30.54$\pm$ 0.70 12.61$\pm$ 0.40 0.718$\pm$ 0.006 14.00$\pm$ 0.19 / 14.12$\pm$ 0.17 15.25$\pm$ 0.28 / 14.84$\pm$ 0.23 Diffusion 14.4 37.03$\pm$ 1.24 7.48$\pm$ 0.16 0.668$\pm$ 0.006 12.87$\pm$ 0.17 / 13.01$\pm$ 0.16 15.80$\pm$ 0.85 / 15.07$\pm$ 0.74
+5 Experiments
 The main evaluation results are presented in Table 2. In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
-5.1. Long-horizon Prediction Quality
+5.1 Long-horizon Prediction Quality
 According to Table 2, autoregressive approaches such as IFTPP, RMTPP, and NHP excel at next-event prediction (MAE and accuracy), while horizon prediction methods like HYPRO and Diffusion perform better in terms of OTD and T-mAP.
@@
 Consequently, the potential of Transformer architectures in the MTPP domain remains insufficiently explored.
-Figure 5. Comparison of predicted label entropies, with the dashed horizontal line indicating the ground truth entropy.
-5.2. Metrics Comparison
+5.2 Metrics Comparison
 Table 2 shows that the OTD metric can yield low values even for simple rule-based baselines.
@@
 Meanwhile, the strong correlation of next-event prediction quality and long-horizon performance on MIMIC-IV can be attributed to the large number of events with identical timestamps, which makes precise event ordering crucial and favors autoregressive approaches like IFTPP.
-Table 3. Examples of predicted sequences (labels only).
-Method Transactions MIMIC-IV Retweet Amazon StackOverflow IFTPP 6, 6, 6, 6, 6, 6 14, 12, 4, 6, 11, 10, 1, 3, 5, 8, 9, 16 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 11, 0, 11, 0, 0, 0, 0, 0, 0, 0, 0 7, 17, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 RMTPP 6, 6, 6, 6, 6, 6 14, 4, 6, 11 0, 0, 0, 0, 0, 0, 0 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11 17, 3, 3, 3, 3, 3 NHP 6, 6, 6, 6, 6 4, 6 1, 1, 1, 1, 1, 1, 1, 1 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 AttNHP 1, 1, 1, 1, 1 4, 6, 10, 1 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 3, 0, 4, 1, 1, 1, 0, 0, 0, 0, 0, 0 3, 3, 3, 3, 3, 3, 3 ODE 1, 6, 6, 6, 6, 6 4, 6, 1, 3, 5, 2, 7, 4, 6 0, 0, 0, 0, 0, 0, 0 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 HYPRO 3, 6, 6, 1, 6, 1, 6 24, 1, 3, 5, 17, 8, 2, 20 0, 0, 1, 0, 0, 1, 0 1, 11, 0, 0, 11, 11, 11, 5, 1, 0, 5, 11 17, 5, 3, 5, 13, 3, 5 Diffusion 6, 6, 6, 6, 6, 6, 6 14, 4, 6, 1, 1, 3, 3, 3, 8, 3, 2, 2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 0, 0, 8, 5, 11, 11, 5, 5, 0, 5, 5, 11 3, 3, 3, 3, 3, 3, 3, 3
-5.3. Mode Collapse
-We found that most models tend to predict a single label or repetitive patterns in long-horizon tasks, as shown in Table 3.
-To quantify this effect, we measured the entropies of predicted labels (Fig.5).
+5.3 Mode Collapse
+Figure 5: Comparison of predicted label entropies for different methods, with the dashed horizontal line indicating the ground truth entropy. The results for all 5 datasets are presented in K.
+We found that most models tend to predict a single label or repetitive patterns in long-horizon tasks (see examples from K).
+To quantify the effect of mode collapse, we measured the entropies of predicted labels, as illustratedin Figure 5.
 Autoregressive methods generally exhibit low prediction diversity across all datasets except MIMIC-IV, where strong dependencies enable the model to learn typical sequences.
@@
 Sampling with an increased temperature can potentially solve this issue by offering a trade-off between predictions quality and diversity.
-Additional examples of predicted sequences are presented in Appendix G.
-6. Limitations and Future Work
-The proposed benchmark has several shortcomings.
-The training process for some methods is relatively slow.
-For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future.
-Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training.
-Additionally, the list of implemented methods can be extended, for example, by incorporating multi-task models (Deshpande et al., 2021).
-Our benchmark encourages future research to develop methods capable of solving next-event and long-horizon prediction tasks simultaneously, as discussed in Section 5.1.
-One promising direction involves adapting concepts from our statistical baselines for deep models, thereby surpassing baseline performance.
-A key question in this effort is the correct alignment between predicted and ground truth events during training, as discussed in the context of T-mAP (Section 3).
-Future work could also target the reduction of mode collapse and the enhancement of predictions diversity, as highlighted in Section 5.3.
-On the other hand, the T-mAP metric can be applied in domains beyond event sequences, such as action recognition (Su and Hua, 2017).
-T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW.
-T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error.
-Our theoretical justification of T-mAP can potentially be adapted for mAP estimation algorithms in computer vision.
-7. Conclusion
-In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting.
+6 Conclusion
+In this paper, we proposed the HoTPP benchmark to assess the quality of long-horizon events forecasting.
 HoTPP provides optimized implementations of the state-of-the-art methods ranging from statistical baselines to ODE and Diffusion (Sec.4), along with diverse metrics for both next-event and long-horizon prediction (Sec.3).
@@
 This advance has the potential to substantially enhance practical applications in various industries, driving both innovation and more informed decision-making.
+7 Limitations and Future Work
+The proposed benchmark has several shortcomings.
+The training process for some methods is relatively slow.
+For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future.
+Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training.
+Additionally, the list of implemented methods can be extended, for example, by incorporating multi-task models [6].
+Our benchmark encourages future research to develop methods capable of solving next-event and long-horizon prediction tasks simultaneously, as discussed in Section 5.1.
+One promising direction involves adapting concepts from our statistical baselines for deep models, thereby surpassing baseline performance.
+A key question in this effort is the correct alignment between predicted and ground truth events during training, as discussed in the context of T-mAP (Section 3).
+Future work could also target the reduction of mode collapse and the enhancement of predictions diversity, as highlighted in Section 5.3.
+On the other hand, the T-mAP metric can be applied in domains beyond event sequences, such as action recognition [28].
+T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW.
+T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error.
+Our theoretical justification of T-mAP can potentially be adapted for mAP estimation algorithms in computer vision.
 Appendix A Modeling Marked Temporal Point Processes
@@
 […]
-The derivation is provided in (Rizoiu et al., 2017).
+The derivation is provided in [24].
 Different TPPs are characterized by their intensity function $\lambda(t_{i})$.
@@
 A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
-(5) $\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$
+$\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$ (5)
 where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function.
@@
 Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach.
-For details on the implementation of sampling, please refer to (Rizoiu et al., 2017).
+For details on the implementation of sampling, please refer to [24].
 Recent research has focused on modeling complex intensity functions.
@@
 These approaches differ in the type of neural network used and the model of the intensity function.
-Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei and Eisner, 2017) and Neural ODEs (Rubanova et al., 2019).
+Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP [19] and Neural ODEs [25].
 The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
 Some methods evaluate the next event time distribution without using intensity functions.
-For example, intensity-free (Shchur et al., 2019) represents the distribution as a mixture of Gaussians or through normalizing flows.
+For example, intensity-free [27] represents the distribution as a mixture of Gaussians or through normalizing flows.
 Instead of predicting the distribution directly, other approaches solve a regression problem using MAE or RMSE loss.
 However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction.
-Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution (Bishop, 1994).
+Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution [3].
 In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
 Rescoring with HYPRO.
-HYPRO (Xue et al., 2022) is an extension applicable to any sequence prediction method capable of sampling.
+HYPRO [34] is an extension applicable to any sequence prediction method capable of sampling.
 HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample.
@@
 Diffusion.
-Recently, diffusion-based approaches have been applied to NLP tasks (Li et al., 2022).
-This idea has been extended to event sequences by adapting best practices from NLP (Zhou et al., 2025).
+Recently, diffusion-based approaches have been applied to NLP tasks [15].
+This idea has been extended to event sequences by adapting best practices from NLP [40].
 In such methods, events are first mapped to continuous embeddings, where the diffusion process operates, then mapped back to event space.
@@
 Training involves a combination of MSE loss in the continuous space, MAE loss for time reconstruction, and cross-entropy loss for label prediction.
-In HoTPP, we provide the first public implementation of the approach proposed by (Zhou et al., 2025).
+In HoTPP, we provide the first public implementation of the approach proposed by [40].
 MostPopular Baseline.
@@
 Consequently, for long-horizon forecasting, it assigns soft probabilities to each predicted event, thereby maximizing the T-mAP score.
-A.1. Methods Excluded from Comparison
-We attempted to incorporate ContiFormer (Chen et al., 2023) into our benchmark but encountered multiple challenges.
+Methods excluded from comparison.
+We attempted to incorporate ContiFormer [4] into our benchmark but encountered multiple challenges.
 The official implementation provides scripts only for training on a toy spiral dataset, making it impossible to reproduce the method’s performance on real-world data.
 Moreover, training even on this toy example required approximately three hours on an NVIDIA RTX 4060, suggesting that ContiFormer cannot be efficiently scaled to large datasets such as Transactions.
-In addition, contrary to the claims in the paper, ODE (Rubanova et al., 2019) and continuous-time transformers (Yang et al., 2022) were already discussed in earlier work.
+In addition, contrary to the claims in the paper, ODE [25] and continuous-time transformers [36] were already discussed in earlier work.
 Therefore, we excluded ContiFormer from our benchmark in favor of ODE and AttNHP.
-We also attempted to include NJDTPP (Zhang et al., 2024) in our benchmark, but encountered several challenges.
+We also attempted to include NJDTPP [38] in our benchmark, but encountered several challenges.
 Even on StackOverflow, our smallest dataset, the inference time spanned two hours for a single test set prediction.
@@
 Then any optimal matching $m^{\prime}\in M(\mathcal{G}^{\prime})$ will either have a size greater than the sizes of matchings in $M(\mathcal{G})$ or have a total weight equal to that of matchings in $M(\mathcal{G})$.
+Figure 6: Illustration of the iterative process in Lemma A.
 Proof.
-Figure 6. Illustration of the iterative process in Lemma A.
 If the size of $m^{\prime}$ is greater than the size of matchings in $M(\mathcal{G})$, then the lemma holds.
@@
 If $m^{\prime}$ includes vertex $b$, then we can walk through the graph with the following process:
-  1. (1)
+  1. 1.
     Start from the new vertex $b$.
-  2. (2)
+  2. 2.
     If we are at a ground truth event $b_{i}$ and there is an edge $(a_{i},b_{i})$ in matching $m^{\prime}$, move to the vertex $a_{i}$.
-  3. (3)
+  3. 3.
     If we are at a prediction $a_{i}$ and there is an edge $(a_{i},b_{i})$ in matching $m$, move to the vertex $b_{i}$.
@@
 For any threshold $h$, there exists an optimal matching in the graph $\mathcal{G}_{h}$ such that it is a subset of an optimal matching in the full graph $\mathcal{G}$:
-(6) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
+$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (6)
 Proof.
@@
 If there is no edge containing vertex $i$ in the matching $m(\mathcal{G})$, then $m(\mathcal{G})$ is also the optimal matching in $\mathcal{G}_{h}$ and theorem holds.
-Consider the case when matching $m(\mathcal{G})$ contains an edge $(i,j)$.
-Let $\hat{m}_{h}=m(\mathcal{G})\setminus\{(i,j)\}$, i.e., the matching without edge $(i,j)$.
-If it is optimal for the graph $\mathcal{G}_{h}$, then it will satisfy the theorem.
-Otherwise, there is an optimal matching $m_{h}$ with either (a) more vertices or (b) a smaller total weight than $\hat{m}_{h}$.
+[…]
 In (a), matching $m_{h}$ has a size greater than the size of $\hat{m}_{h}$.
@@
 By definition, multiclass T-mAP is the average of the average precision (AP) values for each label $l$:
-(7) $\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$
+$\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$ (7)
 Consider AP computation for a particular label $l$. AP is computed as the area under the precision-recall curve:
-(8) $\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$
+$\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$ (8)
 where $\mathrm{Rec}_{i}$ is the $i$-th recall value in a sorted sequence and $\mathrm{Prec}$ is the corresponding precision. Iteration is done among all distinct recall values ($\mathrm{Rec}_{0}=0$).
@@
 The resulting algorithm for AP computation for each label $l$ involves the following steps:
-  1. (1)
+  1. 1.
     Compute the optimal matching between predictions and ground truth events.
-  2. (2)
+  2. 2.
     Collect (a) scores of matched predictions, (b) scores of unmatched predictions, and (c) the number of unmatched ground truth events.
-  3. (3)
+  3. 3.
     Assign a positive label to matched predictions and a negative label to unmatched ones.
-  4. (4)
+  4. 4.
     Evaluate maximum recall as the fraction of matched ground truth events.
-  5. (5)
+  5. 5.
     Find the area under the precision-recall curve for the constructed binary classification problem from item 3 and multiply it by the maximum recall value.
@@
 Specifically, the optimal matching seeks to minimize the total weight in the following form:
-(9) $C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$
+$C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$ (9)
 A linear transformation of logits with a positive scaling factor will adjust the total weight accordingly, but the optimal matching will remain unchanged.
 Since the matching is performed independently for each class, we conclude that T-mAP is invariant to linear calibration.
-Appendix C HoTPP Benchmark Details
-C.1. Architecture
+Appendix C T-mAP Hyperparameters
+The dependency of T-mAP on the $\delta$ parameter for different datasets is presented in Figure 7.
+Figure 7: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value.
+Appendix D T-mAP for Highly Irregular Sequences
+This section analyzes a toy dataset containing highly irregular event sequences.
+This dataset includes a single label, which aims to predict timestamps.
+Most time intervals in the dataset are zero, with only 5% of intervals equal to one.
+We compare three simple baselines:
+  • ZeroStep, which predicts events with timestamps identical to the last observed event (zero intervals);
+  • UnitStep, which predicts events with a unit time step (the largest time step in the dataset);
+  • MeanStep predicts events using the average time step computed from historical data.
+Evaluation results are shown in Fig.
+8.
+The results indicate that the MAE and OTD metrics assign the lowest error to the ZeroStep baseline, which simply repeats the last event without accounting for the dataset’s irregularity.
+In contrast, T-mAP identifies the MeanStep baseline as the most effective, as it is the only method that analyzes historical data and incorporates timestamp statistics (mean interval) into its predictions.
+These findings suggest that T-mAP is a more appropriate metric for assessing the ability of methods to predict irregular events.
+Figure 8: Comparison of simple baselines on the Toy dataset with highly irregular time intervals. MAE and OTD metrics represent error values, while T-mAP measures model quality.
+Appendix E T-mAP for Long-tail Prediction Problems
+In this section, we assess the capability of various evaluation metrics to capture long-tail prediction quality, specifically the ability of models to predict rare classes.
+Unlike OTD, T-mAP evaluates each class independently, allowing for different aggregation strategies.
+The standard T-mAP computes a macro average, where the quality for each event class contributes equally to the final score.
+Additionally, the HoTPP benchmark includes a weighted variant of T-mAP, where the weights are proportional to class frequencies.
+Fig.
+9 compares the performance of IFTPP and RMTPP on the Transactions dataset, which includes 203 classes.
+The results show that all metrics, except macro T-mAP, remain unaffected as the dataset size increases beyond 60 classes.
+In contrast, macro T-mAP effectively evaluates the ability of models to predict across all available classes, highlighting its suitability for long-tail prediction tasks.
+Figure 9: Comparison of various metrics on the Transactions dataset across subsets with fewer event classes.
+Appendix F HoTPP Benchmark Details
+F.1 Architecture
 HoTPP incorporates best practices from extensible and reproducible ML pipelines.
-It leverages PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments.
-Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility.
+It leverages PyTorch Lightning [9] as the core training library, ensuring reproducibility and portability across various computing environments.
+Additionally, HoTPP utilizes the Hydra configuration library [35] to enhance extensibility.
 The overall architecture is illustrated in Fig.
-7.
+10.
 HoTPP supports both discrete-time and continuous-time models as well as RNN and Transformer architectures.
 Implementing a new method requires only adding essential components, while the rest can be configured through Hydra and configuration files.
-Figure 7. The architecture of the HoTPP library.
-C.2. Metrics
+Figure 10: The architecture of the HoTPP library.
+F.2 Metrics
 MTPP models are typically evaluated based on the accuracy of next-event predictions.
@@
 Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for the IFTPP and Diffusion models.
-Previous works have advanced long-horizon evaluation using OTD (Mei et al., 2019; Xue et al., 2023).
+Previous works have advanced long-horizon evaluation using OTD [20, 33].
 In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.1.
-C.3. Backbones
+F.3 Backbones
 We use three types of architectures in our experiments.
-The first is the GRU network (Cho et al., 2014), one of the top-performing neural architectures for event sequences (Babaev et al., 2022).
-We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei and Eisner, 2017).
+The first is the GRU network [5], one of the top-performing neural architectures for event sequences [1].
+We also implement the continuous-time LSTM (CT-LSTM) from the NHP method [19].
 Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases.
 An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence.
-Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
-C.4. Datasets
+Lastly, we implement the AttNHP continuous-time transformer model [36].
+F.4 Datasets
 Transactions [^2] https://huggingface.co/datasets/dllllb/age-group-prediction, Retweet [^3] https://huggingface.co/datasets/easytpp/retweet, Amazon [^4] https://huggingface.co/datasets/easytpp/amazon, and StackOverflow [^5] https://huggingface.co/datasets/easytpp/stackoverflow datasets were obtained from the HuggingFace repository.
@@
 Additionally, we address the issue of multiple diagnoses occurring in a single event by sorting timestamps for reproducibility.
-C.5. Performance improvements
+Appendix G Performance improvements
 The HoTPP benchmark provides highly optimized training and inference procedures for the efficient evaluation of datasets containing up to tens of millions of events.
@@
 Additionally, HoTPP offers a GPU implementation of the Hungarian algorithm, which also finds applications in computer vision.
-Table 4. HoTPP computation speed improvements in terms of seconds per batch.
+Table 3: HoTPP computation speed improvements in terms of seconds per batch.
 Implementation Study
-RNN autoreg. inference CT-LSTM inference CT-LSTM train ODE inference ODE train
+RNN autoreg.
+inference
+CT-LSTM
+inference
+CT-LSTM
+train
+ODE
+inference
+ODE
+train
 Baseline 2.44 0.0158 0.0519 0.08 0.162
@@
 Impr. 17x 6x 2x 8x 3x
-Appendix D Training details
-We trained each model using the Adam optimizer, which has a learning rate 0.001, and a scheduler that reduces the learning rate by 20% after every five epochs.
-Gradient clipping was applied, and the maximum L2-norm was set to 1.
-We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060.
-Each method was trained on a single GPU.
-The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset and up to 15 hours for NHP on Transactions.
-Multi-seed evaluations took approximately five times longer to complete.
-Appendix E Hyperparameters
-Dataset-specific training parameters are listed in Table 5.
-Table 5. Training hyperparameters.
-Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens
-Transactions 30 1200 256 512 512 $\rightarrow$ 256
-MIMIC-IV 30 64 16 64 64
-Retweet 30 264 16 64 64
-Amazon 60 94 32 64 64
-StackOverflow 60 101 32 64 64
-Appendix F Datasets Analysis
+Appendix H Datasets Analysis
 HoTPP incorporates datasets from various domains, including finance, social networks, and healthcare. Below, we provide additional details highlighting the key differences among these datasets.
@@
 The low next-event accuracy of methods based on NHP loss (NHP, AttNHP, and ODE) on MIMIC-IV can be attributed to their independent modeling of each event class, which results in a random ordering of events with identical timestamps.
-Table 6. Time step percentiles.
+Table 4: Time step percentiles.
 Dataset 1% 5% 10% 50% 90% 95% 99%
@@
 StackOverflow 0.0003 0.01 0.05 0.51 2.16 2.97 5.08
+Appendix I Training details
+We trained each model using the Adam optimizer, which has a learning rate 0.001, and a scheduler that reduces the learning rate by 20% after every five epochs.
+Gradient clipping was applied, and the maximum L2-norm was set to 1.
+We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060.
+Each method was trained on a single GPU.
+The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset and up to 15 hours for NHP on Transactions.
+Multi-seed evaluations took approximately five times longer to complete.
+Appendix J Hyperparameters
+Dataset-specific training parameters are listed in Table 5:
+Table 5: Training hyperparameters.
+Dataset Num epochs Max Seq.
+Len.
+Label Emb.
+Size Hidden Size Head hiddens Transactions 30 1200 256 512 512 $\rightarrow$ 256 MIMIC-IV 30 64 16 64 64 Retweet 30 264 16 64 64 Amazon 60 94 32 64 64 StackOverflow 60 101 32 64 64
 Discrete Timestamps.
@@
 The exact values are provided in the configuration files included with the source code.
-Appendix G Qualitative Analysis of Predictions
+Appendix K Predictions Diversity
+Figure 11: Comparison of predicted label entropies, with the dashed horizontal line indicating the ground truth entropy.
 Table 6 presents example predictions from various methods on the Transactions, MIMIC-IV, and Retweet datasets.
@@
 Addressing the mentioned limitation in future research could lead to the development of methods capable of producing more varied and realistic sequences in high-uncertainty scenarios.
-Table 7. Example predictions (labels only).
-Method Seq. ID Transactions MIMIC-IV Retweet
-IFTPP 0 3, 1, 3, 1, 3, 1, 3, 3, 3 10, 27, 23, 22, 1, 27, 3, 28, 26, 25, 23, 30 1, 1
-1 6, 3, 6, 6, 6, 6, 6 11, 10 0, 0, 0, 0, 0, 0, 0, 0, 0
-2 3, 1, 3, 1, 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
-RMTPP 0 3, 1, 3, 1, 3, 1 10, 27, 23, 22, 1, 30, 27, 5, 8, 9, 16, 15 1, 1, 1, 1
-1 6, 6, 6, 6, 6, 6 11, 10, 1 0, 0, 0, 0, 0, 0, 0, 0
-2 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
-NHP 0 1, 1, 1, 1, 1, 3, 1, 1, 1, 1 6 1, 1, 1, 1
-1 1, 1, 6, 6, 6, 6, 6 6, 6, 6 1, 1, 1, 1, 1, 1, 1
-2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 6, 6 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
-ODE 0 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 10, 1, 3, 5, 2, 4, 6, 1 1, 1, 1, 1
-1 6, 6, 6, 6, 6 1, 3, 5, 2, 4, 6, 1, 3, 5, 2, 4, 6 0, 0, 0, 0, 0, 0, 0
-2 1, 1, 1, 1, 1, 1, 1 2, 4, 6, 1, 3, 5, 2, 4, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
-HYPRO 0 3, 1, 16, 3, 12, 1 23, 27, 22, 1, 28, 26, 25, 23 0, 0, 1, 0
-1 1, 32, 6, 6, 1, 6 1, 28, 25, 23, 3, 26, 22, 5 0, 0, 0, 1, 1, 0, 1
-2 3, 5, 1, 3, 5 15, 13, 2, 7, 4, 1, 3, 5 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0
-Diffusion 0 1, 3, 1, 1, 1, 3, 18, 1 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0
-1 6, 6, 6, 6, 6, 6, 6 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 0, 0, 1, 1, 1, 1, 0, 1, 0
-2 8, 1, 3, 8, 3, 8, 1, 3, 8, 1, 1, 8 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1
-Appendix H T-mAP for Highly Irregular Sequences
-This section analyzes a toy dataset containing highly irregular event sequences.
-This dataset includes a single label, which aims to predict timestamps.
-Most time intervals in the dataset are zero, with only 5% of intervals equal to one.
-We compare three simple baselines:
-  • ZeroStep, which predicts events with timestamps identical to the last observed event (zero intervals);
-  • UnitStep, which predicts events with a unit time step (the largest time step in the dataset);
-  • MeanStep predicts events using the average time step computed from historical data.
-Evaluation results are shown in Fig.
-8.
-The results indicate that the MAE and OTD metrics assign the lowest error to the ZeroStep baseline, which simply repeats the last event without accounting for the dataset’s irregularity.
-In contrast, T-mAP identifies the MeanStep baseline as the most effective, as it is the only method that analyzes historical data and incorporates timestamp statistics (mean interval) into its predictions.
-These findings suggest that T-mAP is a more appropriate metric for assessing the ability of methods to predict irregular events.
-Figure 8. Comparison of simple baselines on the Toy dataset with highly irregular time intervals. MAE and OTD metrics represent error values, while T-mAP measures model quality.
-Appendix I Long-tail Prediction
-In this section, we assess the capability of various evaluation metrics to capture long-tail prediction quality, specifically the ability of models to predict rare classes.
-Unlike OTD, T-mAP evaluates each class independently, allowing for different aggregation strategies.
-The standard T-mAP computes a macro average, where the quality for each event class contributes equally to the final score.
-Additionally, the HoTPP benchmark includes a weighted variant of T-mAP, where the weights are proportional to class frequencies.
-Fig.
-9 compares the performance of IFTPP and RMTPP on the Transactions dataset, which includes 203 classes.
-The results show that all metrics, except macro T-mAP, remain unaffected as the dataset size increases beyond 60 classes.
-In contrast, macro T-mAP effectively evaluates the ability of models to predict across all available classes, highlighting its suitability for long-tail prediction tasks.
-Figure 9. Comparison of various metrics on the Transactions dataset across subsets with fewer event classes.
+The quantitative results for all 5 datasets are presented in Figure 11. For details, please refer to Section 5.3.
+Table 6: Example predictions (labels only).
+Method Seq.
+ID Transactions MIMIC-IV Retweet IFTPP 0 3, 1, 3, 1, 3, 1, 3, 3, 3 10, 27, 23, 22, 1, 27, 3, 28, 26, 25, 23, 30 1, 1 1 6, 3, 6, 6, 6, 6, 6 11, 10 0, 0, 0, 0, 0, 0, 0, 0, 0 2 3, 1, 3, 1, 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 RMTPP 0 3, 1, 3, 1, 3, 1 10, 27, 23, 22, 1, 30, 27, 5, 8, 9, 16, 15 1, 1, 1, 1 1 6, 6, 6, 6, 6, 6 11, 10, 1 0, 0, 0, 0, 0, 0, 0, 0 2 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 NHP 0 1, 1, 1, 1, 1, 3, 1, 1, 1, 1 6 1, 1, 1, 1 1 1, 1, 6, 6, 6, 6, 6 6, 6, 6 1, 1, 1, 1, 1, 1, 1 2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 6, 6 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ODE 0 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 10, 1, 3, 5, 2, 4, 6, 1 1, 1, 1, 1 1 6, 6, 6, 6, 6 1, 3, 5, 2, 4, 6, 1, 3, 5, 2, 4, 6 0, 0, 0, 0, 0, 0, 0 2 1, 1, 1, 1, 1, 1, 1 2, 4, 6, 1, 3, 5, 2, 4, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 HYPRO 0 3, 1, 16, 3, 12, 1 23, 27, 22, 1, 28, 26, 25, 23 0, 0, 1, 0 1 1, 32, 6, 6, 1, 6 1, 28, 25, 23, 3, 26, 22, 5 0, 0, 0, 1, 1, 0, 1 2 3, 5, 1, 3, 5 15, 13, 2, 7, 4, 1, 3, 5 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0 Diffusion 0 1, 3, 1, 1, 1, 3, 18, 1 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1 6, 6, 6, 6, 6, 6, 6 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 0, 0, 1, 1, 1, 1, 0, 1, 0 2 8, 1, 3, 8, 3, 8, 1, 3, 8, 1, 1, 8 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1